### v1

# GIST: Greedy Independent Set Thresholding for Diverse Data Summarization

Abstract
We propose a novel subset selection task called min-distance diverse data summarization (MDDS), which has a wide variety of applications in machine learning, e.g., data sampling and feature selection. Given a set of points in a metric space, the goal is to maximize an objective that combines the total utility of the points and a diversity term that captures the minimum distance between any pair of selected points, subject to the constraint $|S|\leq k$. For example, the points may correspond to training examples in a data sampling problem, e.g., learned embeddings of images extracted from a deep neural network. This work presents the GIST algorithm, which achieves a $\nicefrac{{2}}{{3}}$-approximation guarantee for MDDS by approximating a series of maximum independent set problems with a bicriteria greedy algorithm. We also prove a complementary $(\nicefrac{{2}}{{3}}+\varepsilon)$-hardness of approximation, for any $\varepsilon>0$. Finally, we provide an empirical study that demonstrates GIST outperforms existing methods for MDDS on synthetic data, and also for a real-world image classification experiment the studies single-shot subset selection for ImageNet.
1 Introduction
Subset selection is a challenging optimization problem with a wide variety of applications in machine learning, including feature selection, recommender systems, news aggregation, drug discovery, data summarization, and designing pretraining sets for large language models (Anil et al., 2023). Data sampling in particular is a salient problem due to unprecedented and continuous data collection. For example, LiDAR and imaging devices in one self-driving vehicle can easily capture ~80 terabytes of data per day (Kazhamiaka et al., 2021).
In most subset selection tasks, we rely on the weight (or utility) of the objects to rank one over the other, and also to avoid selecting duplicate or near-duplicate objects. If we select a small subset, then we also want to ensure that the selected subset is a good representation of the original set. These utility, diversity, and coverage criteria can be expressed through objective functions, and the interesting research lies in developing efficient algorithms with strong approximation guarantees. The underlying machinery used in constrained subset selection algorithms shares many similarities with techniques from other areas of combinatorial optimization such as submodular maximization, $k$-center clustering, and convex hull approximations.
In this work, we study the problem of selecting a set of points in a metric space that maximizes an objective that combines their utility and a minimum pairwise-distance diversity measure. Concretely, given a set of points $V$, weight function $w:V\rightarrow\mathbb{R}_{\geq 0}$, and distance function $\textnormal{dist}:V\times V\rightarrow\mathbb{R}_{\geq 0}$, our goal is to solve the cardinality-constrained optimization problem:
$S^{*}$ $=\operatorname*{arg\,max}_{S\subseteq V}~{}\frac{1}{k}\sum_{v\in S}w(v)+\texttt{div}(S)$
subject $\hskip 13.6572pt\text{subject to}~{}~{}|S|\leq k$
where the diversity term $\texttt{div}(S)\coloneqq\min_{u,v\in S:u\neq v}\textnormal{dist}(u,v)$ encourages maximizing the minimum distance between all pairs of distinct points in the selected subset. Bhaskara et al. (2016) call $\texttt{div}(S)$ the min-min diversity objective and Meinl et al. (2011) call it $p$-dispersion maximization, where it is known to admit a $\nicefrac{{1}}{{2}}$-approximation ratio subject to the strict cardinality constraint $|S|=k$.
Other diversity objectives such as the sum-of-min-distances $\sum_{u\in S}\min_{v\in S\setminus\{u\}}\textnormal{dist}(u,v)$ and sum-of-all-distances $\sum_{u\in S}\sum_{v\in S}\textnormal{dist}(u,v)$ have also been studied (Meinl et al., 2011; Bhaskara et al., 2016). While previous works have looked at maximizing the diversity objective $\texttt{div}(S)$, our formulation that introduces the utility term has not been studied before. In particular, Bhaskara et al. (2016) observe that min-min is a powerful diversity objective, but point out a major challenge: it is highly non-monotonic, i.e., it can exhibit a stark decrease in the objective value by adding a single new point. Our work shows that the use of a utility term along with the min-min can alleviate some of these pitfalls, and allows us to inherit the strong benefits of this strict diversity objective.
1.1 Our contributions
We summarize our main contributions below:
  • We propose a novel subset selection problem called min-distance diverse data summarization (MDDS), where the goal is to maximize an objective function that combines a linear weight term with the min-min distance diversity term $\texttt{div}(S)$ subject to the cardinality constraint $|S|\leq k$. This problem has a wide variety of applications in machine learning, e.g., data denoising and active learning.
  • We give a $\nicefrac{{2}}{{3}}$-approximation algorithm for MDDS called GIST that approximates a series of maximum independent set problems on geometric intersection graphs with a bicriteria greedy algorithm and returns the best solution.
  • We prove a complementary $(\nicefrac{{2}}{{3}}+\varepsilon)$-hardness of approximation, for any $\varepsilon>0$, for general metric spaces by reducing from approximate maximum clique. We also prove that MDDS is APX-complete for Euclidean metrics.
  • Our experiments show that GIST overcomes the shortcomings of existing methods for MDDS on synthetic data. We also demonstrate how GIST can be used to create better single-shot subsets of training data for an image classification benchmark on ImageNet (Russakovsky et al., 2015) compared to margin sampling and $k$-center algorithms.
1.2 Background and related work
Diversity objectives.
A closely related problem is $k$-center, which aims to find a subset of $k$ centers that minimize the largest distance of any point to its closest center in $S$. Concretely, the objective is $\max_{v\in V}\min_{c\in S}\textnormal{dist}(v,c)$, and a greedy $2$-approximation algorithm is widely used in active learning applications (Sener and Savarese, 2018). The main difference between our objective function and $k$-center is the ability to benefit from the utility term, e.g., uncertainty values in an active learning setting. The second difference is that we relax the need to cover every point in the original dataset, which is particularly beneficial for large datasets, where we prefer to avoid instability due to certain corner or outlier points.
Bhaskara et al. (2016) study various distance-based diversity metrics subject to a cardinality constraint. In particular, they give a $\nicefrac{{1}}{{8}}$-approximation algorithm for the sum-min diversity objective defined as $\sum_{u\in S}\min_{v\in S\setminus\{u\}}\textnormal{dist}(u,v)$ by rounding a linear programming relaxation, and they complement this result by proving a conditional $\nicefrac{{1}}{{2}}$-hardness of approximation under the planted clique assumption. They also study the min-min objective $\min_{u,v\in S:u\neq v}\textnormal{dist}(u,v)$, which is the same as $\texttt{div}(S)$. They do not have a linear weight function in their objective, so they enforce a strict cardinality constraint $|S|=k$; otherwise, it is always better to output at most one or two points. Meinl et al. (2011) study the same objective subject to $|S|=k$ under the name $p$-dispersion (also called max-min or remote-edge) diversity maximization. This cardinality issue does not arise in our formulation since we need to optimize the trade-off between competing linear weight (utility) and diversity terms.
Moumoulidou et al. (2020), Addanki et al. (2022), and Mahabadi and Trajanovski (2023) initiated the study of diversity maximization under fairness constraints. Given a partitioning of the dataset, they ensure that a prespecified number of points $k_{i}$ are selected from each group $i\in[m]$. This fairness constraint considerably differentiates their problem setting from ours (i.e., $|S|=k=\sum_{i=1}^{m}k_{i}$). Moumoulidou et al. (2020); Addanki et al. (2022) maximize the $\texttt{div}(S)$ objective, and Mahabadi and Trajanovski (2023) consider the sum-of-pairwise distances $\sum_{u\in S}\sum_{v\in S}\textnormal{dist}(u,v)$ and sum-min objectives. None of these works, however, add a linear weight function to the objective.
Data sampling.
In the realm of dataset curation and active learning, there have been several lines of work that study the combination of utility and diversity terms. Ash et al. (2020) invoke $k$-means++ seeding over the gradient space of a model in order to balance uncertainty and diversity. Wei et al. (2015) introduce several submodular objectives, e.g., facility location, and use them to diversify a set of uncertain examples in each active learning iteration. Citovsky et al. (2021) cluster examples which are represented by embeddings extracted from the penultimate layer of a partially trained DNN, and then use these clusters to diversify uncertain examples in each iteration. Our work differs from these and several others (e.g., Kirsch et al. (2019); Zhdanov (2019)) in that we directly incorporate both the utility and diversity terms in our objective function.
2 Preliminaries
We are given a set of $n$ points $V\subseteq M$ in the metric space $(M,\textnormal{dist})$, where $\textnormal{dist}:M\times M\rightarrow\mathbb{R}_{\geq 0}$ is the distance function. We are also given weights $w:V\rightarrow\mathbb{R}_{\geq 0}$ representing the utility of each point. We extend these functions to take subsets $S\subseteq V$ as input in the standard way: $w(S)=\sum_{v\in S}w(v)$ and $\textnormal{dist}(u,S)=\min_{v\in S}\textnormal{dist}(u,v)$. For the sake of completeness, we define $\textnormal{dist}(u,\varnothing)=\infty$.
Intersection graph.
Let $G_{d}(V)$ denote the intersection graph of $V$ for distance threshold $d$, i.e., the graph with node set $V$ and edge set
$E=\{(u,v)\in V^{2}:\textnormal{dist}(u,v)<d\}.$
This graph is well-defined for any metric space and distance $d\in\mathbb{R}$. Note that for any independent set $S$ of $G_{d}(V)$ and pair of distinct nodes $u,v\in S$, we have $\textnormal{dist}(u,v)\geq d$.
Definition 2.1.
This work focuses on the min-distance diversity reward function
$\texttt{div}(S)$ =$=\begin{cases}\min_{u,v\in S:u\neq v}\textnormal{dist}(u,v)&\text{if }|S|\geq 2,\\ \max_{u,v\in V}\textnormal{dist}(u,v)&\text{if }|S|\leq 1.\end{cases}$
This reward function encourages selecting a well-separated set of points (generalizing the notion of data deduplication). To make $\texttt{div}(S)$ well-defined if $S$ is empty or consists of a singleton point, we set the diversity to be the diameter of the whole ground set $V$, implying that it is always beneficial to select at least two points to maximize the value of the selected set.
For any $\alpha\in[0,1]$, we define the objective function
$f(S)=\alpha\frac{1}{k}\sum_{v\in S}w(v)+(1-\alpha)\texttt{div}(S),$ (1)
which allows us to control the trade-off between the average utility and the diversity of set $S$. The goal of the min-distance diverse data summarization problem (MDDS) is to maximize $f(S)$ subject to a cardinality constraint $k$:
$S^{*}=\operatorname*{arg\,max}_{S\subseteq V:|S|\leq k}f(S).$
Let $\textnormal{OPT}=f(S^{*})$ denote the optimal objective value.
Simple $\nicefrac{{1}}{{2}}$-approximation algorithm.
Equation 1 has two parts, each of which is easy to optimize in isolation. To maximize the $w(S)$ term, we simply select the $k$ heaviest points. In contrast, to maximize the $\texttt{div}(S)$ term, we select two points whose distance is the diameter of $V$. Outputting a subset that corresponds to the maximum of these two terms gives a $\nicefrac{{1}}{{2}}$-approximation guarantee. To see this, observe that
$\textnormal{ALG}_{\textnormal{simple}}$ $=\max\left\{\alpha\frac{1}{k}\max_{|S|\leq k}w(S),(1-\alpha)\max_{u,v\in V}\textnormal{dist}(u,v)\right\}$
$\geq\frac{1}{2}\cdot\alpha\frac{1}{k}\max_{|S|\leq k}w(S)+\frac{1}{2}\cdot(1-\alpha)\max_{u,v\in V}\textnormal{dist}(u,v)$
$\geq\frac{1}{2}\cdot\left(\alpha\frac{1}{k}w(S^{*})+(1-\alpha)\texttt{div}(S^{*})\right)=\frac{1}{2}\cdot\textnormal{OPT},$
where $\texttt{div}(S)$ is defined in Definition 2.1. Our main goals are to improve over this simple baseline and prove complementary hardness results.
3 Algorithm
We now present the GIST algorithm, which achieves a $(\nicefrac{{2}}{{3}}-\varepsilon)$-approximation for the MDDS problem. It works by repeatedly calling the GreedyIndependentSet heuristic to find an independent set of the intersection graph $G_{d}(V)$ for various distance thresholds $d$. The subroutine GreedyIndependentSet sorts the points by their weight in a non-increasing order breaking ties arbitrarily. It starts with an empty set of selected points, and then it scans the points according to the sorted order. For each point, GreedyIndependentSet selects the point as long as its distance to the set of selected points so far is at least $d$ and the number of selected points does not exceed $k$.
[…]
Algorithm 1 Diversified data summarization using greedy-weighted maximum independent sets.
1:function GIST(points $V$, weights $w:V\rightarrow\mathbb{R}_{\geq 0}$, budget $k$, error $\varepsilon$)
2:     Initialize $S\leftarrow\textnormal{{GreedyIndependentSet}}(V,w,0,k)$triangleright $k$ heaviest points
3:     Let $d_{\max}=\max_{u,v\in V}\textnormal{dist}(u,v)$ be the diameter of $V$
4:     Let $T\leftarrow\{u,v\}$ be two points such that $\textnormal{dist}(u,v)=d_{\max}$
5:     if $f(T)>f(S)$ and $k\geq 2$ then
6:         Update $S\leftarrow T$
7:     Let $D\leftarrow\{(1+\varepsilon)^{i}\cdot\varepsilon d_{\max}/2:(1+\varepsilon)^{i}\leq 2/\varepsilon\mbox{ and }i\in\mathbb{Z}_{\geq 0}\}$triangleright distance thresholds
8:     for threshold $d\in D$ do
9:         Set $T\leftarrow\textnormal{{GreedyIndependentSet}}(V,w,d,k)$
10:         if $f(T)\geq f(S)$ then
11:              Update $S\leftarrow T$               return $S$
1:function GreedyIndependentSet(points $V$, weights $w:V\rightarrow\mathbb{R}_{\geq 0}$, distance $d$, budget $k$)
2:     Initialize $S\leftarrow\varnothing$
3:     for $v\in V$ sorted by non-increasing weight do
4:         if $\textnormal{dist}(v,S)\geq d$ and $|S|<k$ then
5:              Update $S\leftarrow S\cup\{v\}$               return $S$
Theorem 3.1.
For any $\varepsilon>0$, GIST outputs a set $S\subseteq V$ with $|S|\leq k$ and $f(S)\geq(\nicefrac{{2}}{{3}}-\varepsilon)\cdot\textnormal{OPT}$.
The main building block in our design and analysis of GIST is the following lemma, which is inspired by the greedy $2$-approximation algorithm for metric $k$-centers.
Lemma 3.2.
[…]
Proof.
Let $k^{\prime}=|T|$ and $t_{1},t_{2},\dots,t_{k^{\prime}}$ be the points in $T$ in the order that GreedyIndependentSet selected them. Let $B_{i}=\{v\in V:\textnormal{dist}(t_{i},v)<d^{\prime}\}$ be the points in $V$ contained in the radius-$d^{\prime}$ open ball around $t_{i}$. First, we show that each $B_{i}$ contains at most one point in $S_{d}^{*}$. If this is not true, then some $B_{i}$ contains two different points $u,v\in S_{d}^{*}$. Since $\textnormal{dist}(\cdot,\cdot)$ is a metric, this means
$\textnormal{dist}(u,v)$ $\leq\textnormal{dist}(u,t_{i})+\textnormal{dist}(t_{i},v)$
$<d^{\prime}+d^{\prime}$
$\leq d/2+d/2$
$=d,$
which contradicts the assumption that $S_{d}^{*}$ is an independent set of $G_{d}(V)$. Note that it is possible to have $B_{i}\cap B_{j}\neq\varnothing$, for $i\neq j$, since these balls consider all points in $V$.
Now let $C_{i}\subseteq V$ be the set of uncovered (by the open balls) points that become covered when GreedyIndependentSet selects $t_{i}$. Concretely, $C_{i}=B_{i}\setminus(B_{1}\cup\cdots\cup B_{i-1})$. Each $C_{i}$ contains at most one point in $S_{d}^{*}$ since $|B_{i}\cap S_{d}^{*}|\leq 1$. Moreover, if $s\in C_{i}\cap S_{d}^{*}$, then $w(t_{i})\geq w(s)$ because the points are sorted in non-increasing order and selected if uncovered.
Let $A=C_{1}\cup\cdots\cup C_{k^{\prime}}$ be the set of points covered by the algorithm. For each point $s\in S_{d}^{*}\cap A$, there is exactly one covering set $C_{i}$ corresponding to $s$. It follows that
$\sum_{s\in S_{d}^{*}\cap A}w(s)\leq\sum_{i\in[k^{\prime}]:S_{d}^{*}\cap C_{i}\neq\varnothing}w(t_{i}).$ (2)
It remains to account for the points in $S_{d}^{*}\setminus A$. Note that if we have any such points, then $|T|=k$ since the points in $S_{d}^{*}\setminus A$ are uncovered at the end of the algorithm. Further, for any $t_{i}\in T$ and $s\in S_{d}^{*}\setminus A$, we have $w(t_{i})\geq w(s)$ since $t_{i}$ was selected and the points are sorted by non-increasing weight. Therefore, we can assign each $s\in S_{d}^{*}\setminus A$ to a unique $C_{i}$ such that $C_{i}\cap S_{d}^{*}=\varnothing$. It follows that
$\sum_{s\in S_{d}^{*}\setminus A}w(s)\leq\sum_{i\in[k^{\prime}]:S_{d}^{*}\cap C_{i}=\varnothing}w(t_{i}).$ (3)
Adding the two sums together in Equation 2 and (3) completes the proof. ∎
Equipped with this bicriteria approximation guarantee, we can now analyze the approximation ratio of the GIST algorithm.
Proof of Theorem 3.1.
Let $d^{*}$ be the minimum distance between two distinct points in $S^{*}$. There are two cases: $d^{*}<\varepsilon d_{\max}$ or $d^{*}\geq\varepsilon d_{\max}$. For the first case, outputting the $k$ heaviest points (Line 2) achieves a $(1-\varepsilon)$-approximation. To see this, observe that
$\textnormal{OPT}\geq(1-\alpha)d_{\max}>(1-\alpha)\cdot\frac{d^{*}}{\varepsilon}\implies\varepsilon\cdot\textnormal{OPT}>(1-\alpha)d^{*}.$
The sum of the $k$ heaviest points upper bounds $w(S^{*})$, so it follows that
$\textnormal{ALG}\geq\alpha\frac{1}{k}w(S^{*})=\textnormal{OPT}-(1-\alpha)d^{*}>(1-\varepsilon)\textnormal{OPT}.$
Therefore, we focus on the second case where $d^{*}\geq\varepsilon d_{\max}$. It follows that GIST considers a threshold $d=(1+\varepsilon)^{i}\cdot\varepsilon d_{\max}/2\in D$ such that
$\frac{d^{*}}{2}\in[d,(1+\varepsilon)d)\implies\frac{1}{1+\varepsilon}\cdot\frac{d^{*}}{2}<d\leq\frac{d^{*}}{2}.$
[…]
$f(T)$ $\geq\alpha\frac{1}{k}w(S^{*})+(1-\alpha)d\geq\alpha\frac{1}{k}w(S^{*})+(1-\alpha)\frac{d^{*}}{2(1+\varepsilon)}.$ (4)
The potential max-diameter update to $S$ on Line 5 gives us
$\textnormal{ALG}\geq(1-\alpha)d_{\max}\geq(1-\alpha)d^{*}.$ (5)
Putting Equation 4 and (5) together, for any $p\in[0,1]$, it holds that
ALG $\geq p\left[\alpha\frac{1}{k}w(S^{*})+(1-\alpha)\frac{d^{*}}{2(1+\varepsilon)}\right]+(1-p)(1-\alpha)d^{*}$
$=p\cdot\alpha\frac{1}{k}w(S^{*})+\left(1-p+\frac{p}{2(1+\varepsilon)}\right)\cdot(1-\alpha)d^{*}.$
To maximize the approximation ratio as $\varepsilon\rightarrow 0$, solve $p=1-p/2$ to get $p=2/3$. It follows that
ALG $\geq\frac{2}{3}\cdot\alpha\frac{1}{k}w(S^{*})+\left(1-\frac{2}{3}+\frac{1}{3(1+\varepsilon)}\right)\cdot(1-\alpha)d^{*}$ (6)
$=\frac{2}{3}\cdot\alpha\frac{1}{k}w(S^{*})+\frac{1}{3}\left(1+\frac{1}{1+\varepsilon}\right)\cdot(1-\alpha)d^{*}$
$\geq\frac{2}{3}\cdot\alpha\frac{1}{k}w(S^{*})+\frac{1}{3}\left(2-\varepsilon\right)\cdot(1-\alpha)d^{*}$
$\geq\frac{2}{3}\left(1-\varepsilon\right)\cdot\alpha\frac{1}{k}w(S^{*})+\frac{2}{3}\left(1-\varepsilon\right)\cdot(1-\alpha)d^{*}$
$\geq\left(\frac{2}{3}-\varepsilon\right)\cdot\textnormal{OPT},$
which completes the proof. ∎
Remark 3.3.
We can modify GIST to achieve an exact $\nicefrac{{2}}{{3}}$-approximation by trying all pairwise distance thresholds in $D=\{\textnormal{dist}(u,v)/2:u,v\in V\}$. This requires $O(n^{2})$ calls to GreedyIndependentSet compared to $O(\log_{1+\varepsilon}(1/\varepsilon))$, but otherwise the same analysis holds for $\varepsilon=0$.
Finally, we show how to implement GIST in nearly linear time for low-dimensional Euclidean metrics using cover trees for exact nearest-neighbor queries (Beygelzimer et al., 2006; Izbicki and Shelton, 2015) and a $(1-\varepsilon)$-approximation algorithm for point-set diameters $d_{\max}$ (Barequet and Har-Peled, 2001; Imanparast et al., 2018).
Theorem 3.4.
For $V\subseteq\mathbb{R}^{d}$ and Euclidean distance, GIST runs in time $O(n^{3}k+n^{2}d)$. Furthermore, for fixed dimension $d$, the running time can be improved to $O(n\log n\cdot\log_{1+\varepsilon}(1/\varepsilon)\cdot\textnormal{poly}(d)+1/\varepsilon^{d-1})$.
We give the proof in Section A.1 and note that the key property is that cover trees allow us to insert a point into a nearest-neighbor index in $O(c^{12}\log n)$ time and find a nearest neighbor to a query point in $O(c^{6}\log n)$ time, where $c=\Theta(d)$ is the doubling dimension of $d$-dimensional Euclidean space.
4 Hardness of approximation
[…]
  1. 1.
    There is no polynomial-time $(\nicefrac{{2}}{{3}}+\varepsilon)$-approximation algorithm for general metric spaces.
  2. 2.
    There is no polynomial-time $(\nicefrac{{3}}{{4}}+\varepsilon)$-approximation algorithm for general metric spaces, even if $k\geq|V|/4$.
  3. 3.
    APX-completeness for the Euclidean metric, i.e., there is no polynomial-time approximation scheme (PTAS) for this objective.
4.1 General metric spaces
Our first result builds on the hardness of approximating the size of the maximum clique, and it tightly complements the $\nicefrac{{2}}{{3}}$-approximation guarantee that we can achieve with GIST.
Theorem 4.1.
[…]
Proof.
[…]
  • YES instance: Graph $G$ has a clique of size $k$.
  • NO instance: Graph $G$ does not have a clique of size greater than $\delta k$.
We reduce this instance of the maximum clique decision problem to MDDS with objective function (1) as follows. Represent each vertex of graph $G$ with a point in our ground set. The distance between a pair of points is $2$ if there is an edge between their corresponding vertices in $G$, and it is $1$ otherwise.
Use the same threshold value of $k$ (in the YES and NO instance above) for the cardinality constraint on set $S$, and set each weight $w(v)=1$. In a YES instance, selecting a clique of size $k$ as set $S$ results in the maximum possible value of the objective:
$\textnormal{OPT}=\alpha\cdot\frac{1}{k}\cdot k+(1-\alpha)\cdot 2=2-\alpha.$ (7)
In a NO instance, the best objective value that can be achieved in polynomial-time is the maximum of the following two scenarios: (a) selecting $k$ points with minimum distance $1$, or (b) selecting at most $\delta k$ vertices forming a clique with minimum distance $2$. The maximum value obtained by any polynomial-time algorithm is then
ALG $=\max\left\{\alpha+(1-\alpha)\cdot 1,\alpha\cdot\delta+(1-\alpha)\cdot 2\right\}$
$=\max\left\{1,2-(2-\delta)\alpha\right\}.$
These two terms become equal for $\alpha=1/(2-\delta)$. Thus, the gap between the maximum value any algorithm can achieve in the NO case and the optimum value in the YES case is
$\frac{1}{2-\alpha}=\frac{1}{2-\nicefrac{{1}}{{(2-\delta)}}}=\frac{2-\delta}{3-2\delta}.$
To complete the proof, it suffices to show that the ratio above is at most $\nicefrac{{2}}{{3}}+\varepsilon$. We separate the $\nicefrac{{2}}{{3}}$ term as follows:
$\frac{2-\delta}{3-2\delta}=\frac{\nicefrac{{2}}{{3}}\cdot(3-2\delta)+\nicefrac{{\delta}}{{3}}}{3-2\delta}=\frac{2}{3}+\frac{\delta}{9-6\delta}.$
Therefore, we must choose a value of $\delta$ satisfying $\nicefrac{{\delta}}{{(9-6\delta)}}\leq\varepsilon$. Since $\delta\leq 1$, the denominator $9-6\delta$ is positive. Equivalently, we want to satisfy:
$\frac{9-6\delta}{\delta}=\frac{9}{\delta}-6\geq\frac{1}{\varepsilon}.$
By setting $\delta<9\varepsilon/(1+6\varepsilon)$, we satisfy the required inequality and achieve the inapproximability gap in the theorem statement. ∎
4.2 Large subsets
In many data curation problems, we are interested in instances of MDDS where $k=\Omega(|V|)$. Even in this restricted case, we can show hardness of approximation.
Theorem 4.2.
[…]
The proof of this theorem uses the vertex cover instance constructed in the proof of Håstad (2001, Theorem 8.1) and is given in Section B.1.
4.3 Euclidean metrics
[…]
Lemma 4.3 (Alimonti and Kann (2000, Theorem 3.2)).
The maximum independent set problem for graphs with degree at most $3$ is APX-complete.
Similar to our reduction from maximum clique in Section 4.1 for general metric spaces, this construction uses a node embedding function $h_{g}(v)$ to encode graph adjacency in Euclidean space.
Lemma 4.4.
Let $G=(V,E)$ be a simple undirected graph with $n=|V|$, $m=|E|$, and maximum degree $\Delta$. There exists an embedding $h_{G}:V\rightarrow\mathbb{R}^{n+m}$ such that if $\{u,v\}\in E$ then
$\lVert h_{G}(u)-h_{G}(v)\rVert_{2}\leq 1-\frac{1}{2(\Delta+1)},$
and if $\{u,v\}\not\in E$ then $\lVert h_{G}(u)-h_{G}(v)\rVert_{2}=1$.
Theorem 4.5.
The MDDS problem is APX-complete for Euclidean metrics.
We sketch the main idea of our reduction and defer the complete proofs of Lemma 4.4 and Theorem 4.5 to Appendix B. Let $\mathcal{I}(G)$ be the set of independent sets of $G$. A set of embedded nodes $S\subseteq V$ (with at least two nodes) of a graph $G$ with maximum degree $\Delta\leq 3$ has the property:
  • If $S\in\mathcal{I}(G)$: $\texttt{div}(S)=\min_{u,v\in S:u\neq v}\lVert h_{G}(u)-h_{G}(v)\rVert_{2}=1$.
  • If $S\not\in\mathcal{I}(G)$: $\texttt{div}(S)=\min_{u,v\in S:u\neq v}\lVert h_{G}(u)-h_{G}(v)\rVert_{2}\leq 1-1/(2(\Delta+1))\leq 1-\nicefrac{{1}}{{8}}$.
This gap between the two cases is at least $\nicefrac{{1}}{{8}}$ (i.e., a universal constant) for all such graphs, and allows us to show that MDDS inherits the APX-complete hardness of bounded-degree max independent set.
5 Experiments
5.1 Warm-up: Synthetic dataset
Setup.
We start with a simple dataset of weighted random points to see the shortcomings of existing methods for MDDS. For a fixed dimension $d$, we generate $n=1000$ points $\mathbf{x}_{i}\in\mathbb{R}^{d}$, where each $\mathbf{x}_{i}\sim\mathcal{N}(\mathbf{0},\mathbf{I}_{d})$ and each point is assigned an i.i.d. uniform continuous weight $w_{i}\sim\mathcal{U}_{[0,1]}$.
We consider three baseline algorithms: random, simple, and greedy. The random algorithm chooses a random set of $k$ points, permutes them, and returns the best prefix. simple is the $\nicefrac{{1}}{{2}}$-approximation algorithm in Section 2. greedy builds a set $S\subseteq V$ one element at a time by selecting the point with index $i^{*}_{t}=\operatorname*{arg\,max}_{i\in V\setminus S_{t-1}}f(S_{t-1}\cup\{i\})-f(S_{t-1})$ at each step $t\in[k]$. Then it returns the best prefix $S=\operatorname*{arg\,max}_{t\in[k]}f(S_{t})$, as this sequence of objective values is not necessarily monotone.
Figure 1: Synthetic data with $n=1000$ and $d=64$. Plots of $f(S_{\text{ALG}})$ for $k\in[n]$, $\alpha\in\{0.90,0.95\}$.
Results.
We plot the values of $f(S)$ for our baseline algorithms and GIST for all $k\in[n]$ in Figure 1. We set $\alpha\in\{0.90,0.95\}$ to balance the two competing objective terms (which are easy to optimize in isolation). [^2] If $\mathbf{x},\mathbf{y}\sim\mathcal{N}(\mathbf{0},\mathbf{I}_{d})$ are i.i.d., then $\mathbb{E}[\lVert\mathbf{x}-\mathbf{y}\rVert_{2}^{2}]=2d$, so $\frac{\alpha}{2}\approx(1-\alpha)\sqrt{2d}\implies\alpha\approx 1-\frac{1}{2\sqrt{2d}+1}=0.957$. There is a clear separation in solution quality for each $k\in[n]$ when $\alpha=0.95$. Note that these objective values are not directly comparable as a function of $k$ due to the $\nicefrac{{1}}{{k}}$ normalizing factor (hence the discontinuities), but both terms in the objective decreasing on average in $k$ for a fixed $\alpha$: mean utility and min distance of the selected points. GIST noticeably outperforms simple, greedy, and random. The poor performance of greedy is somewhat surprising since it is normally competitive for submodular-style objectives. simple also outperforms greedy for all but the extreme values of $k$, i.e., very small or close to $n$, but is always worse relative to GIST.
5.2 Image classification
The following data sampling experiment compares the top-1 image classification accuracy achieved by different single-shot subset selection algorithms.
Setup.
We use the standard vision dataset ImageNet (Russakovsky et al., 2015), which contains ~1.3 million images and 1000 classes. We select 10% of the data points at random and use them to train an initial ResNet-56 model $\bm{\theta}_{0}$ (He et al., 2016). Then we use this initial model to compute an uncertainty value and 2048-dimensional embedding for each example. The uncertainty value of $\mathbf{x}_{i}$ is $u_{i}=1-\Pr(y=b\mid\mathbf{x}_{i};~{}\bm{\theta}_{0})$, where $b$ is the best class label according to the initial model. Finally, we use the fast maximum inner product search of Guo et al. (2020) to build a $\Delta$-nearest neighbor graph in the embedding space using $\Delta=10$ and cosine distance. We present all of the model training hyperparameters in Appendix C.
We compare GIST with three state-of-the-art benchmarks:
  • random: We select samples from the dataset uniformly at random (and without replacement). While being a simple and lightweight approach, this implicitly encodes diversity in many settings and provides good solutions.
  • margin (Roth and Small, 2006): Margin sampling selects the top-$k$ points based on how hard they are to classify, i.e., using the margin scores $m_{i}=1-(\Pr(y=b\mid\mathbf{x}_{i};~{}\bm{\theta}_{0})-\Pr(y=b^{\prime}\mid\mathbf{x}_{i};~{}\bm{\theta}_{0}))$, which measure the difference between the probability of the best and second-best class label $b^{\prime}$ for an example.
  • $k$-center (Sener and Savarese, 2018): We run the classic greedy algorithm for $k$-center on the $\Delta$-nearest neighbor graph. The distance between non-adjacent nodes is implicitly assumed to be very large, which is acceptable in the large budget regime $k=\Omega(n)$.
For this task, GIST uses uncertainty values as its weights $w_{i}=u_{i}$ and the same graph as $k$-center, and sets $\alpha=0.85$.
Results.
We run each algorithm with cardinality constraint $k$ on the full dataset to get a subset of examples that we use to train a new ResNet-56 model. We report the top-1 classification accuracy of these models in Table 1.
$k$ (of data) random margin $k$-center GIST
30 65.02 67.64 67.32 67.70
40 68.92 70.45 70.98 71.20
50 70.86 73.41 72.51 73.67
60 72.39 74.70 74.78 75.16
70 72.92 75.57 75.62 75.85
80 74.36 75.39 76.26 76.49
90 75.10 76.11 76.02 76.50
Table 1: Top-1 classification accuracy (%) on ImageNet for different single-shot data downsampling algorithms. The cardinality constraint $k$ is expressed as a fraction of the ~1.3 million examples.
Conclusion
This work proposes a novel subset selection problem called MDDS that combines the total utility of the selected points with the $\texttt{div}(S)=\min_{u,v\in S:u\neq v}\textnormal{dist}(u,v)$ diversity objective. We design and analyze the GIST algorithm, which achieves a $\nicefrac{{2}}{{3}}$-approximation guarantee for general metric spaces by solving a series of maximum independent set instances on geometric intersection graphs with the GreedyIndependentSet bicriteria approximation algorithm. We also give a nearly linear-time implementation of GIST for points in low-dimensional Euclidean space. We complement our algorithm with a $(\nicefrac{{2}}{{3}}+\varepsilon)$-hardness of approximation, for any $\varepsilon>0$, in general metric spaces. We also prove hardness of approximation in the practical cases of Euclidean metrics or $k=\Omega(n)$. Finally, we present an empirical study showing that GIST steadily outperforms baseline methods for MDDS (in particular greedy) on a simple set of Gaussian points with random weights. We conclude with an experiment comparing the top-1 image classification accuracy achieved by GIST and three other state-of-the-art data sampling methods for ImageNet.
Appendix A Missing analysis for Section 3
A.1 Proof of Theorem 3.4
See 3.4
Proof.
First, observe that we can compute and store all pairwise distances in $O(n^{2}d)$ time and $O(n^{2})$ space. This is useful in the naive implementation because we can then look up the values $\textnormal{dist}(u,v)$ in constant time in any subroutine.
In the nearly linear-time implementation, we skip this precomputation step and use cover trees for fast exact nearest-neighbor search in low-dimensional Euclidean space (Beygelzimer et al., 2006): inserting a point into the index takes $O(c^{12}\log n)$ time and finding a nearest neighbor in the index to a query point takes $O(c^{6}\log n)$ time, where $c=\Theta(d)$ is the doubling dimension of $d$-dimensional Euclidean space.
Thus, we can speed up the following subroutines in Algorithm 1 to have near-linear running times in a fixed dimension:
  • Computing $d_{\min}$: $O(n^{2})\rightarrow O(n\cdot d^{12}\log n)$ after discarding duplicate points.
  • GreedyIndependentSet: $O(nk)\rightarrow O(n\cdot d^{12}\log n)$
  • Evaluating $f(S)$: $O(k^{2})\rightarrow O(k\cdot d^{12}\log k)$
Taking advantage of the error $\varepsilon$ in our $(\nicefrac{{2}}{{3}}-\varepsilon)$-approximation guarantee, we use the approximate point-set diameter algorithm in Imanparast et al. (2018, Theorem 1) to compute a $(1-\varepsilon)$-approximate diameter, together with two points realizing it. This speeds up the naive running time of Lines 34 in Algorithm 1 as follows:
  • Computing a $(1-\varepsilon)$-approximate $d_{\max}$: $O(n^{2})\rightarrow O\left(n\cdot 2^{d}d+\frac{(2\sqrt{d})^{2d}}{\varepsilon^{d-1}}\right)$
Since we have a smaller diameter $(1-\varepsilon)d_{\max}$, we need to show that it is compatible with the proof of Theorem 3.1. Concretely, this means modifying Eq. (5) and (6), and observing that the inequality
$\left(1-\frac{2}{3}+\frac{1}{3(1+\varepsilon)}\right)(1-\alpha)\cdot(1-\varepsilon)d^{*}\geq\left(\frac{2}{3}-\varepsilon\right)(1-\alpha)d^{*}$
still holds. Lastly, the number of calls to GreedyIndependentSet (i.e., distance thresholds that GIST tries) is $\min\{\lceil\log_{1+\varepsilon}(2/\varepsilon)\rceil,n^{2}\}$. Putting everything together proves the claim. ∎
Appendix B Missing analysis for Section 4
B.1 Proof of Theorem 4.2
See 4.2
Proof.
We use the vertex cover instance constructed in Håstad (2001, Theorem 8.1). For any $\varepsilon_{0}>0$, they construct a graph $G$ with $2^{r+2}$ nodes for some integer $r$ and show that it is NP-hard to distinguish between the following two cases:
  • YES case: Graph $G$ has an independent set of size $2^{r}(1-\varepsilon_{0})$.
  • NO case: Graph $G$ has no independent set of size larger than $2^{r}(\nicefrac{{1}}{{2}}+\varepsilon_{0})$.
We mimic the general structure of the proof of Theorem 4.1. Add a point for each vertex in graph $G$ and set its weight to $1$. Set the distance between two points to $1$ if they share an edge in the graph, and set it to $2$ otherwise. Lastly, let $k=2^{r}$. Since there are $2^{r+2}$ nodes in the graph, $k=|V|/4$ in this case. In the YES case, objective function (1) can achieve the following value by selecting the independent set of size $2^{r}(1-\varepsilon_{0})$:
OPT $\geq\alpha\cdot\frac{1}{2^{r}}\cdot(2^{r}(1-\varepsilon_{0}))+(1-\alpha)\cdot 2$
$=2-(1+\varepsilon_{0})\cdot\alpha.$
In the NO case, the largest value that can be achieved is the best of two scenarios: (a) selecting $k$ points with minimum distance $1$, or (b) selecting an independent set of size at most $2^{r}(\nicefrac{{1}}{{2}}+\varepsilon_{0})$ and achieving minimum distance $2$. This results in the following upper bound for the objective:
ALG $\leq\max\left\{\alpha+(1-\alpha)\cdot 1,\alpha\cdot(\nicefrac{{1}}{{2}}+\varepsilon_{0})+(1-\alpha)\cdot 2\right\}$
$=\max\left\{1,2-(\nicefrac{{3}}{{2}}-\varepsilon_{0})\cdot\alpha\right\}.$
The largest hardness gap occurs when the two terms of the $\max$ expression become equal, which happens at $\alpha=1/(\nicefrac{{3}}{{2}}-\varepsilon_{0})$ and yields the upper bound $\textnormal{ALG}\leq 1$. The optimum value OPT will be lower bounded by:
OPT $\geq 2-(1+\varepsilon_{0})\cdot\alpha$
$=2-\frac{1+\varepsilon_{0}}{\nicefrac{{3}}{{2}}-\varepsilon_{0}}$
$=\frac{2-3\varepsilon_{0}}{\nicefrac{{3}}{{2}}-\varepsilon_{0}}.$
The desired inapproximability gap of $\nicefrac{{3}}{{4}}+\varepsilon$ in Theorem 4.2 is achieved when the ratio between the upper bound on ALG and lower bound for OPT drops below it. In other words, we want to set $\varepsilon_{0}$ such that the following inequality holds:
$\frac{2-3\varepsilon_{0}}{\nicefrac{{3}}{{2}}-\varepsilon_{0}}>\frac{1}{\nicefrac{{3}}{{4}}+\varepsilon}.$
We set $\varepsilon_{0}\coloneqq\varepsilon$ and prove that the above inequality holds. We first note that both the numerator and denominator of the left side ratio are positive. Therefore, we can invert both ratios and swap their places to achieve the following equivalent inequality:
$\frac{3}{4}+\varepsilon>\frac{\nicefrac{{3}}{{2}}-\varepsilon_{0}}{2-3\varepsilon_{0}}\iff\varepsilon>\frac{\nicefrac{{3}}{{2}}-\varepsilon_{0}}{2-3\varepsilon_{0}}-\frac{3}{4}=\frac{6-4\varepsilon_{0}-6+9\varepsilon_{0}}{8-12\varepsilon_{0}}>\frac{5\varepsilon_{0}}{8}.$
The left-hand side inequality $\varepsilon>\nicefrac{{5\varepsilon_{0}}}{{8}}$ holds by our choice of $\varepsilon_{0}=\varepsilon$. ∎
B.2 Proof of Lemma 4.4
Let $G=(V,E)$ be a simple undirected graph. Our goal is to embed the vertices of $V$ into $\mathbb{R}^{d}$, for some $d\geq 1$, in a way that encodes the adjacency structure of $G$. Concretely, we want to construct a function $h_{G}:V\rightarrow\mathbb{R}^{d}$ such that:
  • $\lVert h_{G}(u)-h_{G}(v)\rVert_{2}=1$ if $\{u,v\}\not\in E$, and
  • $\lVert h_{G}(u)-h_{G}(v)\rVert_{2}\leq 1-\varepsilon_{G}$ if $\{u,v\}\in E$,
for the largest possible value of $\varepsilon_{G}\in(0,1]$.
Construction.
First, let $n=|V|$ and $m=|E|$. Augment $G$ by adding a self-loop to each node to get $G^{\prime}=(V,E^{\prime})$. We embed $V$ using $G^{\prime}$ since each node now has positive degree. Let $\deg^{\prime}(v)$ be the degree of $v$ in $G^{\prime}$ and $N^{\prime}(v)$ be the neighborhood of $v$ in $G^{\prime}$.
Define a total ordering on $E^{\prime}$ (e.g., lexicographically by sorted endpoints $\{u,v\}$). Each edge $e\in E^{\prime}$ corresponds to an index in the embedding dimension $d\coloneqq|E^{\prime}|=m+n$. We consider the embedding function that acts as a degree-normalized adjacency vector:
$h_{G}(v)_{e}=\begin{cases}\sqrt{\frac{1}{2\deg^{\prime}(v)}}&\text{if $v\in e$},\\ 0&\text{if $v\not\in e$}.\end{cases}$ (8)
See 4.4
Proof.
If $\{u,v\}\not\in E$, then we have
$\lVert h_{G}(u)-h_{G}(v)\rVert_{2}^{2}$ $=\sum_{e\in N^{\prime}(u)}\left(\sqrt{\frac{1}{2\deg^{\prime}(u)}}-0\right)^{2}+\sum_{e\in N^{\prime}(v)}\left(\sqrt{\frac{1}{2\deg^{\prime}(v)}}-0\right)^{2}$
$=\left(\frac{1}{2}\sum_{e\in N^{\prime}(u)}\frac{1}{\deg^{\prime}(u)}\right)+\left(\frac{1}{2}\sum_{e\in N^{\prime}(v)}\frac{1}{\deg^{\prime}(v)}\right)$
$=\frac{1}{2}+\frac{1}{2}$
$=1.$
This follows because the only index where both embeddings can be nonzero is $\{u,v\}$, if it exists.
Now suppose that $\{u,v\}\in E$. It follows that
$\lVert h_{G}(u)-h_{G}(v)\rVert_{2}^{2}$
$=\sum_{e\in N^{\prime}(u)\setminus\{v\}}\frac{1}{2\deg^{\prime}(u)}+\sum_{e\in N^{\prime}(v)\setminus\{u\}}\frac{1}{2\deg^{\prime}(v)}+\left(\sqrt{\frac{1}{2\deg^{\prime}(u)}}-\sqrt{\frac{1}{2\deg^{\prime}(v)}}\right)^{2}$
$=\frac{\deg^{\prime}(u)-1}{2\deg^{\prime}(u)}+\frac{\deg^{\prime}(v)-1}{2\deg^{\prime}(v)}+\left(\frac{1}{2\deg^{\prime}(u)}+\frac{1}{2\deg^{\prime}(v)}-2\sqrt{\frac{1}{4\deg^{\prime}(u)\deg^{\prime}(v)}}\right)$
$=\frac{1}{2}+\frac{1}{2}-\sqrt{\frac{1}{\deg^{\prime}(u)\deg^{\prime}(v)}}$
$\leq 1-\frac{1}{\Delta+1}.$
The previous inequality follows from $\deg^{\prime}(v)=\deg(v)+1\leq\Delta+1$. For any $x\in[0,1]$, we have
$\sqrt{1-x}\leq 1-\frac{x}{2},$
so it follows that
$\lVert h_{G}(u)-h_{G}(v)\rVert_{2}\leq\sqrt{1-\frac{1}{\Delta+1}}\leq 1-\frac{1}{2(\Delta+1)},$
which completes the proof. ∎
B.3 Proof of Theorem 4.5
See 4.5
Proof.
[…]
  • YES instance: Graph $G$ has an independent set of size $k$.
  • NO instance: Graph $G$ does not have an independent set of size greater than $(1-\varepsilon_{0})k$.
We reduce this instance of bounded-degree maximum independent set to MDDS with objective function (1) as follows. Embed each node of the graph $G$ into Euclidean space using the function $h_{G}(v)$ in Lemma 4.4. We use the same threshold value of $k$ (between YES and NO instances above) for the cardinality constraint on set $S$, and we set each weight $w(v)=1$. In a YES instance, selecting an independent set of size $k$ as the set $S$ results in the maximum value of objective (1):
$\textnormal{OPT}=\alpha\cdot\frac{1}{k}\cdot k+(1-\alpha)\cdot 1=1,$
since $\lVert h_{G}(u)-h_{G}(v)\rVert_{2}=1$ for any two distinct points $u,v\in S$ since there is no edge between $u$ and $v$ in graph $G$.
In a NO instance, the best objective value that can be achieved in polynomial-time is the maximum of the following two scenarios: (a) selecting $k$ points with minimum distance at most $81-1/(2(\Delta+1))=1-1/$ 8, or (b) selecting at most $(1-\varepsilon_{0})k$ vertices forming an independent set with minimum distance equal to $1$. The maximum value obtained by any polynomial-time algorithm is then
ALG $=\max\left\{\alpha(1-\varepsilon_{0})+(1-\alpha)\cdot 1,\alpha+(1-\alpha)\left(1-1/8\right)\right\}$
$=\max\left\{1-\varepsilon_{0}\cdot\alpha,(7+\alpha)/8\right\}.$
These two terms become equal for $\alpha=1/(1+8\varepsilon_{0})$. Therefore, the gap between the maximum value any algorithm can achieve in the NO case and the optimum value in the YES case is upper bounded by
$1-\varepsilon_{0}\cdot\alpha=1-\frac{\varepsilon_{0}}{\nicefrac{{1}}{{\varepsilon_{0}}}+8}=1-\varepsilon_{1}.$
Since $\varepsilon_{0}>0$ is a constant, $\varepsilon_{1}\coloneqq\varepsilon_{0}/(\nicefrac{{1}}{{\varepsilon_{0}}}+8)>0$ is also a positive constant. This completes the proof of APX-completeness. ∎
Appendix C Missing details for Section 5
Hyperparameters for ImageNet classification.
We generate predictions and embeddings for all points using a coarsely-trained ResNet-56 model (He et al., 2016) trained on a random 10% subset of ImageNet (Russakovsky et al., 2015). We use SGD with Nesterov momentum 0.9 with 450/90 epochs. The base learning rate is 0.1, and is reduced by a tenth at 5, 30, 69, and 80. We extract the penultimate layer features to produce $2048$-dimensional embeddings of each image.

### v2

@@
 Abstract
-We propose a novel subset selection task called min-distance diverse data summarization (MDDS), which has a wide variety of applications in machine learning, e.g., data sampling and feature selection.
-Given a set of points in a metric space, the goal is to maximize an objective that combines the total utility of the points and a diversity term that captures the minimum distance between any pair of selected points, subject to the constraint $|S|\leq k$.
-For example, the points may correspond to training examples in a data sampling problem, e.g., learned embeddings of images extracted from a deep neural network.
-This work presents the GIST algorithm, which achieves a $\nicefrac{{2}}{{3}}$-approximation guarantee for MDDS by approximating a series of maximum independent set problems with a bicriteria greedy algorithm.
-We also prove a complementary $(\nicefrac{{2}}{{3}}+\varepsilon)$-hardness of approximation, for any $\varepsilon>0$.
-Finally, we provide an empirical study that demonstrates GIST outperforms existing methods for MDDS on synthetic data, and also for a real-world image classification experiment the studies single-shot subset selection for ImageNet.
+We introduce a novel subset selection problem called min-distance diversification with monotone submodular utility (MDMS), which has a wide variety of applications in machine learning, e.g., data sampling and feature selection.
+Given a set of points in a metric space, the goal of MDMS is to maximize an objective function combining a monotone submodular utility term and a min-distance diversity term between any pair of selected points, subject to a cardinality constraint.
+We propose the GIST algorithm, which achieves a $\nicefrac{{1}}{{2}}$-approximation guarantee for MDMS by approximating a series of maximum independent set problems with a bicriteria greedy algorithm.
+We also prove that it is NP-hard to approximate to within a factor of $0.5584$.
+Finally, we demonstrate that GIST outperforms existing benchmarks for on a real-world image classification task that studies single-shot subset selection for ImageNet.
 1 Introduction
-Subset selection is a challenging optimization problem with a wide variety of applications in machine learning, including feature selection, recommender systems, news aggregation, drug discovery, data summarization, and designing pretraining sets for large language models (Anil et al., 2023).
-Data sampling in particular is a salient problem due to unprecedented and continuous data collection.
+Subset selection is a challenging combinatorial optimization problem with a wide variety of applications in machine learning, including feature selection, recommender systems, drug discovery, data summarization, and designing pretraining sets for large language models (Team et al., 2024).
+Data sampling in particular is a salient problem due to unprecedented and continuous streams of data collection.
 For example, LiDAR and imaging devices in one self-driving vehicle can easily capture ~80 terabytes of data per day (Kazhamiaka et al., 2021).
-In most subset selection tasks, we rely on the weight (or utility) of the objects to rank one over the other, and also to avoid selecting duplicate or near-duplicate objects.
-If we select a small subset, then we also want to ensure that the selected subset is a good representation of the original set.
-These utility, diversity, and coverage criteria can be expressed through objective functions, and the interesting research lies in developing efficient algorithms with strong approximation guarantees.
-The underlying machinery used in constrained subset selection algorithms shares many similarities with techniques from other areas of combinatorial optimization such as submodular maximization, $k$-center clustering, and convex hull approximations.
-In this work, we study the problem of selecting a set of points in a metric space that maximizes an objective that combines their utility and a minimum pairwise-distance diversity measure.
-Concretely, given a set of points $V$, weight function $w:V\rightarrow\mathbb{R}_{\geq 0}$, and distance function $\textnormal{dist}:V\times V\rightarrow\mathbb{R}_{\geq 0}$, our goal is to solve the cardinality-constrained optimization problem:
-$S^{*}$ $=\operatorname*{arg\,max}_{S\subseteq V}~{}\frac{1}{k}\sum_{v\in S}w(v)+\texttt{div}(S)$
-subject $\hskip 13.6572pt\text{subject to}~{}~{}|S|\leq k$
-where the diversity term $\texttt{div}(S)\coloneqq\min_{u,v\in S:u\neq v}\textnormal{dist}(u,v)$ encourages maximizing the minimum distance between all pairs of distinct points in the selected subset.
+In many subset selection tasks, we rely on the utility (or weight) of an object to rank one over the other, and also to avoid selecting duplicate or near-duplicate objects.
+If we select a small subset, then we must ensure that the selected set is a good representation of the original set.
+These utility, diversity, and coverage criteria are expressed through objective functions, and the interesting research lies in developing efficient algorithms with strong approximation guarantees.
+The algorithmic tools used in constrained subset selection shares many similarities with techniques from other areas of combinatorial optimization such as submodular maximization, $k$-center clustering, and convex hull approximations.
+This work introduces a novel subset selection problem called min-distance diversification with monotone submodular utility (MDMS).
+Given $n$ points $V$ in a metric space, a monotone submodular function $g:2^{V}\rightarrow\mathbb{R}_{\geq 0}$, and a cardinality constraint $k$, the goal is to solve
+$S^{*}$ $=\operatorname*{arg\,max}_{S\subseteq V}~{}g(S)+\texttt{div}(S)$ (1)
+subject $\hskip 13.6572pt\text{subject to}~{}~{}|S|\leq k,$
+where the diversity term $\texttt{div}(S)=\min_{u,v\in S:u\neq v}\textnormal{dist}(u,v)$ encourages maximizing the min distance between any two distinct points in the selected set.
+Meinl et al.
+(2011) call $\texttt{div}(S)$ the $p$-dispersion measure, and Bhaskara et al.
+(2016) call it the min-min diversity objective, where is known to admit a $\nicefrac{{1}}{{2}}$-approximation ratio subject to the strict cardinality constraint $|S|=k$.
+Other diversity terms such as $\sum_{u\in S}\min_{v\in S\setminus\{u\}}\textnormal{dist}(u,v)$ and $\sum_{u\in S}\sum_{v\in S}\textnormal{dist}(u,v)$ have also been studied (Meinl et al., 2011; Bhaskara et al., 2016; Borodin et al., 2017).
+While previous works have explored maximizing $\texttt{div}(S)$ in isolation, MDMS introduces a utility term and has not been studied before.
 Bhaskara et al.
-(2016) call $\texttt{div}(S)$ the min-min diversity objective and Meinl et al.
-(2011) call it $p$-dispersion maximization, where it is known to admit a $\nicefrac{{1}}{{2}}$-approximation ratio subject to the strict cardinality constraint $|S|=k$.
-Other diversity objectives such as the sum-of-min-distances $\sum_{u\in S}\min_{v\in S\setminus\{u\}}\textnormal{dist}(u,v)$ and sum-of-all-distances $\sum_{u\in S}\sum_{v\in S}\textnormal{dist}(u,v)$ have also been studied (Meinl et al., 2011; Bhaskara et al., 2016).
-While previous works have looked at maximizing the diversity objective $\texttt{div}(S)$, our formulation that introduces the utility term has not been studied before.
-In particular, Bhaskara et al.
-(2016) observe that min-min is a powerful diversity objective, but point out a major challenge: it is highly non-monotonic, i.e., it can exhibit a stark decrease in the objective value by adding a single new point.
-Our work shows that the use of a utility term along with the min-min can alleviate some of these pitfalls, and allows us to inherit the strong benefits of this strict diversity objective.
+(2016) remark that $\texttt{div}(S)$ is a powerful diversity objective, but point out a major challenge: it is highly non-monotonic, i.e., it can exhibit a stark decrease in the objective value by adding a single new point.
+Our work shows combining a monotone submodular utility term with $\texttt{div}(S)$ mitigates some of the challenges associated with this diversity measure, making it easier to leverage its properties.
 1.1 Our Contributions
 We summarize our main contributions below:
-  • We propose a novel subset selection problem called min-distance diverse data summarization (MDDS), where the goal is to maximize an objective function that combines a linear weight term with the min-min distance diversity term $\texttt{div}(S)$ subject to the cardinality constraint $|S|\leq k$.
-This problem has a wide variety of applications in machine learning, e.g., data denoising and active learning.
-  • We give a $\nicefrac{{2}}{{3}}$-approximation algorithm for MDDS called GIST that approximates a series of maximum independent set problems on geometric intersection graphs with a bicriteria greedy algorithm and returns the best solution.
-  • We prove a complementary $(\nicefrac{{2}}{{3}}+\varepsilon)$-hardness of approximation, for any $\varepsilon>0$, for general metric spaces by reducing from approximate maximum clique.
-We also prove that MDDS is APX-complete for Euclidean metrics.
-  • Our experiments show that GIST overcomes the shortcomings of existing methods for MDDS on synthetic data.
-We also demonstrate how GIST can be used to create better single-shot subsets of training data for an image classification benchmark on ImageNet (Russakovsky et al., 2015) compared to margin sampling and $k$-center algorithms.
-1.2 Background and related work
+  • In Section 2, we formalize the MDMS problem and present a simple $0.387$-approximation algorithm as a warm-up, demonstrating how the competing terms in the objective function interact.
+  • In Section 3, we present the GIST algorithm, which achieves a $\nicefrac{{1}}{{2}}$-approximation guarantee for MDMS by approximating a series of maximum independent set problems with a bicriteria greedy algorithm and returning the best solution.
+In the special case of linear utility functions $g(S)=\sum_{v\in S}w(v)$, we prove that GIST offers a stronger $\nicefrac{{2}}{{3}}$-approximation guarantee.
+[…]
+  • In Section 5, our experiments show that GIST outperforms baselines for MDMS on synthetic data (in particular the classic greedy algorithm).
+Then we show that GIST can be used to build better single-shot subsets of training data for an image classification benchmark on ImageNet compared to margin sampling and $k$-center algorithms, demonstrating the benefit of optimizing for a blend of utility and diversity.
+1.2 Related Work
 Diversity objectives.
-A closely related problem is $k$-center, which aims to find a subset of $k$ centers that minimize the largest distance of any point to its closest center in $S$.
+A closely related problem is $k$-center, which aims to find a subset of $k$ centers that minimize the largest distance of any point to its nearest center in $S$.
 Concretely, the objective is $\max_{v\in V}\min_{c\in S}\textnormal{dist}(v,c)$, and a greedy $2$-approximation algorithm is widely used in active learning applications (Sener and Savarese, 2018).
-The main difference between our objective function and $k$-center is the ability to benefit from the utility term, e.g., uncertainty values in an active learning setting.
-The second difference is that we relax the need to cover every point in the original dataset, which is particularly beneficial for large datasets, where we prefer to avoid instability due to certain corner or outlier points.
+The main difference between MDMS and $k$-center is the ability to benefit from the utility term, e.g., uncertainty values in an active learning setting.
+The second key difference is that we relax the requirement to cover every point in the dataset, which can be helpful for large datasets where we want to avoid instability due to certain corner or outlier points.
 Bhaskara et al.
 (2016) study various distance-based diversity metrics subject to a cardinality constraint.
-In particular, they give a $\nicefrac{{1}}{{8}}$-approximation algorithm for the sum-min diversity objective defined as $\sum_{u\in S}\min_{v\in S\setminus\{u\}}\textnormal{dist}(u,v)$ by rounding a linear programming relaxation, and they complement this result by proving a conditional $\nicefrac{{1}}{{2}}$-hardness of approximation under the planted clique assumption.
-They also study the min-min objective $\min_{u,v\in S:u\neq v}\textnormal{dist}(u,v)$, which is the same as $\texttt{div}(S)$.
-They do not have a linear weight function in their objective, so they enforce a strict cardinality constraint $|S|=k$; otherwise, it is always better to output at most one or two points.
+In particular, they give a $\nicefrac{{1}}{{8}}$-approximation algorithm for the sum-min diversity objective defined as $\sum_{u\in S}\min_{v\in S\setminus\{u\}}\textnormal{dist}(u,v)$ by rounding a linear programming relaxation.
+They complement this result by proving a conditional $\nicefrac{{1}}{{2}}$-hardness of approximation under the planted clique assumption.
+They also study the min-min objective $\min_{u,v\in S:u\neq v}\textnormal{dist}(u,v)$, but they enforce a strict cardinality constraint $|S|=k$; otherwise, it is optimal to only output one or two points.
 Meinl et al.
 (2011) study the same objective subject to $|S|=k$ under the name $p$-dispersion (also called max-min or remote-edge) diversity maximization.
-This cardinality issue does not arise in our formulation since we need to optimize the trade-off between competing linear weight (utility) and diversity terms.
+This exact cardinality issue does not arise in MDMS because we need to balance the trade-off between competing monotone submodular utility and diversity terms.
 Moumoulidou et al.
 (2020), Addanki et al.
-(2022), and Mahabadi and Trajanovski (2023) initiated the study of diversity maximization under fairness constraints.
-Given a partitioning of the dataset, they ensure that a prespecified number of points $k_{i}$ are selected from each group $i\in[m]$.
+(2022), and Mahabadi and Trajanovski (2023) study diversity maximization under fairness constraints.
+Given a partitioning of the dataset, they ensure that a specified number of points $k_{i}$ are selected from each group $i\in[m]$.
 This fairness constraint considerably differentiates their problem setting from ours (i.e., $|S|=k=\sum_{i=1}^{m}k_{i}$).
@@
 (2020); Addanki et al.
-(2022) maximize the $\texttt{div}(S)$ objective, and Mahabadi and Trajanovski (2023) consider the sum-of-pairwise distances $\sum_{u\in S}\sum_{v\in S}\textnormal{dist}(u,v)$ and sum-min objectives.
-None of these works, however, add a linear weight function to the objective.
+(2022) maximize $\texttt{div}(S)$ in this setting.
+Mahabadi and Trajanovski (2023) consider sum-of-all pairwise distances $\sum_{u\in S}\sum_{v\in S}\textnormal{dist}(u,v)$ and sum-min objectives.
+However, none of these works add a utility term to the objective.
+Borodin et al.
+(2017) optimize the combination of a submodular utility and sum-of-all pairwise distances in the selected set, making it the most similar work to MDMS.
+Unlike our $\texttt{div}(S)$ term, max-sum diversity is not very sensitive to duplicate points.
+In fact, adding $m$ copies of the same point to a set raises the increment of the diversity term by a factor of $m$ in their model.
+By using $\texttt{div}(S)$ instead, we can strictly avoid duplication across all selected points.
 Data sampling.
@@
 Citovsky et al.
-(2021) cluster examples which are represented by embeddings extracted from the penultimate layer of a partially trained DNN, and then use these clusters to diversify uncertain examples in each iteration.
+(2021) cluster examples that are represented by embeddings extracted from the penultimate layer of a partially trained DNN, and use these clusters to diversify uncertain examples in each iteration.
 Our work differs from these and several others (e.g., Kirsch et al.
-(2019); Zhdanov (2019)) in that we directly incorporate both the utility and diversity terms in our objective function.
+(2019); Zhdanov (2019)) in that we directly incorporate the utility and diversity terms in our objective function.
 2 Preliminaries
-We are given a set of $n$ points $V\subseteq M$ in the metric space $(M,\textnormal{dist})$, where $\textnormal{dist}:M\times M\rightarrow\mathbb{R}_{\geq 0}$ is the distance function.
-We are also given weights $w:V\rightarrow\mathbb{R}_{\geq 0}$ representing the utility of each point.
-We extend these functions to take subsets $S\subseteq V$ as input in the standard way: $w(S)=\sum_{v\in S}w(v)$ and $\textnormal{dist}(u,S)=\min_{v\in S}\textnormal{dist}(u,v)$.
+Here we formalize the MDMS problem and present a simple $0.387$-approximation algorithm as a starting point.
+Submodular set function.
+For a set function $g:2^{V}\rightarrow\mathbb{R}$ and any $S,T\subseteq V$, let $g(T\mid S)=g(S\cup T)-g(T)$ denote the marginal gain of $g$ at $T$ with respect to $S$.
+A function $g$ is submodular if for every $S\subseteq T\subseteq V$ and $v\in V\setminus T$, we have $g(v\mid S)\geq g(v\mid T)$, where we overload the marginal gain notation for singleton sets.
+A submodular function $g$ is monotone if for every $S\subseteq T\subseteq V$, $g(S)\leq g(T)$.
+[…]
+Min-distance diversity.
+For a set of $n$ points $V$ in a metric space, define the min-distance diversity function
+$\texttt{div}(S)$ =$=\begin{cases}\min_{u,v\in S:u\neq v}\textnormal{dist}(u,v)&\text{if }|S|\geq 2,\\ \max_{u,v\in V}\textnormal{dist}(u,v)&\text{if }|S|\leq 1.\end{cases}$
+For $\texttt{div}(S)$ to be well-defined, we set the diversity to be the diameter of $V$ if $|S|\leq 1$, i.e., $d_{\max}=\max_{u,v\in V}\textnormal{dist}(u,v)$.
+This implies that it is always beneficial to select at least two points to maximize the overall objective.
+We extend the distance function to take subsets $S\subseteq V$ as input, i.e., $\textnormal{dist}(u,S)=\min_{v\in S}\textnormal{dist}(u,v)$.
 For the sake of completeness, we define $\textnormal{dist}(u,\varnothing)=\infty$.
-Intersection graph.
-Let $G_{d}(V)$ denote the intersection graph of $V$ for distance threshold $d$, i.e., the graph with node set $V$ and edge set
-$E=\{(u,v)\in V^{2}:\textnormal{dist}(u,v)<d\}.$
-This graph is well-defined for any metric space and distance $d\in\mathbb{R}$.
-Note that for any independent set $S$ of $G_{d}(V)$ and pair of distinct nodes $u,v\in S$, we have $\textnormal{dist}(u,v)\geq d$.
-Definition 2.1.
-This work focuses on the min-distance diversity reward function
-$\texttt{div}(S)$ =$=\begin{cases}\min_{u,v\in S:u\neq v}\textnormal{dist}(u,v)&\text{if }|S|\geq 2,\\ \max_{u,v\in V}\textnormal{dist}(u,v)&\text{if }|S|\leq 1.\end{cases}$
-This reward function encourages selecting a well-separated set of points (generalizing the notion of data deduplication).
-To make $\texttt{div}(S)$ well-defined if $S$ is empty or consists of a singleton point, we set the diversity to be the diameter of the whole ground set $V$, implying that it is always beneficial to select at least two points to maximize the value of the selected set.
-For any $\alpha\in[0,1]$, we define the objective function
-$f(S)=\alpha\frac{1}{k}\sum_{v\in S}w(v)+(1-\alpha)\texttt{div}(S),$ (1)
-which allows us to control the trade-off between the average utility and the diversity of set $S$.
-The goal of the min-distance diverse data summarization problem (MDDS) is to maximize $f(S)$ subject to a cardinality constraint $k$:
+Optimization problem.
+Define the objective function
+$f(S)=g(S)+\texttt{div}(S).$ (2)
+The goal of the min-distance diversification with monotone submodular utility (MDMS) problem is to maximize $f(S)$ subject to a cardinality constraint $k$:
 $S^{*}=\operatorname*{arg\,max}_{S\subseteq V:|S|\leq k}f(S).$
 Let $\textnormal{OPT}=f(S^{*})$ denote the optimal objective value.
-Simple $\nicefrac{{1}}{{2}}$-approximation algorithm.
-Equation 1 has two parts, each of which is easy to optimize in isolation.
-To maximize the $w(S)$ term, we simply select the $k$ heaviest points.
-In contrast, to maximize the $\texttt{div}(S)$ term, we select two points whose distance is the diameter of $V$.
-Outputting a subset that corresponds to the maximum of these two terms gives a $\nicefrac{{1}}{{2}}$-approximation guarantee.
-To see this, observe that
-$\textnormal{ALG}_{\textnormal{simple}}$ $=\max\left\{\alpha\frac{1}{k}\max_{|S|\leq k}w(S),(1-\alpha)\max_{u,v\in V}\textnormal{dist}(u,v)\right\}$
-$\geq\frac{1}{2}\cdot\alpha\frac{1}{k}\max_{|S|\leq k}w(S)+\frac{1}{2}\cdot(1-\alpha)\max_{u,v\in V}\textnormal{dist}(u,v)$
-$\geq\frac{1}{2}\cdot\left(\alpha\frac{1}{k}w(S^{*})+(1-\alpha)\texttt{div}(S^{*})\right)=\frac{1}{2}\cdot\textnormal{OPT},$
-where $\texttt{div}(S)$ is defined in Definition 2.1. Our main goals are to improve over this simple baseline and prove complementary hardness results.
+Remark 2.1.
+The objective function $f(S)$ is not submodular. We present a counterexample in Appendix A.
+Intersection graph.
+Let $G_{d}(V)$ be the intersection graph of $V$ for distance threshold $d$, i.e., with nodes $V$ and edges
+$E=\{(u,v)\in V^{2}:\textnormal{dist}(u,v)<d\}.$
+Note that for any independent set $S$ of $G_{d}(V)$ and pair of distinct nodes $u,v\in S$, we have $\textnormal{dist}(u,v)\geq d$.
+2.1 Warm-up: Simple $0.387$-Approximation Algorithm
+The objective function $f(S)$ has two terms, each of which is easy to (approximately) optimize in isolation. For the $g(S)$ term, run the greedy algorithm to get $S_{1}$ satisfying
+$g(S_{1})$ $\geq(1-\nicefrac{{1}}{{e}})\cdot\max_{|S|\leq k}g(S)$
+$\geq(1-\nicefrac{{1}}{{e}})\cdot g(S^{*}).$
+For the $\texttt{div}(S)$ term, let $S_{2}$ be a pair of diametrical points in $V$.
+The algorithm then flips a biased coin and outputs $S_{1}$ with probability $p\approx 0.613$, and $S_{2}$ otherwise.
+This achieves a $0.387$-approximation guarantee:
+$\mathbb{E}[\textnormal{ALG}_{\textnormal{simple}}]$ $\geq p\cdot g(S_{1})+(1-p)\cdot\texttt{div}(S_{2})$
+$\geq p\cdot(1-\nicefrac{{1}}{{e}})\cdot g(S^{*})+(1-p)\cdot\texttt{div}(S^{*})$
+$=\frac{e-1}{2e-1}\cdot\textnormal{OPT},$
+where we take $p=e/(2e-1)$ by solving $p\cdot(1-\nicefrac{{1}}{{e}})=1-p$. One of the main goals of this paper is to improve over this baseline and prove complementary hardness results.
 3 Algorithm
-We now present the GIST algorithm, which achieves a $(\nicefrac{{2}}{{3}}-\varepsilon)$-approximation for the MDDS problem.
-It works by repeatedly calling the GreedyIndependentSet heuristic to find an independent set of the intersection graph $G_{d}(V)$ for various distance thresholds $d$.
-The subroutine GreedyIndependentSet sorts the points by their weight in a non-increasing order breaking ties arbitrarily.
-It starts with an empty set of selected points, and then it scans the points according to the sorted order.
-For each point, GreedyIndependentSet selects the point as long as its distance to the set of selected points so far is at least $d$ and the number of selected points does not exceed $k$.
+In this section, we present the GIST algorithm and prove that it achieves an approximation ratio of $(\nicefrac{{1}}{{2}}-\varepsilon)$ for MDMS.
+GIST works by sweeping over multiple distance thresholds $d$ and calling a GreedyIndependentSet subroutine to find a high-value independent set of the intersection graph $G_{d}(V)$.
+This is in contrast to the simple $0.387$-approximation algorithm that just tries two extreme thresholds $D=\{0,d_{\max}\}$.
+The GreedyIndependentSet heuristic builds a set $S$ (starting from the empty set) by repeatedly adding $v\in V\setminus S$ with the highest marginal gain with respect to the submodular function $g$ while satisfying $\textnormal{dist}(v,S)\geq d$.
+This subroutine runs until either $|S|=k$ or $S$ is a maximal independent set of $G_{d}(V)$.
 […]
-Algorithm 1 Diversified data summarization using greedy-weighted maximum independent sets.
-1:function GIST(points $V$, weights $w:V\rightarrow\mathbb{R}_{\geq 0}$, budget $k$, error $\varepsilon$)
-2:     Initialize $S\leftarrow\textnormal{{GreedyIndependentSet}}(V,w,0,k)$triangleright $k$ heaviest points
+Algorithm 1 Diverse data summarization for submodular utilities using greedy-weighted maximal independent sets.
+1:function GIST(points $V$, monotone submodular function $g:V\rightarrow\mathbb{R}_{\geq 0}$, budget $k$, error $\varepsilon$)
+2:     Initialize $S\leftarrow\textnormal{{GreedyIndependentSet}}(V,g,0,k)$triangleright runs the classic greedy algorithm
 3:     Let $d_{\max}=\max_{u,v\in V}\textnormal{dist}(u,v)$ be the diameter of $V$
@@
 8:     for threshold $d\in D$ do
-9:         Set $T\leftarrow\textnormal{{GreedyIndependentSet}}(V,w,d,k)$
+9:         Set $T\leftarrow\textnormal{{GreedyIndependentSet}}(V,g,d,k)$
 10:         if $f(T)\geq f(S)$ then
 11:              Update $S\leftarrow T$               return $S$
-1:function GreedyIndependentSet(points $V$, weights $w:V\rightarrow\mathbb{R}_{\geq 0}$, distance $d$, budget $k$)
+1:function GreedyIndependentSet(points $V$, monotone submodular function $g:V\rightarrow\mathbb{R}_{\geq 0}$, distance $d$, budget $k$)
 2:     Initialize $S\leftarrow\varnothing$
-3:     for $v\in V$ sorted by non-increasing weight do
-4:         if $\textnormal{dist}(v,S)\geq d$ and $|S|<k$ then
-5:              Update $S\leftarrow S\cup\{v\}$               return $S$
+3:     for $i=1$ to $k$ do
+4:         Let $C\leftarrow\{v\in V\setminus S:\textnormal{dist}(v,S)\geq d\}$
+5:         if $C=\varnothing$ then
+6:              return $S$triangleright $S$ is a maximal independent set of $G_{d}(V)$
+7:         Find $t\leftarrow\operatorname*{arg\,max}_{v\in C}g(v\mid S)$
+8:         Update $S\leftarrow S\cup\{t\}$      return $S$
 Theorem 3.1.
-For any $\varepsilon>0$, GIST outputs a set $S\subseteq V$ with $|S|\leq k$ and $f(S)\geq(\nicefrac{{2}}{{3}}-\varepsilon)\cdot\textnormal{OPT}$.
-The main building block in our design and analysis of GIST is the following lemma, which is inspired by the greedy $2$-approximation algorithm for metric $k$-centers.
+For any $\varepsilon>0$, the GIST algorithm outputs a set $S\subseteq V$ with $|S|\leq k$ and $f(S)\geq(\nicefrac{{1}}{{2}}-\varepsilon)\cdot\textnormal{OPT}$.
+The main building block in our design and analysis of GIST is the following lemma, which is inspired by the greedy $2$-approximation algorithm for metric $k$-centers and adapted for monotone submodular rewards.
 Lemma 3.2.
+Let $S_{d}^{*}$ be a maximum-value set of size at most $k$ with diversity at least $d$. In other words,
+$S_{d}^{*}=\operatorname*{arg\,max}_{S:|S|\leq k,\texttt{div}(S)\geq d}g(S).$
+[…]
+Proof.
+Let $k^{\prime}=|T|$ and $t_{1},t_{2},\dots,t_{k^{\prime}}$ be the points in $T$ in selection order.
+Let $B_{i}=\{v\in V:\textnormal{dist}(t_{i},v)<d^{\prime}\}$ be the points in $V$ in the radius-$d^{\prime}$ open ball around $t_{i}$.
+Since the distance between any pair of points in $B_{i}$ is at most $2d^{\prime}<d$, each set $B_{i}$ contains at most one point in $S_{d}^{*}$.
+We first construct an injective map $h:S_{d}^{*}\rightarrow T$.
+We say that a point $v\in V$ is covered by $B_{i}$ if $v\in B_{i}$.
+For each covered $s\in S_{d}^{*}$, we map it to $t_{i}$, where $i$ is the minimum index for which $B_{i}$ covers $s$.
+If there is an $s\in S_{d}^{*}$ not covered by any set $B_{i}$, then GreedyIndependentSet must have selected $k$ points, i.e., $|T|=k$.
+We map the uncovered points in $S_{d}^{*}$ to arbitrary points in $T$ while preserving the injective property.
+Since $g$ is a monotone submodular function, we have
+$g(S_{d}^{*})-g(T)\leq\sum_{s\in S_{d}^{*}}g(s\mid T).$
+We use $h$ to account for $g(s\mid T)$ in terms of the values we gained in set $T$.
+For any $s\in S_{d}^{*}$, let $T_{s}$ be the set of points added to set $T$ in algorithm GreedyIndependentSet right before the addition of $h(s)$.
+Since GreedyIndependentSet incrementally adds points and never deletes points from $T$, we know $T_{s}\subseteq T$.
+By submodularity, $g(s\mid T)\leq g(s\mid T_{s})$.
+By the greedy nature of the algorithm, we also know that
+$g(h(s)\mid T_{s})\geq g(s\mid T_{s}).$
+We note that the sum of the former terms, $g(h(s)\mid T_{s})$, is at most $g(T)$ since $h$ is injective and $g$ is nondecreasing. Thus, we conclude that
+$g(S_{d}^{*})-g(T)$ $\leq\sum_{s\in S_{d}^{*}}g(s\mid T)$
+$\leq\sum_{s\in S_{d}^{*}}g(h(s)\mid T_{s})$
+$\leq g(T),$
+which completes the proof. ∎
+Equipped with this bicriteria approximation guarantee, we can now analyze the approximation ratio of GIST.
+Proof of Theorem 3.1.
+Let $d^{*}$ be the minimum pairwise distance between points in $S^{*}$.
+The GIST algorithm iterates over a set of thresholds $D$.
+The definition of $D$ implies that at least one threshold $d$ is in the range $[d^{*}/(2(1+\varepsilon)),d^{*}/2)$ unless $d^{*}\leq\varepsilon d_{\max}$.
+We deal with this special case later and focus on the case when such a $d$ exists.
+Let $T$ be the output of GreedyIndependentSet for threshold value $d$. Applying Lemma 3.2 since $d<d^{*}/2$, we know that
+$g(T)\geq\frac{1}{2}\cdot g(S_{d^{*}}^{*})\geq\frac{1}{2}\cdot g(S^{*}).$
+We also know from our choice of $d$ that
+$\texttt{div}(T)\geq\frac{1}{2(1+\varepsilon)}\cdot\texttt{div}(S^{*}).$
+Combining these inequalities gives us
+$f(T)\geq\frac{1}{2(1+\varepsilon)}\cdot\textnormal{OPT}>\left(\frac{1}{2}-\varepsilon\right)\cdot\textnormal{OPT}.$
+It remains to prove the claim for the special case of $d^{*}\leq\varepsilon d_{\max}$.
+Algorithm GIST considers two initial feasible sets and picks the better of the two as the initial value for $T$.
+The first set is the classic greedy solution for the submodular function $g$, ignoring the diversity term.
+This gives the lower bound (Nemhauser et al., 1978):
+$f(T)\geq\left(1-\frac{1}{e}\right)\cdot g(S^{*})\geq\left(1-\frac{1}{e}\right)\cdot\left(\textnormal{OPT}-\varepsilon d_{\max}\right).$
+The other initial set realizes the max pairwise distance $d_{\max}$ and ignores the submodular term, yielding the lower bound:
+$f(T)\geq d_{\max}.$
+Combining these two lower bounds, we get a final bound of
+$f(T)\geq\left(1-\frac{1}{e}\right)\cdot\frac{\textnormal{OPT}}{1+\varepsilon},$
+which completes the proof. ∎
+Remark 3.3.
+We can modify GIST to guarantee an exact $\nicefrac{{1}}{{2}}$-approximation ratio by trying all pairwise thresholds $D^{\prime}=\{\textnormal{dist}(u,v)/2-\varepsilon^{\prime}:u,v\in V\}$, for a small enough $\varepsilon^{\prime}>0$, at the cost of $O(n^{2})$ calls to GreedyIndependentSet instead of $O(\log_{1+\varepsilon}(1/\varepsilon))$.
+3.1 Linear Utility
+We now assume $g(S)=\sum_{v\in S}w(v)$ is a linear function for some weights $w:V\rightarrow\mathbb{R}_{\geq 0}$.
+This is a special case of submodularity, so Theorem 3.1 gives a $(\nicefrac{{1}}{{2}}-\varepsilon)$-approximation.
+However, GIST achieves a stronger approximation ratio in this case.
+Theorem 3.4.
+Let $g(S)=\sum_{v\in S}w(v)$ be a linear function. For any $\varepsilon>0$, GIST outputs a set $S\subseteq V$ with $|S|\leq k$ and $f(S)\geq(\nicefrac{{2}}{{3}}-\varepsilon)\cdot\textnormal{OPT}$.
+We defer the proof to Appendix B, but describe the main differences.
+For linear functions, Lemma 3.2 can be strengthened to show that GreedyIndependentSet outputs a set $T$ such that $g(T)\geq g(S_{d}^{*})$, for any $d^{\prime}<d/2$.
+Then, for the right distance threshold $d\in D$, we get two lower bounds $\textnormal{ALG}\geq\max\left\{g(S^{*})+d^{*}/(2(1+\varepsilon)),d^{*}\right\}$ that can be balanced appropriately.
+Remark 3.5.
+GIST iterates over $O(\log_{1+\varepsilon}(1/\varepsilon))$ distance thresholds.
+For each $d$, it inserts up to $k$ points with $O(n)$ distance and submodular valuation oracle calls per insertion, so these iterations consist of $O(nk\log_{1+\varepsilon}(1/\varepsilon))$ distance and submodular oracle calls.
+For linear functions, we can sort the points in advance by $w(v)$ and iterate over this ordering in all GreedyIndependentSet calls to reduce the number of valuation probes to $O(n\log_{1+\varepsilon}(1/\varepsilon))$.
+4 Hardness of Approximation
+[…]
+Submodular utility.
+There is no polynomial-time $0.5584$-approximation algorithm if $g$ is submodular.
+Linear utility.
+  • There is no polynomial-time $(\nicefrac{{2}}{{3}}+\varepsilon)$-approximation algorithm if $g$ is linear for general metrics, for any $\varepsilon>0$.
+  • APX-completeness for linear utility functions even in the Euclidean metric, i.e., there is no polynomial-time approximation scheme (PTAS) for this problem.
+4.1 General Metric Spaces
+We build on the set cover hardness instances originally designed by Feige (1998); Feige et al.
+(2004); Feige and Vondrák (2010), and further engineered by Kapralov et al.
+(2013).
+In particular, we present the instances designed in Kapralov et al.
+(2013, Section 2) as follows:
+Hardness of max $k$-cover.
+For any constants $c,\varepsilon>0$ and a given collection/family of sets $\mathcal{F}$ partitioned into groups $\mathcal{F}_{1},\mathcal{F}_{2},\cdots,\mathcal{F}_{k}$, it is NP-hard to distinguish between the following two cases:
+  • YES case: there exists $k$ disjoint sets $S_{i}$, one from each $\mathcal{F}_{i}$, whose union covers the entire universe.
+  • NO case: for any $\ell\leq c\cdot k$ sets, their union covers at most a $(1-(1-\nicefrac{{1}}{{k}})^{\ell}+\varepsilon)$-fraction of the universe.
+The construction above can be done such that each set $S_{i}\in\mathcal{F}=\bigcup_{r=1}^{k}\mathcal{F}_{r}$ has the same size.
+We also note that the parameter $k$ should be larger than any constant we set since there is always an exhaustive search algorithm with running time $|\mathcal{F}|^{k}$ to distinguish between the YES and NO cases.
+Therefore, we can assume, e.g., that $k\geq 1/\varepsilon$.
+We use this max $k$-cover instance to prove hardness of approximation for the MDMS problem.
+Theorem 4.1.
+For any $\delta>0$, it is NP-hard to approximate MDMS within a factor of $\frac{2(1-1/e)}{2(1-1/e)+1}+\delta<0.55836+\delta$.
+Proof.
+We construct our instance based on the collection of sets above as follows.
+We set $\varepsilon=\min\{\delta,\delta^{2}/6\}$ and $c=1$.
+As mentioned above, we can assume that $k>\nicefrac{{1}}{{\varepsilon}}+1$ since there is always a brute force polynomial time algorithm for constant $k$ to distinguish between YES and NO instances.
+Suppose there are $n$ sets $S_{1},S_{2},\cdots,S_{n}$ in the collection $\mathcal{F}$.
+We represent these $n$ sets with $n$ points in the MDMS instance.
+We overload the notation to show the corresponding point with $S_{i}$ too.
+For any subset of sets/points $T$, the submodular value $g(T)$ is defined as the cardinality of the union of corresponding sets, i.e., the coverage submodular function.
+In other words, $g(T)=|\bigcup_{S_{i}\in T}S_{i}|$.
+The distances between points/sets are either $d$ or $2d$.
+If two sets are disjoint and belong to two separate partitions $\mathcal{F}_{r}$ and $\mathcal{F}_{r^{\prime}}$, their distance is $2d$.
+Otherwise, we set their distance to $d$.
+So for any two sets $S,S^{\prime}$ belonging to the same group $\mathcal{F}_{r}$, their distance is set to $d$.
+Also, for any two sets with nonempty intersection, we set their distance to $d$ too.
+Any other pair of points will have distance $2d$.
+We set $d$ to be $(1-\nicefrac{{1}}{{e}})U$ where $U$ is the cardinality of the union of all sets, i.e., $U=|\bigcup_{S\in\mathcal{F}}S|$.
+In the YES case, the optimum solution is the family of $k$ disjoint sets from different groups that cover the entire universe.
+The value of optimum in this case is $2d+U=(2(1-\nicefrac{{1}}{{e}})+1)U$.
+The algorithm has two possibilities: (case a) the algorithm gives up on the diversity objective and selects $k$ sets with minimum distance $d$, or (case b) it aims for a diversity term of $2d$.
+In case (a), the submodular value is at most $(1-(1-\nicefrac{{1}}{{k}})^{k}+\varepsilon)U$ because of the property of the NO case.
+We note that if the algorithm finds a set of $k$ sets with union size above this threshold, we can conclude that we have a YES instance, which contradicts the hardness result.
+The limit of the upper bound for the submodular value as $k$ goes to infinity is $1-\nicefrac{{1}}{{e}}+\varepsilon$. We use its expansion series to derive:
+$\left(1-\frac{1}{k}\right)^{k}\geq\frac{1}{e}-\frac{1}{2ek}-\frac{5}{24ek^{2}}-\cdots$
+The sequence of negative terms declines in absolute value with a rate of at least $\nicefrac{{1}}{{k}}$.
+Thus, the absolute value of their total sum is at most the first deductive term $\nicefrac{{1}}{{2ek}}$ times $1/(1-\nicefrac{{1}}{{k}})=k/(k-1)$, which gives the simpler bound:
+$\left(1-\frac{1}{k}\right)^{k}\geq\frac{1}{e}-\frac{1}{2e(k-1)}.$
+Since $k-1$ is at least $\nicefrac{{1}}{{\varepsilon}}$, the submodular value in case (a) does not exceed:
+$\left(1-\frac{1}{e}+\frac{1}{2e(k-1)}+\varepsilon\right)U\leq\left(1-\frac{1}{e}+2\varepsilon\right)U.$
+Recall that $d=(1-\nicefrac{{1}}{{e}})U$, so the ratio of what the algorithm achieves in case (a) and the optimum solution of YES case is at most:
+$\frac{(1-\nicefrac{{1}}{{e}}+2\varepsilon)U+d}{U+2d}$ $=\frac{2(1-\nicefrac{{1}}{{e}})+2\varepsilon}{2(1-\nicefrac{{1}}{{e}})+1}$
+$\leq\frac{2(1-\nicefrac{{1}}{{e}})}{2(1-\nicefrac{{1}}{{e}})+1}+\delta.$
+In case (b), the algorithm is forced to pick only disjoint sets to maintain a minimum distance of $2d$.
+This means if the algorithm picks $\ell$ sets, their union has size $\ell\cdot\nicefrac{{U}}{{k}}$.
+This is true because all sets in $\mathcal{F}$ have the same size and we know in the YES case, the union of $k$ disjoint sets covers the entire universe hence each set has size $\nicefrac{{U}}{{k}}$.
+For the special case of $\ell=1$, we know that only a $\nicefrac{{1}}{{k}}<\varepsilon$ fraction of universe is covered.
+We upper bound the covered fraction in terms of $\varepsilon$ for the other cases.
+The property of the NO case implies the following upper bound on $\ell$:
+$\frac{\ell}{k}\leq 1-\left(1-\frac{1}{k}\right)^{\ell}+\varepsilon.$
+We use the binomial expansion of $(1-\nicefrac{{1}}{{k}})^{\ell}$ and note that each negative term exceeds its following positive term in absolute value.
+This is true because of the cardinality constraint $\ell\leq k$.
+Therefore,
+$\left(1-\frac{1}{k}\right)^{\ell}$ $\geq 1-\frac{\ell}{k}+\frac{\ell(\ell-1)}{2k^{2}}-\frac{\ell(\ell-1)(\ell-2)}{6k^{3}}$
+$\geq 1-\frac{\ell}{k}+\frac{\ell(\ell-1)}{3k^{2}}$
+$\geq 1-\frac{\ell}{k}+\frac{\ell^{2}}{6k^{2}},$
+where the second to last inequality holds since $\ell-2<k$ and the last inequality holds because $\ell\geq 2$ and consequently $\ell-1\geq\nicefrac{{\ell}}{{2}}$.
+We can now revise the initial inequality:
+$\frac{\ell}{k}$ $\leq(1-(1-\nicefrac{{1}}{{k}})^{\ell}+\varepsilon)$
+$\leq 1-1+\frac{\ell}{k}-\frac{\ell^{2}}{6k^{2}}+\varepsilon\implies\frac{\ell^{2}}{6k^{2}}\leq\varepsilon.$
+Thus, the fraction of the universe that the algorithm covers, namely $\nicefrac{{\ell}}{{k}}$, is at most $\sqrt{6\varepsilon}\leq\delta$.
+In case (b), the ratio of what the algorithm achieves, and the optimum solution of YES case is at most:
+$\frac{\delta\cdot U+2d}{U+2d}=\frac{2(1-\nicefrac{{1}}{{\mathrm{e}}})+\delta}{2(1-\nicefrac{{1}}{{\mathrm{e}}})+1}\leq\frac{2(1-\nicefrac{{1}}{{\mathrm{e}}})}{2(1-\nicefrac{{1}}{{\mathrm{e}}})+1}+\delta.$
+This concludes the proof in both cases (a) and (b). ∎
+4.2 Linear Utility
+For the special case of linear utility functions, we prove a tight hardness-of-approximation result complementing the $\nicefrac{{2}}{{3}}$-approximation guarantee in Theorem 3.4.
+Theorem 4.2.
+[…]
+We give the proof in Section C.1.
+Our construction builds on the work of Håstad (1996); Zuckerman (2006), which shows that the maximum clique problem does not admit an efficient $n^{1-\theta}$-approximation algorithm, for any constant $\theta>0$.
+4.3 Euclidean Metric
+[…]
+Lemma 4.3 (Alimonti and Kann (2000, Theorem 3.2)).
+The maximum independent set problem for graphs with degree at most $3$ is APX-complete.
+Our second ingredient is a node embedding function $h_{G}(v)$ that encodes graph adjacency in Euclidean space.
+Lemma 4.4.
+Let $G=(V,E)$ be a simple undirected graph with $n=|V|$, $m=|E|$, and max degree $\Delta$. There exists an embedding $h_{G}:V\rightarrow\mathbb{R}^{n+m}$ such that if $\{u,v\}\in E$ then
+$\lVert h_{G}(u)-h_{G}(v)\rVert_{2}\leq 1-\frac{1}{2(\Delta+1)},$
+and if $\{u,v\}\not\in E$ then $\lVert h_{G}(u)-h_{G}(v)\rVert_{2}=1$.
+Theorem 4.5.
+The MDMS problem is APX-complete for the Euclidean metric even if $g$ is a linear function.
+We sketch the main idea below and defer complete proofs to Appendix C.
+Let $\mathcal{I}(G)$ be the set of independent sets of $G$.
+Using Lemma 4.4, for any set of nodes $S\subseteq V$ with $|S|\geq 2$ in a graph $G$ with max degree $\Delta\leq 3$, we have the property:
+  • $S\in\mathcal{I}(G)\implies\texttt{div}(S)=1$;
+  • $S\not\in\mathcal{I}(G)\implies\texttt{div}(S)\leq 1-\frac{1}{2(\Delta+1)}\leq 1-\nicefrac{{1}}{{8}}$.
+The gap between the cases is at least $\nicefrac{{1}}{{8}}$ for all such graphs (i.e., a universal constant), allowing us to prove that MDMS inherits the APX-completeness of bounded-degree max independent set.
+5 Experiments
+5.1 Synthetic Dataset
+We use a collection of weighted random points and a budget-additive (submodular) utility function to (1) compare against baselines and (2) highlight the non-monotone behavior.
+Setup.
+Generate $n=1000$ points $\mathbf{x}_{i}\in\mathbb{R}^{d}$, for $d=64$, where each $\mathbf{x}_{i}\sim\mathcal{N}(\mathbf{0},\mathbf{I}_{d})$ is i.i.d.
+and assigned a uniform continuous weight $w_{i}\sim\mathcal{U}_{[0,1]}$.
+For $\alpha,\beta\in[0,1]$, the objective function below trades off between the average (capped) utility of the selected points and their min-distance diversity:
+$f(S)=\alpha\cdot\min\left\{\frac{1}{k}\sum_{i\in S}w_{i},\beta\right\}+(1-\alpha)\cdot\texttt{div}(S).$
+We consider three baselines: random, simple, and greedy.
+random chooses a random set of $k$ points, permutes them, and returns the best prefix since the objective function is non-monotone.
+simple is the $0.387$-approximation algorithm in Section 2.1.
+greedy builds $S$ one point at a time by selecting the point with index $i^{*}_{t}=\operatorname*{arg\,max}_{i\in V\setminus S_{t-1}}f(S_{t-1}\cup\{i\})-f(S_{t-1})$ at each step $t\in[k]$, and outputs the best prefix $S=\operatorname*{arg\,max}_{t\in[k]}f(S_{t})$.
+Results.
+We plot the values of $f(S)$ for the baseline methods and GIST for all $k\in[n]$ in Figure 1.
+We set $\alpha=0.95$ to roughly balance the submodular and diversity terms.
+[^3] If $\mathbf{x},\mathbf{y}\sim\mathcal{N}(\mathbf{0},\mathbf{I}_{d})$ are i.i.d., then $\mathbb{E}[\lVert\mathbf{x}-\mathbf{y}\rVert_{2}^{2}]=2d$.
+Setting $\frac{\alpha}{2}\approx(1-\alpha)\sqrt{2d}\implies\alpha\approx 1-\frac{1}{2\sqrt{2d}+1}=0.957$, for $\beta\geq 0.5$.
+We sweep over $\alpha$ and budget $\beta$ in Section D.1, and plot how these hyperparameters change $f(S)$.
+Note that the objective values are not directly comparable as a function of $k$ because of the $\nicefrac{{1}}{{k}}$ normalizing factor (hence the discontinuities), but both terms in the objective are decreasing on average in $k$ for a fixed $\alpha$.
+GIST dominates simple, greedy, and random in all cases.
+The poor performance of greedy is somewhat surprising since it is often competitive for subset selection tasks.
+simple beats greedy for all but the extreme values of $k$, i.e., small or close to $n$, but is always worse than GIST.
+Figure 1: Baseline comparison plotting $f(S_{\text{ALG}})$ for $k\in[n]$ on synthetic data with $n=1000$, $\alpha=0.95$, and $\beta=0.75$.
+5.2 Image Classification
+Our ImageNet data sampling experiment compares the top-1 image classification accuracy achieved by different single-shot subset selection algorithms.
+Setup.
+We use the standard vision dataset ImageNet (Russakovsky et al., 2015) containing ~1.3 million images and 1000 classes.
+We select 10% of the points at random and use them to train an initial ResNet-56 model $\bm{\theta}_{0}$ (He et al., 2016).
+Then we use this initial model to compute an uncertainty value and 2048-dimensional embedding for each example.
+The margin-based uncertainty value of $\mathbf{x}_{i}$ is given by $u_{i}=1-(\Pr(y=b\mid\mathbf{x}_{i};~{}\bm{\theta}_{0})-\Pr(y=b^{\prime}\mid\mathbf{x}_{i};~{}\bm{\theta}_{0}))$, which measure the difference between the probability of the best and second-best class label $b^{\prime}$ for an example.
+Finally, we use the fast maximum inner product search of Guo et al.
+(2020) to build a $\Delta$-nearest neighbor graph in the embedding space using $\Delta=10$ and cosine distance.
+We present all model training hyperparameters in Appendix D.
+We compare GIST with several state-of-the-art benchmarks:
+  • random: We select samples from the dataset uniformly at random (without replacement).
+While this approach is simple and lightweight, it creates diversity in many settings and provides good solutions.
+  • margin (Roth and Small, 2006): Margin sampling selects the top-$k$ points based on how hard they are to classify, i.e., using margin-based uncertainty values $u_{i}$.
+  • $k$-center (Sener and Savarese, 2018): We run the classic greedy algorithm for $k$-center on the $\Delta$-nearest neighbor graph.
+The distance between non-adjacent nodes is implicitly assumed to be very large, which is reasonable in the large budget regime $k=\Omega(n)$.
+  • submod: We select a subset by maximizing the submodular objective function $\alpha_{s}\sum_{i\in S}u_{i}-\beta_{s}\sum_{i,j\in S}s(i,j)$, under the budget constraint $|S|\leq k$.
+Similar pairwise submodular objective functions have been used in (Iyer et al., 2021; Kothawade et al., 2021; Ramalingam et al., 2021).
+Here $s(i,j)$ denotes the similarity between adjacent nodes using the same graph as in $k$-center.
+We tuned for best performance by selecting $\alpha_{s}=0.9$ and $\beta_{s}=0.1$.
+We evaluate GIST using both linear and submodular utility functions.
+Let $f(S)=\alpha\cdot g(S)+(1-\alpha)\cdot\texttt{div}(S)$ for $\alpha\in[0,1]$.
+In Table 1, GIST-linear uses the same linear utility as margin: $g(S)=\sum_{i\in S}u_{i}$ with $\alpha=0.9$.
+GIST uses the same submodular utility as submod: $g(S)=\alpha_{s}\sum_{i\in S}u_{i}-\beta_{s}\sum_{i,j\in S}s(i,j)$ with $\alpha=0.95$, $\alpha_{s}=0.9$, and $\beta_{s}=0.1$.
+We use the same graph as $k$-center for distances.
+Table 1: Top-1 classification accuracy (%) on ImageNet for different single-shot data downsampling algorithms.
+The cardinality constraint $k$ is expressed as a fraction of the ~1.3 million examples.
+The results are the average of 3 trials and the top performance for each $k$ is shown in bold.
+$k$ (of data) random margin $k$-center submod GIST-linear GIST
+30 66.27 65.97 66.70 66.52 66.89 67.23
+40 69.17 69.73 70.06 70.71 70.51 70.75
+50 71.05 72.33 73.09 72.69 72.67 73.14
+60 72.49 73.42 73.60 74.20 74.37 74.30
+70 73.70 74.49 74.24 75.24 75.41 75.32
+80 74.42 75.01 75.17 75.91 75.96 75.45
+90 75.16 75.10 74.97 75.84 76.12 76.03
+Results.
+We run each sampling algorithm with cardinality constraint $k$ on the full dataset to get a subset of examples that we then use to train a new ResNet-56 model.
+We report the top-1 classification accuracy of these models in Table 1.
+GIST with a linear or submodular utility is superior to all baselines.
+Interestingly, there is a cutover value of $k$ where GIST-linear becomes the best algorithm.
+Finally, observe that GIST-linear outperforms margin in all cases.
+This demonstrates how $\texttt{div}(S)$ encourages diversity in the subset of sampled points and ultimately improves downstream model quality.
+Conclusion
+This work presents a novel subset selection problem called MDMS that combines the total utility of the selected points (modeled by a monotone submodular function $g$) with the $\texttt{div}(S)=\min_{u,v\in S:u\neq v}\textnormal{dist}(u,v)$ diversity objective.
+We design and analyze the GIST algorithm, which achieves a $\nicefrac{{1}}{{2}}$-approximation guarantee by solving a series of maximum independent set instances on geometric intersection graphs with the GreedyIndependentSet bicriteria approximation algorithm.
+We complement our algorithm with a $0.5584$-hardness of approximation.
+From the theoretical perspective, it is an interesting open problem to bridge the gap between the algorithmic and hardness results of $0.5$ and $0.5584$ approximations.
+For the special case of a linear utility function $g$, we show that GIST has a $\nicefrac{{2}}{{3}}$-approximation guarantee, and we complement this with a $(\nicefrac{{2}}{{3}}+\varepsilon)$-hardness of approximation, for any $\varepsilon>0$.
+Finally, we present an empirical study showing that GIST steadily outperforms baseline methods for MDMS (in particular greedy) on a simple set of Gaussian points with random weights.
+We conclude by comparing the top-1 image classification accuracy achieved by GIST and three other state-of-the-art data sampling methods for ImageNet, demonstrating the benefit of optimizing for a blend of utility and diversity.
+Appendix A $f(S)$ is Not Submodular
+Consider the instance on four points $N=\{a,b,c,d\}\subseteq\mathbb{R}$ where $a=0$, $b=1$, and $c=d=2$ in the (one-dimensional) Euclidean metric, i.e., all the points are collinear.
+Let $g(S)=0$ for all $S\subseteq V$.
+For $f(S)$ to be submodular, it must hold that for every $S\subseteq T\subseteq N$ and $x\in N\setminus T$,
+$f(S\cup\{x\})-f(S)\geq f(T\cup\{x\})-f(T).$
+However, if $x=b$, $S=\{a,c\}$ and $T=\{a,c,d\}$, we have:
+$f(S\cup\{x\})-f(S)$ $=f(\{a,b,c\})-f(\{a,c\})=1-2=-1$
+$f(T\cup\{x\})-f(T)$ $=f(\{a,b,c,d\})-f(\{a,c,d\})=0-0=0,$
+so $f(S)$ is not submodular.
+In this instance, the function $f$ is not monotone. However, a similar monotone but still not submodular $f$ can be defined by setting $g(S)=\sum_{v\in S}w(v)$ where $w(v)=2$ for every $v\in N$.
+Appendix B Proof of Theorem 3.4
+Lemma B.1.
 […]
@@
 Adding the two sums together in Equation 3 and (4) completes the proof. ∎
-Equipped with this bicriteria approximation guarantee, we can now analyze the approximation ratio of the GIST algorithm.
+See 3.4
 Proof of Theorem 3.4.
 Let $d^{*}$ be the minimum distance between two distinct points in $S^{*}$.
-There are two cases: $d^{*}<\varepsilon d_{\max}$ or $d^{*}\geq\varepsilon d_{\max}$.
-For the first case, outputting the $k$ heaviest points (Line 2) achieves a $(1-\varepsilon)$-approximation.
-To see this, observe that
-$\textnormal{OPT}\geq(1-\alpha)d_{\max}>(1-\alpha)\cdot\frac{d^{*}}{\varepsilon}\implies\varepsilon\cdot\textnormal{OPT}>(1-\alpha)d^{*}.$
-The sum of the $k$ heaviest points upper bounds $w(S^{*})$, so it follows that
-$\textnormal{ALG}\geq\alpha\frac{1}{k}w(S^{*})=\textnormal{OPT}-(1-\alpha)d^{*}>(1-\varepsilon)\textnormal{OPT}.$
-Therefore, we focus on the second case where $d^{*}\geq\varepsilon d_{\max}$. It follows that GIST considers a threshold $d=(1+\varepsilon)^{i}\cdot\varepsilon d_{\max}/2\in D$ such that
+There are two cases: $d^{*}<\varepsilon d_{\max}$ and $d^{*}\geq\varepsilon d_{\max}$.
+In the first case, outputting the $k$ heaviest points (Line 2) achieves a $(1-\varepsilon)$-approximation.
+To see this, first observe that
+$\textnormal{OPT}\geq d_{\max}>\frac{d^{*}}{\varepsilon}\implies\varepsilon\cdot\textnormal{OPT}>d^{*}.$
+The sum of the $k$ heaviest points upper bounds $g(S^{*})$, so we have
+$\textnormal{ALG}\geq g(S^{*})=\textnormal{OPT}-d^{*}>(1-\varepsilon)\textnormal{OPT}.$
+Therefore, we focus on the second case where $d^{*}\geq\varepsilon d_{\max}$. GIST considers a threshold $d=(1+\varepsilon)^{i}\cdot\varepsilon d_{\max}/2\in D$ such that
 $\frac{d^{*}}{2}\in[d,(1+\varepsilon)d)\implies\frac{1}{1+\varepsilon}\cdot\frac{d^{*}}{2}<d\leq\frac{d^{*}}{2}.$
 […]
-$f(T)$ $\geq\alpha\frac{1}{k}w(S^{*})+(1-\alpha)d\geq\alpha\frac{1}{k}w(S^{*})+(1-\alpha)\frac{d^{*}}{2(1+\varepsilon)}.$ (4)
+$f(T)$ $\geq g(S^{*})+d\geq g(S^{*})+\frac{d^{*}}{2(1+\varepsilon)}.$ (5)
 The potential max-diameter update to $S$ on Line 5 gives us
-$\textnormal{ALG}\geq(1-\alpha)d_{\max}\geq(1-\alpha)d^{*}.$ (5)
-Putting Equation 4 and (5) together, for any $p\in[0,1]$, it holds that
-ALG $\geq p\left[\alpha\frac{1}{k}w(S^{*})+(1-\alpha)\frac{d^{*}}{2(1+\varepsilon)}\right]+(1-p)(1-\alpha)d^{*}$
-$=p\cdot\alpha\frac{1}{k}w(S^{*})+\left(1-p+\frac{p}{2(1+\varepsilon)}\right)\cdot(1-\alpha)d^{*}.$
-To maximize the approximation ratio as $\varepsilon\rightarrow 0$, solve $p=1-p/2$ to get $p=2/3$. It follows that
-ALG $\geq\frac{2}{3}\cdot\alpha\frac{1}{k}w(S^{*})+\left(1-\frac{2}{3}+\frac{1}{3(1+\varepsilon)}\right)\cdot(1-\alpha)d^{*}$ (6)
-$=\frac{2}{3}\cdot\alpha\frac{1}{k}w(S^{*})+\frac{1}{3}\left(1+\frac{1}{1+\varepsilon}\right)\cdot(1-\alpha)d^{*}$
-$\geq\frac{2}{3}\cdot\alpha\frac{1}{k}w(S^{*})+\frac{1}{3}\left(2-\varepsilon\right)\cdot(1-\alpha)d^{*}$
-$\geq\frac{2}{3}\left(1-\varepsilon\right)\cdot\alpha\frac{1}{k}w(S^{*})+\frac{2}{3}\left(1-\varepsilon\right)\cdot(1-\alpha)d^{*}$
-$\geq\left(\frac{2}{3}-\varepsilon\right)\cdot\textnormal{OPT},$
+$\textnormal{ALG}\geq d_{\max}\geq d^{*}.$ (6)
+Putting (5) and (6) together, for any $0\leq p\leq 1$, it holds that
+ALG $\geq p\cdot\left(g(S^{*})+\frac{d^{*}}{2(1+\varepsilon)}\right)+(1-p)\cdot d^{*}$
+$=p\cdot g(S^{*})+\left(1-p+\frac{p}{2(1+\varepsilon)}\right)\cdot d^{*}.$
+To maximize the approximation ratio as $\varepsilon\rightarrow 0$, we set $p=2/3$ by solving $p=1-p/2$. It follows that
+ALG $\geq\frac{2}{3}\cdot g(S^{*})+\left(1-\frac{2}{3}+\frac{1}{3(1+\varepsilon)}\right)\cdot d^{*}$ (7)
+$=\frac{2}{3}\cdot g(S^{*})+\frac{1}{3}\left(1+\frac{1}{1+\varepsilon}\right)\cdot d^{*}$
+$\geq\frac{2}{3}\cdot g(S^{*})+\frac{1}{3}\left(2-\varepsilon\right)\cdot d^{*}$
+$\geq\left(\nicefrac{{2}}{{3}}-\varepsilon\right)\cdot\textnormal{OPT},$
 which completes the proof. ∎
-Remark 3.3.
-We can modify GIST to achieve an exact $\nicefrac{{2}}{{3}}$-approximation by trying all pairwise distance thresholds in $D=\{\textnormal{dist}(u,v)/2:u,v\in V\}$.
-This requires $O(n^{2})$ calls to GreedyIndependentSet compared to $O(\log_{1+\varepsilon}(1/\varepsilon))$, but otherwise the same analysis holds for $\varepsilon=0$.
-Finally, we show how to implement GIST in nearly linear time for low-dimensional Euclidean metrics using cover trees for exact nearest-neighbor queries (Beygelzimer et al., 2006; Izbicki and Shelton, 2015) and a $(1-\varepsilon)$-approximation algorithm for point-set diameters $d_{\max}$ (Barequet and Har-Peled, 2001; Imanparast et al., 2018).
-Theorem 3.4.
-For $V\subseteq\mathbb{R}^{d}$ and Euclidean distance, GIST runs in time $O(n^{3}k+n^{2}d)$.
-Furthermore, for fixed dimension $d$, the running time can be improved to $O(n\log n\cdot\log_{1+\varepsilon}(1/\varepsilon)\cdot\textnormal{poly}(d)+1/\varepsilon^{d-1})$.
-We give the proof in Section A.1 and note that the key property is that cover trees allow us to insert a point into a nearest-neighbor index in $O(c^{12}\log n)$ time and find a nearest neighbor to a query point in $O(c^{6}\log n)$ time, where $c=\Theta(d)$ is the doubling dimension of $d$-dimensional Euclidean space.
-4 Hardness of approximation
-[…]
-  1. 1.
-    There is no polynomial-time $(\nicefrac{{2}}{{3}}+\varepsilon)$-approximation algorithm for general metric spaces.
-  2. 2.
-    There is no polynomial-time $(\nicefrac{{3}}{{4}}+\varepsilon)$-approximation algorithm for general metric spaces, even if $k\geq|V|/4$.
-  3. 3.
-    APX-completeness for the Euclidean metric, i.e., there is no polynomial-time approximation scheme (PTAS) for this objective.
-4.1 General metric spaces
-Our first result builds on the hardness of approximating the size of the maximum clique, and it tightly complements the $\nicefrac{{2}}{{3}}$-approximation guarantee that we can achieve with GIST.
-Theorem 4.1.
-[…]
+Appendix C Missing Analysis for Section 4
+C.1 Proof of Theorem: 4.2
+See 4.2
 Proof.
@@
   • NO instance: graph $G$ does not have a clique of size greater than $\delta k$.
-We reduce this instance of the maximum clique decision problem to MDDS with objective function (1) as follows.
+We reduce this instance of the maximum clique decision problem to MDMS with objective function (2) as follows.
 Represent each vertex of graph $G$ with a point in our ground set.
@@
 By setting $\delta<9\varepsilon/(1+6\varepsilon)$, we satisfy the required inequality and achieve the inapproximability gap in the theorem statement. ∎
-4.2 Large subsets
-In many data curation problems, we are interested in instances of MDDS where $k=\Omega(|V|)$. Even in this restricted case, we can show hardness of approximation.
-Theorem 4.2.
-[…]
-The proof of this theorem uses the vertex cover instance constructed in the proof of Håstad (2001, Theorem 8.1) and is given in Section B.1.
-4.3 Euclidean metrics
-[…]
-Lemma 4.3 (Alimonti and Kann (2000, Theorem 3.2)).
-The maximum independent set problem for graphs with degree at most $3$ is APX-complete.
-Similar to our reduction from maximum clique in Section 4.1 for general metric spaces, this construction uses a node embedding function $h_{g}(v)$ to encode graph adjacency in Euclidean space.
-Lemma 4.4.
-Let $G=(V,E)$ be a simple undirected graph with $n=|V|$, $m=|E|$, and maximum degree $\Delta$. There exists an embedding $h_{G}:V\rightarrow\mathbb{R}^{n+m}$ such that if $\{u,v\}\in E$ then
-$\lVert h_{G}(u)-h_{G}(v)\rVert_{2}\leq 1-\frac{1}{2(\Delta+1)},$
-and if $\{u,v\}\not\in E$ then $\lVert h_{G}(u)-h_{G}(v)\rVert_{2}=1$.
-Theorem 4.5.
-The MDDS problem is APX-complete for Euclidean metrics.
-We sketch the main idea of our reduction and defer the complete proofs of Lemma 4.4 and Theorem 4.5 to Appendix B.
-Let $\mathcal{I}(G)$ be the set of independent sets of $G$.
-A set of embedded nodes $S\subseteq V$ (with at least two nodes) of a graph $G$ with maximum degree $\Delta\leq 3$ has the property:
-  • If $S\in\mathcal{I}(G)$: $\texttt{div}(S)=\min_{u,v\in S:u\neq v}\lVert h_{G}(u)-h_{G}(v)\rVert_{2}=1$.
-  • If $S\not\in\mathcal{I}(G)$: $\texttt{div}(S)=\min_{u,v\in S:u\neq v}\lVert h_{G}(u)-h_{G}(v)\rVert_{2}\leq 1-1/(2(\Delta+1))\leq 1-\nicefrac{{1}}{{8}}$.
-This gap between the two cases is at least $\nicefrac{{1}}{{8}}$ (i.e., a universal constant) for all such graphs, and allows us to show that MDDS inherits the APX-complete hardness of bounded-degree max independent set.
-5 Experiments
-5.1 Warm-up: Synthetic dataset
-Setup.
-We start with a simple dataset of weighted random points to see the shortcomings of existing methods for MDDS.
-For a fixed dimension $d$, we generate $n=1000$ points $\mathbf{x}_{i}\in\mathbb{R}^{d}$, where each $\mathbf{x}_{i}\sim\mathcal{N}(\mathbf{0},\mathbf{I}_{d})$ and each point is assigned an i.i.d.
-uniform continuous weight $w_{i}\sim\mathcal{U}_{[0,1]}$.
-We consider three baseline algorithms: random, simple, and greedy.
-The random algorithm chooses a random set of $k$ points, permutes them, and returns the best prefix.
-simple is the $\nicefrac{{1}}{{2}}$-approximation algorithm in Section 2.
-greedy builds a set $S\subseteq V$ one element at a time by selecting the point with index $i^{*}_{t}=\operatorname*{arg\,max}_{i\in V\setminus S_{t-1}}f(S_{t-1}\cup\{i\})-f(S_{t-1})$ at each step $t\in[k]$.
-Then it returns the best prefix $S=\operatorname*{arg\,max}_{t\in[k]}f(S_{t})$, as this sequence of objective values is not necessarily monotone.
-Figure 1: Synthetic data with $n=1000$ and $d=64$. Plots of $f(S_{\text{ALG}})$ for $k\in[n]$, $\alpha\in\{0.90,0.95\}$.
-Results.
-We plot the values of $f(S)$ for our baseline algorithms and GIST for all $k\in[n]$ in Figure 1.
-We set $\alpha\in\{0.90,0.95\}$ to balance the two competing objective terms (which are easy to optimize in isolation).
-[^2] If $\mathbf{x},\mathbf{y}\sim\mathcal{N}(\mathbf{0},\mathbf{I}_{d})$ are i.i.d., then $\mathbb{E}[\lVert\mathbf{x}-\mathbf{y}\rVert_{2}^{2}]=2d$, so $\frac{\alpha}{2}\approx(1-\alpha)\sqrt{2d}\implies\alpha\approx 1-\frac{1}{2\sqrt{2d}+1}=0.957$.
-There is a clear separation in solution quality for each $k\in[n]$ when $\alpha=0.95$.
-Note that these objective values are not directly comparable as a function of $k$ due to the $\nicefrac{{1}}{{k}}$ normalizing factor (hence the discontinuities), but both terms in the objective decreasing on average in $k$ for a fixed $\alpha$: mean utility and min distance of the selected points.
-GIST noticeably outperforms simple, greedy, and random.
-The poor performance of greedy is somewhat surprising since it is normally competitive for submodular-style objectives.
-simple also outperforms greedy for all but the extreme values of $k$, i.e., very small or close to $n$, but is always worse relative to GIST.
-5.2 Image classification
-The following data sampling experiment compares the top-1 image classification accuracy achieved by different single-shot subset selection algorithms.
-Setup.
-We use the standard vision dataset ImageNet (Russakovsky et al., 2015), which contains ~1.3 million images and 1000 classes.
-We select 10% of the data points at random and use them to train an initial ResNet-56 model $\bm{\theta}_{0}$ (He et al., 2016).
-Then we use this initial model to compute an uncertainty value and 2048-dimensional embedding for each example.
-The uncertainty value of $\mathbf{x}_{i}$ is $u_{i}=1-\Pr(y=b\mid\mathbf{x}_{i};~{}\bm{\theta}_{0})$, where $b$ is the best class label according to the initial model.
-Finally, we use the fast maximum inner product search of Guo et al.
-(2020) to build a $\Delta$-nearest neighbor graph in the embedding space using $\Delta=10$ and cosine distance.
-We present all of the model training hyperparameters in Appendix C.
-We compare GIST with three state-of-the-art benchmarks:
-  • random: We select samples from the dataset uniformly at random (and without replacement).
-While being a simple and lightweight approach, this implicitly encodes diversity in many settings and provides good solutions.
-  • margin (Roth and Small, 2006): Margin sampling selects the top-$k$ points based on how hard they are to classify, i.e., using the margin scores $m_{i}=1-(\Pr(y=b\mid\mathbf{x}_{i};~{}\bm{\theta}_{0})-\Pr(y=b^{\prime}\mid\mathbf{x}_{i};~{}\bm{\theta}_{0}))$, which measure the difference between the probability of the best and second-best class label $b^{\prime}$ for an example.
-  • $k$-center (Sener and Savarese, 2018): We run the classic greedy algorithm for $k$-center on the $\Delta$-nearest neighbor graph.
-The distance between non-adjacent nodes is implicitly assumed to be very large, which is acceptable in the large budget regime $k=\Omega(n)$.
-For this task, GIST uses uncertainty values as its weights $w_{i}=u_{i}$ and the same graph as $k$-center, and sets $\alpha=0.85$.
-Results.
-We run each algorithm with cardinality constraint $k$ on the full dataset to get a subset of examples that we use to train a new ResNet-56 model.
-We report the top-1 classification accuracy of these models in Table 1.
-$k$ (of data) random margin $k$-center GIST
-30 65.02 67.64 67.32 67.70
-40 68.92 70.45 70.98 71.20
-50 70.86 73.41 72.51 73.67
-60 72.39 74.70 74.78 75.16
-70 72.92 75.57 75.62 75.85
-80 74.36 75.39 76.26 76.49
-90 75.10 76.11 76.02 76.50
-Table 1: Top-1 classification accuracy (%) on ImageNet for different single-shot data downsampling algorithms. The cardinality constraint $k$ is expressed as a fraction of the ~1.3 million examples.
-Conclusion
-This work proposes a novel subset selection problem called MDDS that combines the total utility of the selected points with the $\texttt{div}(S)=\min_{u,v\in S:u\neq v}\textnormal{dist}(u,v)$ diversity objective.
-We design and analyze the GIST algorithm, which achieves a $\nicefrac{{2}}{{3}}$-approximation guarantee for general metric spaces by solving a series of maximum independent set instances on geometric intersection graphs with the GreedyIndependentSet bicriteria approximation algorithm.
-We also give a nearly linear-time implementation of GIST for points in low-dimensional Euclidean space.
-We complement our algorithm with a $(\nicefrac{{2}}{{3}}+\varepsilon)$-hardness of approximation, for any $\varepsilon>0$, in general metric spaces.
-We also prove hardness of approximation in the practical cases of Euclidean metrics or $k=\Omega(n)$.
-Finally, we present an empirical study showing that GIST steadily outperforms baseline methods for MDDS (in particular greedy) on a simple set of Gaussian points with random weights.
-We conclude with an experiment comparing the top-1 image classification accuracy achieved by GIST and three other state-of-the-art data sampling methods for ImageNet.
-Appendix A Missing analysis for Section 3
-A.1 Proof of Theorem 3.4
-See 3.4
-Proof.
-First, observe that we can compute and store all pairwise distances in $O(n^{2}d)$ time and $O(n^{2})$ space.
-This is useful in the naive implementation because we can then look up the values $\textnormal{dist}(u,v)$ in constant time in any subroutine.
-In the nearly linear-time implementation, we skip this precomputation step and use cover trees for fast exact nearest-neighbor search in low-dimensional Euclidean space (Beygelzimer et al., 2006): inserting a point into the index takes $O(c^{12}\log n)$ time and finding a nearest neighbor in the index to a query point takes $O(c^{6}\log n)$ time, where $c=\Theta(d)$ is the doubling dimension of $d$-dimensional Euclidean space.
-Thus, we can speed up the following subroutines in Algorithm 1 to have near-linear running times in a fixed dimension:
-  • Computing $d_{\min}$: $O(n^{2})\rightarrow O(n\cdot d^{12}\log n)$ after discarding duplicate points.
-  • GreedyIndependentSet: $O(nk)\rightarrow O(n\cdot d^{12}\log n)$
-  • Evaluating $f(S)$: $O(k^{2})\rightarrow O(k\cdot d^{12}\log k)$
-Taking advantage of the error $\varepsilon$ in our $(\nicefrac{{2}}{{3}}-\varepsilon)$-approximation guarantee, we use the approximate point-set diameter algorithm in Imanparast et al.
-(2018, Theorem 1) to compute a $(1-\varepsilon)$-approximate diameter, together with two points realizing it.
-This speeds up the naive running time of Lines 34 in Algorithm 1 as follows:
-  • Computing a $(1-\varepsilon)$-approximate $d_{\max}$: $O(n^{2})\rightarrow O\left(n\cdot 2^{d}d+\frac{(2\sqrt{d})^{2d}}{\varepsilon^{d-1}}\right)$
-Since we have a smaller diameter $(1-\varepsilon)d_{\max}$, we need to show that it is compatible with the proof of Theorem 3.1.
-Concretely, this means modifying Eq.
-(5) and (6), and observing that the inequality
-$\left(1-\frac{2}{3}+\frac{1}{3(1+\varepsilon)}\right)(1-\alpha)\cdot(1-\varepsilon)d^{*}\geq\left(\frac{2}{3}-\varepsilon\right)(1-\alpha)d^{*}$
-still holds.
-Lastly, the number of calls to GreedyIndependentSet (i.e., distance thresholds that GIST tries) is $\min\{\lceil\log_{1+\varepsilon}(2/\varepsilon)\rceil,n^{2}\}$.
-Putting everything together proves the claim.
-∎
-Appendix B Missing analysis for Section 4
-B.1 Proof of Theorem 4.2
-See 4.2
-Proof.
-We use the vertex cover instance constructed in Håstad (2001, Theorem 8.1).
-For any $\varepsilon_{0}>0$, they construct a graph $G$ with $2^{r+2}$ nodes for some integer $r$ and show that it is NP-hard to distinguish between the following two cases:
-  • YES case: Graph $G$ has an independent set of size $2^{r}(1-\varepsilon_{0})$.
-  • NO case: Graph $G$ has no independent set of size larger than $2^{r}(\nicefrac{{1}}{{2}}+\varepsilon_{0})$.
-We mimic the general structure of the proof of Theorem 4.1.
-Add a point for each vertex in graph $G$ and set its weight to $1$.
-Set the distance between two points to $1$ if they share an edge in the graph, and set it to $2$ otherwise.
-Lastly, let $k=2^{r}$.
-Since there are $2^{r+2}$ nodes in the graph, $k=|V|/4$ in this case.
-In the YES case, objective function (1) can achieve the following value by selecting the independent set of size $2^{r}(1-\varepsilon_{0})$:
-OPT $\geq\alpha\cdot\frac{1}{2^{r}}\cdot(2^{r}(1-\varepsilon_{0}))+(1-\alpha)\cdot 2$
-$=2-(1+\varepsilon_{0})\cdot\alpha.$
-In the NO case, the largest value that can be achieved is the best of two scenarios: (a) selecting $k$ points with minimum distance $1$, or (b) selecting an independent set of size at most $2^{r}(\nicefrac{{1}}{{2}}+\varepsilon_{0})$ and achieving minimum distance $2$.
-This results in the following upper bound for the objective:
-ALG $\leq\max\left\{\alpha+(1-\alpha)\cdot 1,\alpha\cdot(\nicefrac{{1}}{{2}}+\varepsilon_{0})+(1-\alpha)\cdot 2\right\}$
-$=\max\left\{1,2-(\nicefrac{{3}}{{2}}-\varepsilon_{0})\cdot\alpha\right\}.$
-The largest hardness gap occurs when the two terms of the $\max$ expression become equal, which happens at $\alpha=1/(\nicefrac{{3}}{{2}}-\varepsilon_{0})$ and yields the upper bound $\textnormal{ALG}\leq 1$.
-The optimum value OPT will be lower bounded by:
-OPT $\geq 2-(1+\varepsilon_{0})\cdot\alpha$
-$=2-\frac{1+\varepsilon_{0}}{\nicefrac{{3}}{{2}}-\varepsilon_{0}}$
-$=\frac{2-3\varepsilon_{0}}{\nicefrac{{3}}{{2}}-\varepsilon_{0}}.$
-The desired inapproximability gap of $\nicefrac{{3}}{{4}}+\varepsilon$ in Theorem 4.2 is achieved when the ratio between the upper bound on ALG and lower bound for OPT drops below it.
-In other words, we want to set $\varepsilon_{0}$ such that the following inequality holds:
-$\frac{2-3\varepsilon_{0}}{\nicefrac{{3}}{{2}}-\varepsilon_{0}}>\frac{1}{\nicefrac{{3}}{{4}}+\varepsilon}.$
-We set $\varepsilon_{0}\coloneqq\varepsilon$ and prove that the above inequality holds.
-We first note that both the numerator and denominator of the left side ratio are positive.
-Therefore, we can invert both ratios and swap their places to achieve the following equivalent inequality:
-$\frac{3}{4}+\varepsilon>\frac{\nicefrac{{3}}{{2}}-\varepsilon_{0}}{2-3\varepsilon_{0}}\iff\varepsilon>\frac{\nicefrac{{3}}{{2}}-\varepsilon_{0}}{2-3\varepsilon_{0}}-\frac{3}{4}=\frac{6-4\varepsilon_{0}-6+9\varepsilon_{0}}{8-12\varepsilon_{0}}>\frac{5\varepsilon_{0}}{8}.$
-The left-hand side inequality $\varepsilon>\nicefrac{{5\varepsilon_{0}}}{{8}}$ holds by our choice of $\varepsilon_{0}=\varepsilon$. ∎
-B.2 Proof of Lemma 4.4
+C.2 Proof of Lemma 4.4
 Let $G=(V,E)$ be a simple undirected graph.
@@
 Construction.
-First, let $n=|V|$ and $m=|E|$.
+Let $n=|V|$ and $m=|E|$.
 Augment $G$ by adding a self-loop to each node to get $G^{\prime}=(V,E^{\prime})$.
@@
 Now suppose that $\{u,v\}\in E$. It follows that
-$\lVert h_{G}(u)-h_{G}(v)\rVert_{2}^{2}$
-$=\sum_{e\in N^{\prime}(u)\setminus\{v\}}\frac{1}{2\deg^{\prime}(u)}+\sum_{e\in N^{\prime}(v)\setminus\{u\}}\frac{1}{2\deg^{\prime}(v)}+\left(\sqrt{\frac{1}{2\deg^{\prime}(u)}}-\sqrt{\frac{1}{2\deg^{\prime}(v)}}\right)^{2}$
+$\lVert h_{G}(u)-h_{G}(v)\rVert_{2}^{2}$ $=\sum_{e\in N^{\prime}(u)\setminus\{v\}}\frac{1}{2\deg^{\prime}(u)}+\sum_{e\in N^{\prime}(v)\setminus\{u\}}\frac{1}{2\deg^{\prime}(v)}+\left(\sqrt{\frac{1}{2\deg^{\prime}(u)}}-\sqrt{\frac{1}{2\deg^{\prime}(v)}}\right)^{2}$
 $=\frac{\deg^{\prime}(u)-1}{2\deg^{\prime}(u)}+\frac{\deg^{\prime}(v)-1}{2\deg^{\prime}(v)}+\left(\frac{1}{2\deg^{\prime}(u)}+\frac{1}{2\deg^{\prime}(v)}-2\sqrt{\frac{1}{4\deg^{\prime}(u)\deg^{\prime}(v)}}\right)$
@@
 which completes the proof. ∎
-B.3 Proof of Theorem 4.5
+C.3 Proof of Theorem 4.5
 See 4.5
@@
   • NO instance: graph $G$ does not have an independent set of size greater than $(1-\varepsilon_{0})k$.
-We reduce this instance of bounded-degree maximum independent set to MDDS with objective function (1) as follows.
+We reduce this instance of bounded-degree maximum independent set to MDMS with objective function (2) as follows.
 Embed each node of the graph $G$ into Euclidean space using the function $h_{G}(v)$ in Lemma 4.4.
@@
 $1-\varepsilon_{0}\cdot\alpha=1-\frac{\varepsilon_{0}}{\nicefrac{{1}}{{\varepsilon_{0}}}+8}=1-\varepsilon_{1}.$
-Since $\varepsilon_{0}>0$ is a constant, $\varepsilon_{1}\coloneqq\varepsilon_{0}/(\nicefrac{{1}}{{\varepsilon_{0}}}+8)>0$ is also a positive constant. This completes the proof of APX-completeness. ∎
-Appendix C Missing details for Section 5
+Since $\varepsilon_{0}>0$ is a constant, $\varepsilon_{1}\coloneqq\varepsilon_{0}/(\nicefrac{{1}}{{\varepsilon_{0}}}+8)>0$ is also a constant. This completes the proof of APX-completeness. ∎
+Appendix D Additional Details for Section 5
+D.1 Synthetic Dataset
+We extend our comparison of baseline algorithms in Section 5.1 by sweeping over values of $\alpha,\beta$ in the objective function.
+Figure 2: Baseline comparison with $\alpha\in(0.85,0.90,0.95,1.00)$ and $\beta=0.75$.
+Figure 3: Baseline comparison with $\alpha=0.95$ and $\beta\in(0.60,0.70,0.80,0.90)$.
+D.2 Image Classification
 Hyperparameters for ImageNet classification.
@@
 We extract the penultimate layer features to produce $2048$-dimensional embeddings of each image.
+We use the same hyperparameters as the original ResNet paper (He et al., 2016) with budgets and one-shot subset selection experiments designed in the same manner as Ramalingam et al.
+(2021).

### v3

@@
-GIST: Greedy Independent Set Thresholding for Diverse Data Summarization
+GIST: Greedy Independent Set Thresholding for Max-Min Diversification with Submodular Utility
 Abstract
-We introduce a novel subset selection problem called min-distance diversification with monotone submodular utility (MDMS), which has a wide variety of applications in machine learning, e.g., data sampling and feature selection.
-Given a set of points in a metric space, the goal of MDMS is to maximize an objective function combining a monotone submodular utility term and a min-distance diversity term between any pair of selected points, subject to a cardinality constraint.
-We propose the GIST algorithm, which achieves a $\nicefrac{{1}}{{2}}$-approximation guarantee for MDMS by approximating a series of maximum independent set problems with a bicriteria greedy algorithm.
-We also prove that it is NP-hard to approximate to within a factor of $0.5584$.
-Finally, we demonstrate that GIST outperforms existing benchmarks for on a real-world image classification task that studies single-shot subset selection for ImageNet.
+This work studies a novel subset selection problem called max-min diversification with monotone submodular utility (MDMS), which has a wide range of applications in machine learning, e.g., data sampling and feature selection.
+Given a set of points in a metric space, the goal of MDMS is to maximize $f(S)=g(S)+\lambda\cdot\texttt{div}(S)$ subject to a cardinality constraint $|S|\leq k$, where $g(S)$ is a monotone submodular function and $\texttt{div}(S)=\min_{u,v\in S:u\neq v}\textnormal{dist}(u,v)$ is the max-min diversity objective.
+We propose the GIST algorithm, which gives a $\nicefrac{{1}}{{2}}$-approximation guarantee for MDMS by approximating a series of maximum independent set problems with a bicriteria greedy algorithm.
+We also prove that it is NP-hard to approximate within a factor of $0.5584$.
+Finally, we show in our empirical study that GIST outperforms state-of-the-art benchmarks for a single-shot data sampling task on ImageNet.
 1 Introduction
-Subset selection is a challenging combinatorial optimization problem with a wide variety of applications in machine learning, including feature selection, recommender systems, drug discovery, data summarization, and designing pretraining sets for large language models (Team et al., 2024).
-Data sampling in particular is a salient problem due to unprecedented and continuous streams of data collection.
-For example, LiDAR and imaging devices in one self-driving vehicle can easily capture ~80 terabytes of data per day (Kazhamiaka et al., 2021).
-In many subset selection tasks, we rely on the utility (or weight) of an object to rank one over the other, and also to avoid selecting duplicate or near-duplicate objects.
-If we select a small subset, then we must ensure that the selected set is a good representation of the original set.
-These utility, diversity, and coverage criteria are expressed through objective functions, and the interesting research lies in developing efficient algorithms with strong approximation guarantees.
-The algorithmic tools used in constrained subset selection shares many similarities with techniques from other areas of combinatorial optimization such as submodular maximization, $k$-center clustering, and convex hull approximations.
-This work introduces a novel subset selection problem called min-distance diversification with monotone submodular utility (MDMS).
-Given $n$ points $V$ in a metric space, a monotone submodular function $g:2^{V}\rightarrow\mathbb{R}_{\geq 0}$, and a cardinality constraint $k$, the goal is to solve
-$S^{*}$ $=\operatorname*{arg\,max}_{S\subseteq V}~{}g(S)+\texttt{div}(S)$ (1)
-subject $\hskip 13.6572pt\text{subject to}~{}~{}|S|\leq k,$
-where the diversity term $\texttt{div}(S)=\min_{u,v\in S:u\neq v}\textnormal{dist}(u,v)$ encourages maximizing the min distance between any two distinct points in the selected set.
-Meinl et al.
-(2011) call $\texttt{div}(S)$ the $p$-dispersion measure, and Bhaskara et al.
-(2016) call it the min-min diversity objective, where is known to admit a $\nicefrac{{1}}{{2}}$-approximation ratio subject to the strict cardinality constraint $|S|=k$.
-Other diversity terms such as $\sum_{u\in S}\min_{v\in S\setminus\{u\}}\textnormal{dist}(u,v)$ and $\sum_{u\in S}\sum_{v\in S}\textnormal{dist}(u,v)$ have also been studied (Meinl et al., 2011; Bhaskara et al., 2016; Borodin et al., 2017).
-While previous works have explored maximizing $\texttt{div}(S)$ in isolation, MDMS introduces a utility term and has not been studied before.
-Bhaskara et al.
-(2016) remark that $\texttt{div}(S)$ is a powerful diversity objective, but point out a major challenge: it is highly non-monotonic, i.e., it can exhibit a stark decrease in the objective value by adding a single new point.
-Our work shows combining a monotone submodular utility term with $\texttt{div}(S)$ mitigates some of the challenges associated with this diversity measure, making it easier to leverage its properties.
+Subset selection is a ubiquitous and challenging problem at the intersection of machine learning and combinatorial optimization, finding applications in areas such as feature selection, recommendation systems, and data summarization, including the critical task of designing pretraining sets for large language models [59, 3].
+Data sampling in particular is an increasingly important problem given the unprecedented and continuous streams of data collection.
+For example, one self-driving vehicle generates ~80 terabytes of data daily from LiDAR and imaging devices [34], and academic datasets have scaled dramatically from 1.2M images in ImageNet [54] to 5B in Laion [56].
+Subset selection often involves balancing competing objectives: we rely on the utility (or weight) of individual items to prioritize them, while simultaneously trying to avoid selecting duplicate or near-duplicate items to ensure diversity.
+When selecting a small subset, this process should guarantee that the chosen set is a good representation of the original dataset—often called coverage.
+The intricate trade-offs between utility, diversity, and coverage are often expressed through an objective function.
+It is often a significant challenge to design efficient algorithms with strong approximation guarantees for constrained subset selection problems, even when leveraging tools and techniques from combinatorial optimization such as submodular maximization, $k$-center clustering, and convex hull approximations.
+This paper studies a novel subset selection problem called max-min diversification with monotone submodular utility (MDMS).
+Given $n$ points $V$ in a metric space, a nonnegative monotone submodular function $g:2^{V}\rightarrow\mathbb{R}_{\geq 0}$, diversity strength $\lambda\geq 0$, and cardinality constraint $k$, our goal is to solve
+$S^{*}$ $=\operatorname*{arg\,max}_{S\subseteq V}~g(S)+\lambda\cdot\texttt{div}(S)$ (1)
+$\hskip 13.6572pt\text{subject to}~~|S|\leq k,$
+where $\texttt{div}(S)=\min_{u,v\in S:u\neq v}\textnormal{dist}(u,v)$ is the max-min diversification term [2, 40, 38].
+Maximizing the submodular term $g(S)$ aims to select points that are the most valuable or informative.
+A purely utility-driven algorithm, however, can lead to redundancies where very similar points are chosen.
+To counteract this, we add the $\lambda\cdot\texttt{div}(S)$ term to encourage selecting points that are well spread out in the metric space, which effectively penalizes subsets with closely clustered elements.
+Note that $\lambda$ is a knob that controls the trade-off between these two terms, acting as a regularization strength for subset selection problems.
+The most relevant prior work is by Borodin et al.
+[13].
+They combine a monotone submodular utility and the max-sum diversity term $\sum_{u,v\in S}\textnormal{dist}(u,v)$, in contrast to our max-min $\texttt{div}(S)$ in (1).
+They give a greedy $\nicefrac{{1}}{{2}}$-approximation algorithm and extend their results to general matroid constraints via local search.
+An earlier contribution by Gollapudi and Sharma [26] combined a linear utility (instead of a monotone submodular function) and the max-sum diversity term, subject to the strict constraint $|S|=k$, to design a ranking system for search engines.
+These approaches focus on maximizing the sum of pairwise distances, but our max-min formulation ensures that the selected points have high utility and are maximally separated, leading to less redundant and more representative subsets.
 1.1 Our contributions
-We summarize our main contributions below:
-  • In Section 2, we formalize the MDMS problem and present a simple $0.387$-approximation algorithm as a warm-up, demonstrating how the competing terms in the objective function interact.
-  • In Section 3, we present the GIST algorithm, which achieves a $\nicefrac{{1}}{{2}}$-approximation guarantee for MDMS by approximating a series of maximum independent set problems with a bicriteria greedy algorithm and returning the best solution.
+We summarize the main contributions of this work below:
+  • In Section˜2, we formalize the MDMS problem and present a simple $0.387$-approximation algorithm as a warm-up to show how the competing terms in the objective function interact.
+  • In Section˜3, we present the GIST algorithm, which achieves a $\nicefrac{{1}}{{2}}$-approximation guarantee for MDMS by approximating a series of maximum independent set problems with a bicriteria greedy algorithm and returning the best solution.
 In the special case of linear utility functions $g(S)=\sum_{v\in S}w(v)$, we prove that GIST offers a stronger $\nicefrac{{2}}{{3}}$-approximation guarantee.
-[…]
-  • In Section 5, our experiments show that GIST outperforms baselines for MDMS on synthetic data (in particular the classic greedy algorithm).
+  • In Section˜4, we prove that it is NP-hard to approximate MDMS to within a factor of $0.5584$ via a careful reduction from the maximum coverage problem.
+For linear utilities, we match the guarantees of GIST and prove a tight $(\nicefrac{{2}}{{3}}+\varepsilon)$-hardness of approximation, for any $\varepsilon>0$, assuming $\text{P}\neq\text{NP}$.
+In the more restricted case of a linear utility function and the Euclidean metric, we prove APX-completeness.
+  • In Section˜5, our experiments show that GIST outperforms baselines for MDMS on synthetic data (in particular the classic greedy algorithm).
 Then we show that GIST can be used to build better single-shot subsets of training data for an image classification benchmark on ImageNet compared to margin sampling and $k$-center algorithms, demonstrating the benefit of optimizing for a blend of utility and diversity.
 1.2 Related work
-Diversity objectives.
-A closely related problem is $k$-center, which aims to find a subset of $k$ centers that minimize the largest distance of any point to its nearest center in $S$.
-Concretely, the objective is $\max_{v\in V}\min_{c\in S}\textnormal{dist}(v,c)$, and a greedy $2$-approximation algorithm is widely used in active learning applications (Sener and Savarese, 2018).
-The main difference between MDMS and $k$-center is the ability to benefit from the utility term, e.g., uncertainty values in an active learning setting.
-The second key difference is that we relax the requirement to cover every point in the dataset, which can be helpful for large datasets where we want to avoid instability due to certain corner or outlier points.
+Submodular maximization.
+Submodular maximization subject to a cardinality constraint $k$ is an NP-hard problem, i.e., $\max_{S\subseteq V:|S|\leq k}g(S)$.
+For monotone submodular functions, Nemhauser et al.
+[47] proved that the greedy algorithm achieves a $(1-\nicefrac{{1}}{{e}})$-approximation, which is asymptotically optimal unless $\text{P}=\text{NP}$ [23].
+The non-monotone case is substantially less understood, but Buchbinder and Feldman [14] recently gave a 0.401-approximation algorithm and Qi [48] improved the hardness of approximation to 0.478.
+Over the last decade, there has also been significant research on distributed algorithms for submodular maximization in the MapReduce [44, 43, 9, 10, 39, 33] and low-adaptivity models [7, 19, 22, 8, 5, 16, 17].
+Diversity maximization.
+The related max-min diversification problem $\max_{S\subseteq V:|S|=k}\texttt{div}(S)$ has a rich history in operations research due to its connection to facility location, and is also called the $p$-dispersion problem.
+It uses the strict constraint $|S|=k$; otherwise, if $|S|\leq k$ we can maximize $\texttt{div}(S)$ with two diametrical points.
+Tamir [58] gave a simple greedy $\nicefrac{{1}}{{2}}$-approximation algorithm for max-min diversification, and Ravi et al.
+[51] proved $(\nicefrac{{1}}{{2}}+\varepsilon)$-inapproximability, for any $\varepsilon>0$, unless $\text{P}=\text{NP}$.
+Indyk et al.
+[30] gave a distributed $\nicefrac{{1}}{{3}}$-approximation algorithm via composable coresets for diversity and coverage maximization.
+Borassi et al.
+[12] designed a $\nicefrac{{1}}{{5}}$-approximation algorithm in the sliding window model.
+For asymmetric distances, Kumpulainen et al.
+[38] recently designed a $1/(6k)$-approximation algorithm.
+There are also many related works on mixed-integer programming formulations, heuristics, and applications to drug discovery [20, 52, 55, 60, 37, 40, 61, 42].
 Bhaskara et al.
-(2016) study various distance-based diversity metrics subject to a cardinality constraint.
-In particular, they give a $\nicefrac{{1}}{{8}}$-approximation algorithm for the sum-min diversity objective defined as $\sum_{u\in S}\min_{v\in S\setminus\{u\}}\textnormal{dist}(u,v)$ by rounding a linear programming relaxation.
-They complement this result by proving a conditional $\nicefrac{{1}}{{2}}$-hardness of approximation under the planted clique assumption.
-They also study the min-min objective $\min_{u,v\in S:u\neq v}\textnormal{dist}(u,v)$, but they enforce a strict cardinality constraint $|S|=k$; otherwise, it is optimal to only output one or two points.
-Meinl et al.
-(2011) study the same objective subject to $|S|=k$ under the name $p$-dispersion (also called max-min or remote-edge) diversity maximization.
-This exact cardinality issue does not arise in MDMS because we need to balance the trade-off between competing monotone submodular utility and diversity terms.
-Moumoulidou et al.
-(2020), Addanki et al.
-(2022), and Mahabadi and Trajanovski (2023) study diversity maximization under fairness constraints.
-Given a partitioning of the dataset, they ensure that a specified number of points $k_{i}$ are selected from each group $i\in[m]$.
-This fairness constraint considerably differentiates their problem setting from ours (i.e., $|S|=k=\sum_{i=1}^{m}k_{i}$).
-Moumoulidou et al.
-(2020); Addanki et al.
-(2022) maximize $\texttt{div}(S)$ in this setting.
-Mahabadi and Trajanovski (2023) consider sum-of-all pairwise distances $\sum_{u\in S}\sum_{v\in S}\textnormal{dist}(u,v)$ and sum-min objectives.
-However, none of these works add a utility term to the objective.
-Borodin et al.
-(2017) optimize the combination of a submodular utility and sum-of-all pairwise distances in the selected set, making it the most similar work to MDMS.
-Unlike our $\texttt{div}(S)$ term, max-sum diversity is not very sensitive to duplicate points.
-In fact, adding $m$ copies of the same point to a set raises the increment of the diversity term by a factor of $m$ in their model.
-By using $\texttt{div}(S)$ instead, we can strictly avoid duplication across all selected points.
+[11] studied the sum-min diversity problem $\max_{S\subseteq V:|S|\leq k}\sum_{u\in S}\textnormal{dist}(u,S\setminus\{u\})$, providing the first constant-factor approximation algorithm with a $\nicefrac{{1}}{{8}}$ guarantee.
+Their algorithm is based on a novel linear-programming relaxation and generalizes to matroid constraints.
+They also proved the first inapproximability result of $\nicefrac{{1}}{{2}}$ under the planted clique assumption.
+Diversity maximization has also recently been studied with fairness constraints.
+Given a partition of the points into $m$ groups, they ensure $k_{i}\in[\ell_{i},u_{i}]$ points are selected from each group $i\in[m]$, subject to $|S|=k=\sum_{i=1}^{m}k_{i}$.
+This line of work can be categorized as studying fair max-min [46, 2, 61], sum-min [11, 41], and max-sum [1, 41] diversity objectives.
 Data sampling.
 In the realm of dataset curation and active learning, there have been several lines of work that study the combination of utility and diversity terms.
+Greedy coreset methods based on $k$-center have been successfully used for data selection with and without utility terms [50, 57].
 Ash et al.
-(2020) invoke $k$-means++ seeding over the gradient space of a model in order to balance uncertainty and diversity.
+[6] use $k$-means++ seeding over the gradient space of a model to balance uncertainty and diversity.
 Wei et al.
-(2015) introduce several submodular objectives, e.g., facility location, and use them to diversify a set of uncertain examples in each active learning iteration.
+[62] introduce several submodular objectives, e.g., facility location, and use them to diversify a set of uncertain examples in each active learning iteration.
 Citovsky et al.
-(2021) cluster examples that are represented by embeddings extracted from the penultimate layer of a partially trained DNN, and use these clusters to diversify uncertain examples in each iteration.
+[18] cluster examples represented by embeddings extracted from the penultimate layer of a partially trained DNN, and use these clusters to diversify uncertain examples in each iteration.
 Our work differs from these and several others (e.g., Kirsch et al.
-(2019); Zhdanov (2019)) in that we directly incorporate the utility and diversity terms in our objective function.
+[35], Zhdanov [63]) in that we directly incorporate the utility and diversity terms into the objective function.
 2 Preliminaries
-Here we formalize the MDMS problem and present a simple $0.387$-approximation algorithm as a starting point.
-Submodular set function.
-For a set function $g:2^{V}\rightarrow\mathbb{R}$ and any $S,T\subseteq V$, let $g(T\mid S)=g(S\cup T)-g(T)$ denote the marginal gain of $g$ at $T$ with respect to $S$.
-A function $g$ is submodular if for every $S\subseteq T\subseteq V$ and $v\in V\setminus T$, we have $g(v\mid S)\geq g(v\mid T)$, where we overload the marginal gain notation for singleton sets.
+Submodular function.
+For any $g:2^{V}\rightarrow\mathbb{R}_{\geq 0}$ and $S,T\subseteq V$, let $g(S\mid T)=g(S\cup T)-g(T)$ be the marginal gain of $g$ at $S$ with respect to $T$.
+A function $g$ is submodular if for every $S\subseteq T\subseteq V$ and $v\in V\setminus T$, we have $g(v\mid S)\geq g(v\mid T)$, where we overload the marginal gain notation for singletons.
 A submodular function $g$ is monotone if for every $S\subseteq T\subseteq V$, $g(S)\leq g(T)$.
-[…]
-Min-distance diversity.
-For a set of $n$ points $V$ in a metric space, define the min-distance diversity function
-$\texttt{div}(S)$ =$=\begin{cases}\min_{u,v\in S:u\neq v}\textnormal{dist}(u,v)&\text{if }|S|\geq 2,\\ \max_{u,v\in V}\textnormal{dist}(u,v)&\text{if }|S|\leq 1.\end{cases}$
-For $\texttt{div}(S)$ to be well-defined, we set the diversity to be the diameter of $V$ if $|S|\leq 1$, i.e., $d_{\max}=\max_{u,v\in V}\textnormal{dist}(u,v)$.
-This implies that it is always beneficial to select at least two points to maximize the overall objective.
-We extend the distance function to take subsets $S\subseteq V$ as input, i.e., $\textnormal{dist}(u,S)=\min_{v\in S}\textnormal{dist}(u,v)$.
-For the sake of completeness, we define $\textnormal{dist}(u,\varnothing)=\infty$.
-Optimization problem.
-Define the objective function
-$f(S)=g(S)+\texttt{div}(S).$ (2)
-The goal of the min-distance diversification with monotone submodular utility (MDMS) problem is to maximize $f(S)$ subject to a cardinality constraint $k$:
+Max-min diversity.
+For a set of $n$ points $V$ in a metric space, define the max-min diversity function as
+$\texttt{div}(S)$ $=\begin{cases}\min_{u,v\in S:u\neq v}\textnormal{dist}(u,v)&\text{if }|S|\geq 2,\\ \max_{u,v\in V}\textnormal{dist}(u,v)&\text{if }|S|\leq 1.\end{cases}$
+We take $\texttt{div}(S)$ to be the diameter of $V$ if $|S|\leq 1$ so that it is monotone decreasing.
+We extend the distance function to take subsets $S\subseteq V$ as input in the standard way, i.e., $\textnormal{dist}(u,S)=\min_{v\in S}\textnormal{dist}(u,v)$, and define $\textnormal{dist}(u,\varnothing)=\infty$.
+MDMS problem statement.
+For any nonnegative monotone submodular function $g:2^{V}\rightarrow\mathbb{R}_{\geq 0}$ and $\lambda\geq 0$, let
+$f(S)=g(S)+\lambda\cdot\texttt{div}(S).$ (2)
+The max-min diversification with monotone submodular utility (MDMS) problem is to maximize $f(S)$ subject to a cardinality constraint $k$:
 $S^{*}=\operatorname*{arg\,max}_{S\subseteq V:|S|\leq k}f(S).$
@@
 Remark 2.1.
-The objective function $f(S)$ is not submodular. We present a counterexample in Appendix A.
+The objective function $f(S)$ is not submodular (see Appendix˜A for a counterexample).
 Intersection graph.
-Let $G_{d}(V)$ be the intersection graph of $V$ for distance threshold $d$, i.e., with nodes $V$ and edges
-$E=\{(u,v)\in V^{2}:\textnormal{dist}(u,v)<d\}.$
-Note that for any independent set $S$ of $G_{d}(V)$ and pair of distinct nodes $u,v\in S$, we have $\textnormal{dist}(u,v)\geq d$.
+Let $G_{d}(V)$ be the intersection graph of $V$ for distance threshold $d$, i.e., with nodes $V$ and edges $E=\{(u,v)\in V^{2}:u\neq v,\textnormal{dist}(u,v)<d\}.$ For any independent set $S$ of $G_{d}(V)$ and pair of distinct nodes $u,v\in S$, we have $\textnormal{dist}(u,v)\geq d$.
 2.1 Warm-up: Simple $0.387$-approximation algorithm
-The objective function $f(S)$ has two terms, each of which is easy to (approximately) optimize in isolation. For the $g(S)$ term, run the greedy algorithm to get $S_{1}$ satisfying
+The objective function $f(S)$ is composed of two terms, each of which is easy to (approximately) optimize in isolation.
+For the submodular term $g(S)$, run the greedy algorithm to get $S_{1}$ satisfying
 $g(S_{1})$ $\geq(1-\nicefrac{{1}}{{e}})\cdot\max_{|S|\leq k}g(S)$
@@
 For the $\texttt{div}(S)$ term, let $S_{2}$ be a pair of diametrical points in $V$.
-The algorithm then flips a biased coin and outputs $S_{1}$ with probability $p\approx 0.613$, and $S_{2}$ otherwise.
-This achieves a $0.387$-approximation guarantee:
-$\mathbb{E}[\textnormal{ALG}_{\textnormal{simple}}]$ $\geq p\cdot g(S_{1})+(1-p)\cdot\texttt{div}(S_{2})$
-$\geq p\cdot(1-\nicefrac{{1}}{{e}})\cdot g(S^{*})+(1-p)\cdot\texttt{div}(S^{*})$
+Then, return the better of the two solutions.
+This gives a $0.387$-approximation guarantee because, for any $0\leq p\leq 1$, we have:
+$\textnormal{ALG}_{\textnormal{simple}}$ $=\max\left\{f(S_{1}),f(S_{2})\right\}$
+$\geq p\cdot g(S_{1})+(1-p)\cdot\lambda\cdot\texttt{div}(S_{2})$
+$\geq p\cdot(1-\nicefrac{{1}}{{e}})\cdot g(S^{*})+(1-p)\cdot\lambda\cdot\texttt{div}(S^{*})$
 $=\frac{e-1}{2e-1}\cdot\textnormal{OPT},$
-where we take $p=e/(2e-1)$ by solving $p\cdot(1-\nicefrac{{1}}{{e}})=1-p$. One of the main goals of this paper is to improve over this baseline and prove complementary hardness results.
+where we set p=e/(2e1)p=e/(2e-1) at the end by solving $p\cdot(1-\nicefrac{{1}}{{e}})=1-p$.
+One of our main goals is to improve over this baseline and prove complementary hardness of approximation results.
+One of the most common algorithms for subset selection problems is the greedy algorithm that iteratively adds the item with maximum marginal value with respect to the objective function $f$.
+We show in Appendix B that the greedy algorithm does not provide a constant-factor approximation guarantee.
 3 Algorithm
-In this section, we present the GIST algorithm and prove that it achieves an approximation ratio of $(\nicefrac{{1}}{{2}}-\varepsilon)$ for MDMS.
-GIST works by sweeping over multiple distance thresholds $d$ and calling a GreedyIndependentSet subroutine to find a high-value independent set of the intersection graph $G_{d}(V)$.
-This is in contrast to the simple $0.387$-approximation algorithm that just tries two extreme thresholds $D=\{0,d_{\max}\}$.
-The GreedyIndependentSet heuristic builds a set $S$ (starting from the empty set) by repeatedly adding $v\in V\setminus S$ with the highest marginal gain with respect to the submodular function $g$ while satisfying $\textnormal{dist}(v,S)\geq d$.
+In this section, we present the GIST algorithm for the MDMS problem and prove that it achieves an approximation ratio of $\nicefrac{{1}}{{2}}-\varepsilon$.
+At a high level, GIST works by sweeping over multiple distance thresholds $d$ and calling a GreedyIndependentSet subroutine on the intersection graph $G_{d}(V)$ to find a high-valued maximal independent set.
+This is in contrast to the warm-up $0.387$-approximation algorithm, which only considers the two extreme thresholds $d\in\{0,d_{\max}\}$.
+GreedyIndependentSet builds the set $S$ (starting from the empty set) by iteratively adding $v\in V\setminus S$ with the highest marginal gain with respect to the submodular utility $g$ while satisfying $\textnormal{dist}(v,S)\geq d$.
 This subroutine runs until either $|S|=k$ or $S$ is a maximal independent set of $G_{d}(V)$.
-[…]
-Algorithm 1 Diverse data summarization for submodular utilities using greedy-weighted maximal independent sets.
-1:function GIST(points $V$, monotone submodular function $g:V\rightarrow\mathbb{R}_{\geq 0}$, budget $k$, error $\varepsilon$)
-2:     Initialize $S\leftarrow\textnormal{{GreedyIndependentSet}}(V,g,0,k)$triangleright runs the classic greedy algorithm
+GIST computes multiple solutions and returns the one with maximum value $f(S)$.
+It first runs the classic greedy algorithm for monotone submodular functions to get an initial solution, which is equivalent to calling GreedyIndependentSet with distance threshold $d=0$.
+Then, it considers the set of distance thresholds $D\leftarrow\{(1+\varepsilon)^{i}\cdot\varepsilon d_{\max}/2:(1+\varepsilon)^{i}\leq 2/\varepsilon\mbox{ and }i\in\mathbb{Z}_{\geq 0}\}$.
+The set $D$ contains a threshold close to the target $d^{*}/2$ where $d^{*}=\texttt{div}(S^{*})$.
+[^1] If this is not true, we are in the special case where $d^{*}\leq\varepsilon d_{\max}$, which we analyze separately.
+For each $d\in D$, GIST calls $\textnormal{{GreedyIndependentSet}}(V,g,d,k)$ to find a set $T$ of size at most $k$.
+If $T$ has a larger objective value than the best solution so far, it updates $S\leftarrow T$.
+After iterating over all thresholds in $D$, it returns the highest-value solution among all candidates.
+Algorithm 1 Max-min diversification with submodular utility via greedy weighted independent sets.
+1:function GIST(points $V$, monotone submodular function $g:2^{V}\rightarrow\mathbb{R}_{\geq 0}$, budget $k$, error $\varepsilon$)
+2:  Initialize $S\leftarrow\textnormal{{GreedyIndependentSet}}(V,g,0,k)$ $\triangleright$ classic greedy algorithm
 3:  Let $d_{\max}=\max_{u,v\in V}\textnormal{dist}(u,v)$ be the diameter of $V$
@@
 6:   Update $S\leftarrow T$
-7:     Let $D\leftarrow\{(1+\varepsilon)^{i}\cdot\varepsilon d_{\max}/2:(1+\varepsilon)^{i}\leq 2/\varepsilon\mbox{ and }i\in\mathbb{Z}_{\geq 0}\}$triangleright distance thresholds
+7:  Let $D\leftarrow\{(1+\varepsilon)^{i}\cdot\varepsilon d_{\max}/2:(1+\varepsilon)^{i}\leq 2/\varepsilon\mbox{ and }i\in\mathbb{Z}_{\geq 0}\}$ $\triangleright$ distance thresholds
 8:  for threshold $d\in D$ do
@@
 11:     Update $S\leftarrow T$      return $S$
-1:function GreedyIndependentSet(points $V$, monotone submodular function $g:V\rightarrow\mathbb{R}_{\geq 0}$, distance $d$, budget $k$)
+1:function GreedyIndependentSet(points $V$, monotone submodular function $g:2^{V}\rightarrow\mathbb{R}_{\geq 0}$, distance $d$, budget $k$)
 2:  Initialize $S\leftarrow\varnothing$
@@
 5:   if $C=\varnothing$ then
-6:              return $S$triangleright $S$ is a maximal independent set of $G_{d}(V)$
+6:     return $S$ $\triangleright$ $S$ is a maximal independent set of $G_{d}(V)$
 7:   Find $t\leftarrow\operatorname*{arg\,max}_{v\in C}g(v\mid S)$
@@
 Theorem 3.1.
-For any $\varepsilon>0$, the GIST algorithm outputs a set $S\subseteq V$ with $|S|\leq k$ and $f(S)\geq(\nicefrac{{1}}{{2}}-\varepsilon)\cdot\textnormal{OPT}$.
-The main building block in our design and analysis of GIST is the following lemma, which is inspired by the greedy $2$-approximation algorithm for metric $k$-centers and adapted for monotone submodular rewards.
+For any $\varepsilon>0$, GIST outputs a set $S\subseteq V$ with $|S|\leq k$ and $f(S)\geq(\nicefrac{{1}}{{2}}-\varepsilon)\cdot\textnormal{OPT}$ using $O(nk\log_{1+\varepsilon}(1/\varepsilon))$ submodular value oracle queries.
+The main building block in our design and analysis of GIST is the following lemma, which is inspired by the greedy $2$-approximation algorithm for metric $k$-center and adapted for monotone submodular utility functions.
 Lemma 3.2.
@@
 $S_{d}^{*}=\operatorname*{arg\,max}_{S:|S|\leq k,\texttt{div}(S)\geq d}g(S).$
-[…]
+Let $T$ be the output of $\textnormal{{GreedyIndependentSet}}(V,g,d^{\prime},k)$. If $d^{\prime}<d/2$, then $g(T)\geq g(S_{d}^{*})/2$.
 Proof.
@@
 Since the distance between any pair of points in $B_{i}$ is at most $2d^{\prime}<d$, each set $B_{i}$ contains at most one point in $S_{d}^{*}$.
-We first construct an injective map $h:S_{d}^{*}\rightarrow T$.
-We say that a point $v\in V$ is covered by $B_{i}$ if $v\in B_{i}$.
+First we construct an injective map $h:S_{d}^{*}\rightarrow T$.
+We say that $B_{i}$ covers a point $v\in V$ if $v\in B_{i}$.
 For each covered $s\in S_{d}^{*}$, we map it to $t_{i}$, where $i$ is the minimum index for which $B_{i}$ covers $s$.
@@
 For any $s\in S_{d}^{*}$, let $T_{s}$ be the set of points added to set $T$ in algorithm GreedyIndependentSet right before the addition of $h(s)$.
-Since GreedyIndependentSet incrementally adds points and never deletes points from $T$, we know $T_{s}\subseteq T$.
+Since GreedyIndependentSet iteratively adds points and never deletes points from $T$, we know $T_{s}\subseteq T$.
 By submodularity, $g(s\mid T)\leq g(s\mid T_{s})$.
@@
 We note that the sum of the former terms, $g(h(s)\mid T_{s})$, is at most $g(T)$ since $h$ is injective and $g$ is nondecreasing. Thus, we conclude that
-$g(S_{d}^{*})-g(T)$ $\leq\sum_{s\in S_{d}^{*}}g(s\mid T)$
-$\leq\sum_{s\in S_{d}^{*}}g(h(s)\mid T_{s})$
-$\leq g(T),$
+$g(S_{d}^{*})-g(T)\leq\sum_{s\in S_{d}^{*}}g(s\mid T)\leq\sum_{s\in S_{d}^{*}}g(h(s)\mid T_{s})\leq g(T),$
 which completes the proof. ∎
-Equipped with this bicriteria approximation guarantee, we can now analyze the approximation ratio of GIST.
-Proof of Theorem 3.1.
+Now that we have with this bicriteria approximation, we can analyze the approximation ratio of GIST.
+Proof of Theorem˜3.1.
+We first prove the oracle complexity of GIST.
+There are $1+\log_{1+\varepsilon}(2/\varepsilon)=O(\log_{1+\varepsilon}(1/\varepsilon))$ thresholds in $D$.
+For each threshold, we call GreedyIndependentSet once.
+In each call for $k$ iterations, we find the maximum marginal-value point by scanning (in the worst case) all points.
+This requires at most $nk$ oracle calls yielding the overall oracle complexity of the algorithm.
 Let $d^{*}$ be the minimum pairwise distance between points in $S^{*}$.
 The GIST algorithm iterates over a set of thresholds $D$.
-The definition of $D$ implies that at least one threshold $d$ is in the range $[d^{*}/(2(1+\varepsilon)),d^{*}/2)$ unless $d^{*}\leq\varepsilon d_{\max}$.
+The definition of $D$ implies that at least one threshold $d$ is in the interval $d^{*}/(2(1+\varepsilon)),d^{*}/2)$, unless $d^{*}\leq\varepsilon d_{\max}$.
 We deal with this special case later and focus on the case when such a $d$ exists.
-Let $T$ be the output of GreedyIndependentSet for threshold value $d$. Applying Lemma 3.2 since $d<d^{*}/2$, we know that
+Let $T$ be the output of GreedyIndependentSet for threshold value $d$. Since $d<d^{*}/2$, Lemma˜3.2 implies that
 $g(T)\geq\frac{1}{2}\cdot g(S_{d^{*}}^{*})\geq\frac{1}{2}\cdot g(S^{*}).$
@@
 $f(T)\geq\frac{1}{2(1+\varepsilon)}\cdot\textnormal{OPT}>\left(\frac{1}{2}-\varepsilon\right)\cdot\textnormal{OPT}.$
-It remains to prove the claim for the special case of $d^{*}\leq\varepsilon d_{\max}$.
-Algorithm GIST considers two initial feasible sets and picks the better of the two as the initial value for $T$.
-The first set is the classic greedy solution for the submodular function $g$, ignoring the diversity term.
-This gives the lower bound (Nemhauser et al., 1978):
-$f(T)\geq\left(1-\frac{1}{e}\right)\cdot g(S^{*})\geq\left(1-\frac{1}{e}\right)\cdot\left(\textnormal{OPT}-\varepsilon d_{\max}\right).$
-The other initial set realizes the max pairwise distance $d_{\max}$ and ignores the submodular term, yielding the lower bound:
-$f(T)\geq d_{\max}.$
-Combining these two lower bounds, we get a final bound of
-$f(T)\geq\left(1-\frac{1}{e}\right)\cdot\frac{\textnormal{OPT}}{1+\varepsilon},$
-which completes the proof. ∎
-Remark 3.3.
-We can modify GIST to guarantee an exact $\nicefrac{{1}}{{2}}$-approximation ratio by trying all pairwise thresholds $D^{\prime}=\{\textnormal{dist}(u,v)/2-\varepsilon^{\prime}:u,v\in V\}$, for a small enough $\varepsilon^{\prime}>0$, at the cost of $O(n^{2})$ calls to GreedyIndependentSet instead of $O(\log_{1+\varepsilon}(1/\varepsilon))$.
+It remains to prove the claim for the case when $d^{*}\leq\varepsilon d_{\max}$.
+GIST considers two initial feasible sets and picks the better of the two as the initial value for $T$.
+The first set is the classic greedy solution [47] for the monotone submodular function $g(S)$, and ignores the diversity term.
+It follows that
+$f(T)\geq\left(1-\frac{1}{e}\right)\cdot g(S^{*})\geq\left(1-\frac{1}{e}\right)\cdot\left(\textnormal{OPT}-\lambda\cdot\varepsilon d_{\max}\right).$ (3)
+The second set contains two points with pairwise distance $d_{\max}$, and ignores the submodular term. This yields the lower bound
+$f(T)\geq\lambda\cdot d_{\max}\implies-\lambda\cdot d_{\max}\geq-f(T).$ (4)
+Combining (3) and (4), we get a final bound of $f(T)\geq(1-\frac{1}{e})\cdot\frac{\textnormal{OPT}}{1+\varepsilon},$ which completes the proof. ∎
 3.1 Linear utility
-We now assume $g(S)=\sum_{v\in S}w(v)$ is a linear function for some weights $w:V\rightarrow\mathbb{R}_{\geq 0}$.
-This is a special case of submodularity, so Theorem 3.1 gives a $(\nicefrac{{1}}{{2}}-\varepsilon)$-approximation.
-However, GIST achieves a stronger approximation ratio in this case.
+If $g(S)=\sum_{v\in S}w(v)$ is a linear utility with nonnegative weights $w:V\rightarrow\mathbb{R}_{\geq 0}$, then Theorem˜3.1 gives us a $(\nicefrac{{1}}{{2}}-\varepsilon)$-approximation since this is a special case of submodularity.
+However, GIST offers a stronger approximation ratio under this assumption.
 Theorem 3.3.
-Let $g(S)=\sum_{v\in S}w(v)$ be a linear function. For any $\varepsilon>0$, GIST outputs a set $S\subseteq V$ with $|S|\leq k$ and $f(S)\geq(\nicefrac{{2}}{{3}}-\varepsilon)\cdot\textnormal{OPT}$.
-We defer the proof to Appendix B, but describe the main differences.
-For linear functions, Lemma 3.2 can be strengthened to show that GreedyIndependentSet outputs a set $T$ such that $g(T)\geq g(S_{d}^{*})$, for any $d^{\prime}<d/2$.
-Then, for the right distance threshold $d\in D$, we get two lower bounds $\textnormal{ALG}\geq\max\left\{g(S^{*})+d^{*}/(2(1+\varepsilon)),d^{*}\right\}$ that can be balanced appropriately.
-Remark 3.5.
-GIST iterates over $O(\log_{1+\varepsilon}(1/\varepsilon))$ distance thresholds.
-For each $d$, it inserts up to $k$ points with $O(n)$ distance and submodular valuation oracle calls per insertion, so these iterations consist of $O(nk\log_{1+\varepsilon}(1/\varepsilon))$ distance and submodular oracle calls.
-For linear functions, we can sort the points in advance by $w(v)$ and iterate over this ordering in all GreedyIndependentSet calls to reduce the number of valuation probes to $O(n\log_{1+\varepsilon}(1/\varepsilon))$.
+Let $g(S)=\sum_{v\in S}w(v)$ be a linear function. For any $\varepsilon>0$, GIST returns $S\subseteq V$ with $|S|\leq k$ and $f(S)\geq(\nicefrac{{2}}{{3}}-\varepsilon)\cdot\textnormal{OPT}$.
+We defer the proof to Section˜C.1, but explain the main differences.
+For linear functions, Lemma˜3.2 can be strengthened to show that GreedyIndependentSet outputs a set $T$ such that $g(T)\geq g(S_{d}^{*})$, for any $d^{\prime}<d/2$.
+Then, for some $d\in D$, we get $\textnormal{ALG}\geq\max\left\{g(S^{*})+\lambda\cdot d^{*}/(2(1+\varepsilon)),\lambda\cdot d^{*}\right\}$, and the optimal convex combination of these two lower bounds gives the approximation ratio.
 4 Hardness of approximation
-[…]
-Submodular utility.
-There is no polynomial-time $0.5584$-approximation algorithm if $g$ is submodular.
+We begin by summarizing our hardness results for MDMS. Assuming $\text{P}\neq\text{NP}$, we prove that:
+Submodular utility. There is no polynomial-time $0.5584$-approximation algorithm if $g$ is a nonnegative, monotone submodular function.
 Linear utility.
-  • There is no polynomial-time $(\nicefrac{{2}}{{3}}+\varepsilon)$-approximation algorithm if $g$ is linear for general metrics, for any $\varepsilon>0$.
+  • There is no polynomial-time $(\nicefrac{{2}}{{3}}+\varepsilon)$-approximation algorithm if $g$ is linear, for any $\varepsilon>0$, for general distance metrics.
   • APX-completeness for linear utility functions even in the Euclidean metric, i.e., there is no polynomial-time approximation scheme (PTAS) for this problem.
 4.1 General metric spaces
-We build on the set cover hardness instances originally designed by Feige (1998); Feige et al.
-(2004); Feige and Vondrák (2010), and further engineered by Kapralov et al.
-(2013).
-In particular, we present the instances designed in Kapralov et al.
-(2013, Section 2) as follows:
+We build on the set cover hardness instances originally designed by Feige et al.
+[23, 25, 24], and then further engineered by Kapralov et al.
+[32].
+We present the instances in [32, Section 2] as follows:
 Hardness of max $k$-cover.
@@
 The construction above can be done such that each set $S_{i}\in\mathcal{F}=\bigcup_{r=1}^{k}\mathcal{F}_{r}$ has the same size.
-We also note that the parameter $k$ should be larger than any constant we set since there is always an exhaustive search algorithm with running time $|\mathcal{F}|^{k}$ to distinguish between the YES and NO cases.
+We note that the parameter $k$ should be larger than any constant we set since there is always an exhaustive search algorithm with running time $|\mathcal{F}|^{k}$ to distinguish between the YES and NO cases.
 Therefore, we can assume, e.g., that $k\geq 1/\varepsilon$.
@@
 Theorem 4.1.
-For any $\delta>0$, it is NP-hard to approximate MDMS within a factor of $\frac{2(1-1/e)}{2(1-1/e)+1}+\delta<0.55836+\delta$.
+It is NP-hard to approximate MDMS within a factor of $\frac{2(1-1/e)}{2(1-1/e)+1}+\delta<0.55836+\delta$, for any $\delta>0$.
+Proof sketch.
+We reduce an instance of the max $k$-cover problem to MDMS to demonstrate how the inapproximability result translates.
+For any pair of sets in the collection $\mathcal{F}$, their distance is defined to be $2$ if and only if they are disjoint; otherwise, it is $1$.
+Consequently, a selected sub-collection of $\mathcal{F}$ achieves the higher diversity score of $2$ if and only if all chosen sets are mutually disjoint.
+Given that all sets in $\mathcal{F}$ are of uniform size and considering the upper bound in the NO case on the union coverage for any collection of sets, a polynomial-time algorithm can only find $O(\varepsilon k)$ pairwise disjoint sets.
+Thus, if an algorithm aims for a diversity score of $2$ (by selecting only disjoint sets), it forgoes most of the potential value from the submodular coverage function in the MDMS formulation.
+On the other hand, if the algorithm settles for the lower diversity score of $1$, it still cannot cover more than an approximate fraction $1-(1-\nicefrac{{1}}{{k}})^{k}+\varepsilon\approx 1-\nicefrac{{1}}{{e}}+\varepsilon$ of the universe.
+Combining these two upper bounds on any polynomial-time algorithm’s performance with the fact that, in the YES case, the optimal solution covers the entire universe and achieves diversity score $2$, establishes the claimed hardness of approximation.
+∎
+4.2 Linear utility
+For the special case of linear utility functions, we prove a tight hardness result to complement our $\nicefrac{{2}}{{3}}$-approximation guarantee in Theorem˜3.3.
+Theorem 4.2.
+For any $\varepsilon>0$, there is no polynomial-time $(\nicefrac{{2}}{{3}}+\varepsilon)$-approximation algorithm for the MDMS problem if $g$ is a linear function, unless $\textnormal{P}=\textnormal{NP}$.
+We defer the proof to Section˜D.2.
+Our construction builds on the work of Håstad [28] and Zuckerman [64], which shows that the max clique problem does not admit an efficient $n^{1-\theta}$-approximation algorithm, for any constant $\theta>0$.
+4.3 Euclidean metric
+Our final result is for the Euclidean metric, i.e., $S\subseteq\mathbb{R}^{d}$ and $\textnormal{dist}(\mathbf{x},\mathbf{y})=\lVert\mathbf{x}-\mathbf{y}\rVert_{2}$.
+We build on a result of Alimonti and Kann [4] showing that the size of a max independent set in a bounded-degree graph cannot be approximated to within a constant factor $1-\varepsilon$, for any $\varepsilon>0$, unless $\text{P}=\text{NP}$.
+Lemma 4.3 (Alimonti and Kann [4, Theorem 3.2]).
+The maximum independent set problem for graphs with degree at most $3$ is APX-complete.
+Our second ingredient is an embedding function $h_{G}(v)$ that encodes graph adjacency in Euclidean space.
+Lemma 4.4.
+Let $G=(V,E)$ be a simple undirected graph with $n=|V|$, $m=|E|$, and max degree $\Delta$. There exists an embedding $h_{G}:V\rightarrow\mathbb{R}^{n+m}$ such that if $\{u,v\}\in E$ then
+$\lVert h_{G}(u)-h_{G}(v)\rVert_{2}\leq 1-\frac{1}{2(\Delta+1)},$
+and if $\{u,v\}\not\in E$ then $\lVert h_{G}(u)-h_{G}(v)\rVert_{2}=1$.
+Theorem 4.5.
+MDMS is APX-complete for the Euclidean metric if $g$ is a linear function.
+We sketch the main idea below and defer proofs to Appendix˜D.
+Let $\mathcal{I}(G)$ be the set of independent sets of $G$.
+Using Lemma˜4.4, for any set of nodes $S\subseteq V$ with $|S|\geq 2$ in a graph $G$ with max degree $\Delta\leq 3$, we have the property:
+  • $S\in\mathcal{I}(G)\implies\texttt{div}(S)=1$;
+  • $S\not\in\mathcal{I}(G)\implies\texttt{div}(S)\leq 1-\frac{1}{2(\Delta+1)}\leq 1-\nicefrac{{1}}{{8}}$.
+The gap is at least $\nicefrac{{1}}{{8}}$ for all such graphs (i.e., a universal constant), so setting $\lambda=1$ allows us to prove that MDMS inherits the APX-completeness of bounded-degree max independent set.
+5 Experiments
+5.1 Warm-up: Synthetic dataset
+We first compare GIST against baseline methods on a simple MDMS task with a normalized budget-additive utility function (i.e., monotone submodular) and a set of weighted Guassian points.
+Setup.
+Generate $n=1000$ points $\mathbf{x}_{i}\in\mathbb{R}^{d}$, for $d=64$, where each $\mathbf{x}_{i}\sim\mathcal{N}(\mathbf{0},\mathbf{I}_{d})$ is i.i.d.
+and assigned a uniform continuous weight $w_{i}\sim\mathcal{U}_{[0,1]}$.
+For $\alpha,\beta\in[0,1]$, the objective function trades off between the average (capped) utility of the points and their min-distance diversity reward:
+$f(S)=\alpha\cdot\min\left\{\frac{1}{k}\sum_{i\in S}w_{i},\,\beta\right\}+(1-\alpha)\cdot\texttt{div}(S).$
+We consider three baseline methods: random, simple, and greedy.
+random selects $k$ random points, permutes them, and returns the best prefix since the objective function $f(S)$ is non-monotone.
+simple is the $0.387$-approximation algorithm in Section˜2.1.
+greedy builds $S$ one point at a time by selecting the point with index $i^{*}_{t}=\operatorname*{arg\,max}_{i\in V\setminus S_{t-1}}f(S_{t-1}\cup\{i\})-f(S_{t-1})$ at each step $t\in[k]$, and returns the best prefix $S=\operatorname*{arg\,max}_{t\in[k]}f(S_{t})$.
+GIST considers all possible $D\leftarrow\{\textnormal{dist}(u,v)/2:u,v\in V\}$ since $n=1000$, which yields an exact $\nicefrac{{2}}{{3}}$-approximation ratio.
+Results.
+Figure 1: $f(S_{\text{ALG}})$ for baseline methods and GIST, for each cardinality constraint $k\in[n]$, on synthetic data with $n=1000$, $\alpha=0.95$, and $\beta=0.75$.
+We plot the values of $f(S_{\text{ALG}})$ for the baseline methods and GIST, for each $k\in[n]$, in Figure˜1.
+GIST dominates simple, greedy, and random in all cases.
+greedy performs poorly for $k\geq 100$, which is surprising since it is normally a competitive method for subset selection tasks.
+simple beats greedy for mid-range values of $k$, i.e., $k\in[250,900]$, but it is always worse than GIST.
+In Section˜E.1, we sweep over $\alpha$ and $\beta$, and show how these hyperparameters affect the $f(S_{\text{ALG}})$ plots.
+Finally, we remark that the plots in Figure˜1 are decreasing in $k$ because (i) we use a normalized budget-additive utility function, and (ii) $\texttt{div}(S)$ is a monotone-decreasing set function.
+5.2 Image classification
+Our ImageNet data sampling experiment compares the top-1 image classification accuracy achieved by different single-shot subset selection algorithms.
+Setup.
+We use the standard vision dataset ImageNet [54] containing ~1.3 million images and 1000 classes.
+We select 10% of the images uniformly at random and use them to train an initial ResNet-56 model $\bm{\theta}_{0}$ [29].
+Then we use the model $\bm{\theta}_{0}$ to compute a 2048-dimensional unit-length embedding $\bm{e}_{i}$ and uncertainty score for each example $\bm{x}_{i}$.
+The margin-based uncertainty score of $\mathbf{x}_{i}$ is given by $u_{i}=1-(\Pr(y=b\mid\mathbf{x}_{i};~\bm{\theta}_{0})-\Pr(y=b^{\prime}\mid\mathbf{x}_{i};~\bm{\theta}_{0}))$, which measures the difference between the probability of the best predicted class label $b$ and second-best label $b^{\prime}$ for an example.
+Finally, we use the fast maximum inner product search of Guo et al.
+[27] to build a $\Delta$-nearest neighbor graph $G$ in the embedding space using $\Delta=100$ and cosine distance, i.e., $\textnormal{dist}(\bm{x}_{i},\bm{x}_{j})=1-\bm{e}_{i}\cdot\bm{e}_{j}$.
+We present all model training hyperparameters in Appendix˜E.
+We compare GIST with several state-of-the-art benchmarks:
+  • random: We draw samples from the dataset uniformly at random without replacement. This is a simple and lightweight approach that promotes diversity in many settings and provides good solutions.
+  • margin [53]: Margin sampling selects the top-$k$ points using the uncertainty scores $u_{i}$, i.e., based on how hard they are to classify.
+It is not incentivized to output a diverse set of training examples.
+  • $k$-center [57]: We run the classic greedy algorithm for $k$-center on $G$. We take the distance between non-adjacent nodes to be the max distance among all pairs of adjacent nodes in $G$.
+  • submod: We select a subset by greedily maximizing the submodular objective function
+    $g(S)=\alpha_{s}\sum_{i\in S}u_{i}-\beta_{s}\sum_{i,j\in S}s(i,j),$ (5)
+    subject to the constraint $|S|\leq k$, where $s(i,j)=1-\textnormal{dist}(\bm{x}_{i},\bm{x}_{j})$ is the cosine similarity between adjacent nodes in $G$.
+Similar pairwise-diversity submodular objective functions have also been used in [45, 21, 31, 36, 49].
+This is a different diversity objective than $\texttt{div}(S)$ that keeps $g(S)$ submodular but allows it to be non-monotone.
+We tuned for best performance by selecting $\alpha_{s}=0.9$ and $\beta_{s}=0.1$.
+Finally, we bootstrap the margin and submod objectives with MDMS and run GIST with $\varepsilon=0.05$:
+  • GIST-margin: Let $f(S)=\alpha\cdot\sum_{i\in S}u_{i}+(1-\alpha)\cdot\texttt{div}(S)$ for $\alpha\in[0,1]$.
+This uses the same linear utility function as margin sampling.
+We optimize for performance and set $\alpha=0.9$.
+  • GIST-submod: Let $f(S)=\alpha\cdot g(S)+(1-\alpha)\cdot\texttt{div}(S)$, where $g(S)$ is the same submodular function in (5) with $\alpha_{s}=0.9$ and $\beta_{s}=0.1$.
+We optimize for performance and set $\alpha=0.95$.
+Table 1: Top-1 classification accuracy (%) on ImageNet for different single-shot data downsampling algorithms.
+The cardinality constraint $k$ is expressed as a percent of the ~1.3 million examples.
+The results are the average of three trials and the top performance for each $k$ is shown in bold.
+$k$ (%) random margin $k$-center submod GIST-margin GIST-submod
+30 66.23 $\pm$ 0.15 65.97 $\pm$ 0.33 66.71 $\pm$ 0.57 66.52 $\pm$ 0.21 66.90 $\pm$ 0.19 67.24 $\pm$ 0.06
+40 69.17 $\pm$ 0.12 69.73 $\pm$ 0.38 70.06 $\pm$ 0.06 70.71 $\pm$ 0.24 70.51 $\pm$ 0.12 70.76 $\pm$ 0.35
+50 71.05 $\pm$ 0.16 72.33 $\pm$ 0.09 73.01 $\pm$ 0.10 72.69 $\pm$ 0.15 72.69 $\pm$ 0.34 73.15 $\pm$ 0.21
+60 72.49 $\pm$ 0.28 73.43 $\pm$ 0.06 73.60 $\pm$ 0.31 74.20 $\pm$ 0.08 74.34 $\pm$ 0.01 74.30 $\pm$ 0.27
+70 73.70 $\pm$ 0.22 74.49 $\pm$ 0.22 74.24 $\pm$ 0.23 75.24 $\pm$ 0.30 75.41 $\pm$ 0.23 75.32 $\pm$ 0.19
+80 74.42 $\pm$ 0.39 75.01 $\pm$ 0.05 75.17 $\pm$ 0.11 75.91 $\pm$ 0.07 75.96 $\pm$ 0.28 75.45 $\pm$ 0.16
+90 75.16 $\pm$ 0.13 75.11 $\pm$ 0.13 75.00 $\pm$ 0.38 75.84 $\pm$ 0.10 76.12 $\pm$ 0.33 76.03 $\pm$ 0.02
+Results.
+We run each sampling algorithm with cardinality constraint $k$ on the full dataset to get a subset of examples that we then use to train a new ResNet-56 model.
+We report the average top-1 classification accuracy of these models in Table˜1.
+GIST with margin or submodular utility is superior to all baselines.
+Interestingly, there is a cut-over value of $k$ where the best algorithm switches from GIST-submod to GIST-margin.
+We also observe that GIST-submod and GIST-margin outperform submod and margin, respectively.
+This demonstrates how $\texttt{div}(S)$ encourages diversity in the set of sampled points and improves downstream model quality.
+The running time of margin and submod is 3–4 minutes per run on average.
+GIST is similar to the margin or submodular algorithms for a given distance threshold $d$.
+The end-to-end running time is dominated by training ImageNet models, which takes more than a few hours even with several accelerators (e.g., GPU/TPU chips).
+Conclusion
+We introduce a novel subset selection problem called MDMS that combines the utility of the selected points (modeled as a monotone submodular function $g$) with the $\texttt{div}(S)=\min_{u,v\in S:u\neq v}\textnormal{dist}(u,v)$ diversity objective.
+We design and analyze the GIST algorithm, which achieves a $\nicefrac{{1}}{{2}}$-approximation guarantee by solving a series of maximal-weight independent set instances on intersection graphs with the GreedyIndependentSet bicriteria-approximation algorithm.
+We complement GIST with a $0.5584$ hardness of approximation.
+It is an interesting open theory problem to close the gap between the $0.5$ approximation ratio and $0.5584$ inapproximability.
+For linear utilities, we show that GIST achieves a $\nicefrac{{2}}{{3}}$-approximation and that it is NP-hard to find a $(\nicefrac{{2}}{{3}}+\varepsilon)$-approximation, for any $\varepsilon>0$.
+Our empirical study starts by comparing GIST to existing methods for MDMS on a simple synthetic task to show the shortcomings of baseline methods (in particular the greedy algorithm).
+Then we compare the top-1 image classification accuracy of GIST and state-of-the-art data sampling methods for ImageNet, demonstrating the benefit of optimizing the trade-off between a submodular utility and max-min diversity.
+Appendix A $f(S)$ is not submodular
+Consider the instance on four points $N=\{a,b,c,d\}\subseteq\mathbb{R}$ where $a=0$, $b=1$, and $c=d=2$ with the (one-dimensional) Euclidean metric, i.e., all the points are collinear.
+Let $g(S)=0$ for all $S\subseteq V$.
+For $f(S)$ to be submodular, it must hold that for every $S\subseteq T\subseteq N$ and $x\in N\setminus T$,
+$f(S\cup\{x\})-f(S)\geq f(T\cup\{x\})-f(T).$
+However, if $x=b$, $S=\{a,c\}$ and $T=\{a,c,d\}$, we have:
+$f(S\cup\{x\})-f(S)$ $=f(\{a,b,c\})-f(\{a,c\})=1-2=-1$
+$f(T\cup\{x\})-f(T)$ $=f(\{a,b,c,d\})-f(\{a,c,d\})=0-0=0,$
+so $f(S)$ is not submodular.
+In this instance, $f(S)$ is not monotone. However, a similar monotone but still not submodular $f(S)$ can be defined by setting $g(S)=\sum_{v\in S}w(v)$ where $w(v)=2$ for every $v\in N$.
+Appendix B Greedy does not give a constant-factor approximation guarantee
+The objective function $f$ is highly non-monotone since the diversity term can decrease as we add items.
+Consequently, the standard greedy algorithm, which rejects any item with a negative marginal gain, can have an arbitrarily poor performance.
+We demonstrate this with a hard instance.
+Let the submodular part of the objective be $g(S)=|S|$.
+For the diversity component, define the distance between a specific pair $(u,v)$ to be $\text{dist}(u,v)=2+2\varepsilon$, while for all other pairs $(x,y)\neq(u,v)$, we have $\text{dist}(x,y)=1+\varepsilon$.
+The greedy algorithm first selects the set $\{u,v\}$, achieving a value of $f(\{u,v\})=g(\{u,v\})+\text{dist}(u,v)=4+2\varepsilon$.
+The algorithm then terminates because adding any subsequent item gives a submodular gain of 1 but causes a diversity loss of $1+\varepsilon$, resulting in a negative marginal gain.
+An optimal solution of size $k$, however, can achieve a value of at least $k$ from the submodular term alone.
+The resulting approximation ratio for greedy is at most $(4+2\varepsilon)/k$, which approaches 0 as $k$ grows.
+Thus, greedy offers no constant-factor approximation guarantee.
+Furthermore, modifying greedy to accept items with negative marginal value does not solve this problem since one can construct instances where an optimal solution value is dominated by a diversity term such that selecting any set of size $k$ reduces the diversity term to zero.
+Appendix C Missing analysis for Section 3
+C.1 Proof of Theorem 3.3
+The following result is a tighter analysis of the bicriteria approximation of GreedyIndependentSet (Lemma˜3.2) if $g(S)$ is a linear function.
+This is the key ingredient for improving the approximation ratio of GIST to $\nicefrac{{2}}{{3}}-\varepsilon$.
+Lemma C.1.
+Let $g(S)=\sum_{v\in S}w(v)$ be a linear function with nonnegative weights $w:V\rightarrow\mathbb{R}_{\geq 0}$.
+Let $S_{d}^{*}$ be a max-weight independent set of the intersection graph $G_{d}(V)$ of size at most $k$.
+If $T$ is the output of $\textnormal{{GreedyIndependentSet}}(V,g,d^{\prime},k)$, for $d^{\prime}\leq d/2$, then $w(T)\geq w(S_{d}^{*})$.
+Proof.
+Let $k^{\prime}=|T|$ and $t_{1},t_{2},\dots,t_{k^{\prime}}$ be the points in $T$ in the order that GreedyIndependentSet selected them.
+Let $B_{i}=\{v\in V:\textnormal{dist}(t_{i},v)<d^{\prime}\}$ be the points in $V$ contained in the radius-$d^{\prime}$ open ball around $t_{i}$.
+First, we show that each $B_{i}$ contains at most one point in $S_{d}^{*}$.
+If this is not true, then some $B_{i}$ contains two different points $u,v\in S_{d}^{*}$.
+Since $\textnormal{dist}(\cdot,\cdot)$ is a metric, this means
+$\textnormal{dist}(u,v)$ $\leq\textnormal{dist}(u,t_{i})+\textnormal{dist}(t_{i},v)$
+$<d^{\prime}+d^{\prime}$
+$\leq d/2+d/2$
+$=d,$
+which contradicts the assumption that $S_{d}^{*}$ is an independent set of $G_{d}(V)$.
+Note that it is possible to have $B_{i}\cap B_{j}\neq\varnothing$, for $i\neq j$, since these balls consider all points in $V$.
+Now let $C_{i}\subseteq V$ be the set of uncovered points (by the open balls) that become covered when GreedyIndependentSet selects $t_{i}$.
+Concretely, $C_{i}=B_{i}\setminus(B_{1}\cup\cdots\cup B_{i-1})$.
+Each $C_{i}$ contains at most one point in $S_{d}^{*}$ since $|B_{i}\cap S_{d}^{*}|\leq 1$.
+Moreover, if $s\in C_{i}\cap S_{d}^{*}$, then $w(t_{i})\geq w(s)$ because the points are sorted in non-increasing order and selected if uncovered.
+Let $A=C_{1}\cup\cdots\cup C_{k^{\prime}}$ be the set of points covered by the algorithm.
+For each point $s\in S_{d}^{*}\cap A$, there is exactly one covering set $C_{i}$ corresponding to $s$.
+It follows that
+$\sum_{s\in S_{d}^{*}\cap A}w(s)\leq\sum_{i\in[k^{\prime}]:S_{d}^{*}\cap C_{i}\neq\varnothing}w(t_{i}).$ (6)
+It remains to account for the points in $S_{d}^{*}\setminus A$.
+If we have any such points, then $|T|=k$ since the points in $S_{d}^{*}\setminus A$ are uncovered at the end of the algorithm.
+Further, for any $t_{i}\in T$ and $s\in S_{d}^{*}\setminus A$, we have $w(t_{i})\geq w(s)$ since $t_{i}$ was selected and the points are sorted by non-increasing weight.
+Therefore, we can assign each $s\in S_{d}^{*}\setminus A$ to a unique $C_{i}$ such that $C_{i}\cap S_{d}^{*}=\varnothing$.
+It follows that
+$\sum_{s\in S_{d}^{*}\setminus A}w(s)\leq\sum_{i\in[k^{\prime}]:S_{d}^{*}\cap C_{i}=\varnothing}w(t_{i}).$ (7)
+Adding the two sums together in (6) and (7) completes the proof. ∎
+See 3.3
+Proof of Theorem˜3.3.
+Let $d^{*}$ be the minimum distance between two distinct points in $S^{*}$.
+There are two cases: $d^{*}\leq\varepsilon d_{\max}$ and $d^{*}>\varepsilon d_{\max}$.
+In the first case, outputting the $k$ heaviest points (Line 2 of GIST) yields a $(1-\varepsilon)$-approximation.
+To see this, first observe that
+$\textnormal{OPT}\geq\lambda\cdot d_{\max}\geq\lambda\cdot\frac{d^{*}}{\varepsilon}\implies\varepsilon\cdot\textnormal{OPT}\geq\lambda\cdot d^{*}.$
+The sum of the $k$ heaviest points upper bounds $g(S^{*})$, so we have
+$\textnormal{ALG}\geq g(S^{*})=\textnormal{OPT}-\lambda\cdot d^{*}\geq(1-\varepsilon)\cdot\textnormal{OPT}.$
+Now we consider the case where $d^{*}>\varepsilon d_{\max}$.
+GIST tries a threshold $d\in[d^{*}/(2(1+\varepsilon)),d^{*}/2)$, so Lemma˜C.1 implies that $\textnormal{{GreedyIndependentSet}}(V,g,d,k)$ outputs a set $T$ such that
+$f(T)$ $\geq g(S^{*})+\lambda\cdot d\geq g(S^{*})+\lambda\cdot\frac{d^{*}}{2(1+\varepsilon)}.$ (8)
+The max-diameter check on Lines 3–6 give us the lower bound
+$\textnormal{ALG}\geq\lambda\cdot d_{\max}\geq\lambda\cdot d^{*}.$ (9)
+Combining (8) and (9), the following inequality holds for any $0\leq p\leq 1$:
+ALG $\geq p\cdot\left[g(S^{*})+\lambda\cdot\frac{d^{*}}{2(1+\varepsilon)}\right]+(1-p)\cdot\lambda\cdot d^{*}$
+$=p\cdot g(S^{*})+\left(1-p+\frac{p}{2(1+\varepsilon)}\right)\cdot\lambda\cdot d^{*}.$
+To maximize the approximation ratio as $\varepsilon\rightarrow 0$, we set $p=2/3$ by solving p=1p/2p=1-p/2. Therefore,
+ALG $\geq\frac{2}{3}\cdot g(S^{*})+\left(1-\frac{2}{3}+\frac{1}{3(1+\varepsilon)}\right)\cdot\lambda\cdot d^{*}$ (10)
+$=\frac{2}{3}\cdot g(S^{*})+\frac{1}{3}\left(1+\frac{1}{1+\varepsilon}\right)\cdot\lambda\cdot d^{*}$
+$\geq\frac{2}{3}\cdot g(S^{*})+\frac{1}{3}\left(2-\varepsilon\right)\cdot\lambda\cdot d^{*}$
+$\geq\left(\frac{2}{3}-\varepsilon\right)\cdot\textnormal{OPT},$
+which completes the proof. ∎
+Appendix D Missing analysis for Section 4
+D.1 Proof of Theorem 4.1
+See 4.1
 Proof.
@@
 We set $d$ to be $(1-\nicefrac{{1}}{{e}})U$ where $U$ is the cardinality of the union of all sets, i.e., $U=|\bigcup_{S\in\mathcal{F}}S|$.
+Finally, we set $\lambda=1$ to complete the construction of the MDMS instance.
 In the YES case, the optimum solution is the family of $k$ disjoint sets from different groups that cover the entire universe.
@@
 $\left(1-\frac{1}{k}\right)^{k}\geq\frac{1}{e}-\frac{1}{2e(k-1)}.$
-Since $k-1$ is at least $\nicefrac{{1}}{{\varepsilon}}$, the submodular value in case (a) does not exceed:
+Since k1k-1 is at least $\nicefrac{{1}}{{\varepsilon}}$, the submodular value in case (a) does not exceed:
 $\left(1-\frac{1}{e}+\frac{1}{2e(k-1)}+\varepsilon\right)U\leq\left(1-\frac{1}{e}+2\varepsilon\right)U.$
@@
 This concludes the proof in both cases (a) and (b). ∎
-4.2 Linear Utility
-For the special case of linear utility functions, we prove a tight hardness-of-approximation result complementing the $\nicefrac{{2}}{{3}}$-approximation guarantee in Theorem 3.4.
-Theorem 4.2.
-[…]
-We give the proof in Section C.1.
-Our construction builds on the work of Håstad (1996); Zuckerman (2006), which shows that the maximum clique problem does not admit an efficient $n^{1-\theta}$-approximation algorithm, for any constant $\theta>0$.
-4.3 Euclidean Metric
-[…]
-Lemma 4.3 (Alimonti and Kann (2000, Theorem 3.2)).
-The maximum independent set problem for graphs with degree at most $3$ is APX-complete.
-Our second ingredient is a node embedding function $h_{G}(v)$ that encodes graph adjacency in Euclidean space.
-Lemma 4.4.
-Let $G=(V,E)$ be a simple undirected graph with $n=|V|$, $m=|E|$, and max degree $\Delta$. There exists an embedding $h_{G}:V\rightarrow\mathbb{R}^{n+m}$ such that if $\{u,v\}\in E$ then
-$\lVert h_{G}(u)-h_{G}(v)\rVert_{2}\leq 1-\frac{1}{2(\Delta+1)},$
-and if $\{u,v\}\not\in E$ then $\lVert h_{G}(u)-h_{G}(v)\rVert_{2}=1$.
-Theorem 4.5.
-The MDMS problem is APX-complete for the Euclidean metric even if $g$ is a linear function.
-We sketch the main idea below and defer complete proofs to Appendix C.
-Let $\mathcal{I}(G)$ be the set of independent sets of $G$.
-Using Lemma 4.4, for any set of nodes $S\subseteq V$ with $|S|\geq 2$ in a graph $G$ with max degree $\Delta\leq 3$, we have the property:
-  • $S\in\mathcal{I}(G)\implies\texttt{div}(S)=1$;
-  • $S\not\in\mathcal{I}(G)\implies\texttt{div}(S)\leq 1-\frac{1}{2(\Delta+1)}\leq 1-\nicefrac{{1}}{{8}}$.
-The gap between the cases is at least $\nicefrac{{1}}{{8}}$ for all such graphs (i.e., a universal constant), allowing us to prove that MDMS inherits the APX-completeness of bounded-degree max independent set.
-5 Experiments
-5.1 Synthetic Dataset
-We use a collection of weighted random points and a budget-additive (submodular) utility function to (1) compare against baselines and (2) highlight the non-monotone behavior.
-Setup.
-Generate $n=1000$ points $\mathbf{x}_{i}\in\mathbb{R}^{d}$, for $d=64$, where each $\mathbf{x}_{i}\sim\mathcal{N}(\mathbf{0},\mathbf{I}_{d})$ is i.i.d.
-and assigned a uniform continuous weight $w_{i}\sim\mathcal{U}_{[0,1]}$.
-For $\alpha,\beta\in[0,1]$, the objective function below trades off between the average (capped) utility of the selected points and their min-distance diversity:
-$f(S)=\alpha\cdot\min\left\{\frac{1}{k}\sum_{i\in S}w_{i},\beta\right\}+(1-\alpha)\cdot\texttt{div}(S).$
-We consider three baselines: random, simple, and greedy.
-random chooses a random set of $k$ points, permutes them, and returns the best prefix since the objective function is non-monotone.
-simple is the $0.387$-approximation algorithm in Section 2.1.
-greedy builds $S$ one point at a time by selecting the point with index $i^{*}_{t}=\operatorname*{arg\,max}_{i\in V\setminus S_{t-1}}f(S_{t-1}\cup\{i\})-f(S_{t-1})$ at each step $t\in[k]$, and outputs the best prefix $S=\operatorname*{arg\,max}_{t\in[k]}f(S_{t})$.
-Results.
-We plot the values of $f(S)$ for the baseline methods and GIST for all $k\in[n]$ in Figure 1.
-We set $\alpha=0.95$ to roughly balance the submodular and diversity terms.
-[^3] If $\mathbf{x},\mathbf{y}\sim\mathcal{N}(\mathbf{0},\mathbf{I}_{d})$ are i.i.d., then $\mathbb{E}[\lVert\mathbf{x}-\mathbf{y}\rVert_{2}^{2}]=2d$.
-Setting $\frac{\alpha}{2}\approx(1-\alpha)\sqrt{2d}\implies\alpha\approx 1-\frac{1}{2\sqrt{2d}+1}=0.957$, for $\beta\geq 0.5$.
-We sweep over $\alpha$ and budget $\beta$ in Section D.1, and plot how these hyperparameters change $f(S)$.
-Note that the objective values are not directly comparable as a function of $k$ because of the $\nicefrac{{1}}{{k}}$ normalizing factor (hence the discontinuities), but both terms in the objective are decreasing on average in $k$ for a fixed $\alpha$.
-GIST dominates simple, greedy, and random in all cases.
-The poor performance of greedy is somewhat surprising since it is often competitive for subset selection tasks.
-simple beats greedy for all but the extreme values of $k$, i.e., small or close to $n$, but is always worse than GIST.
-Figure 1: Baseline comparison plotting $f(S_{\text{ALG}})$ for $k\in[n]$ on synthetic data with $n=1000$, $\alpha=0.95$, and $\beta=0.75$.
-5.2 Image Classification
-Our ImageNet data sampling experiment compares the top-1 image classification accuracy achieved by different single-shot subset selection algorithms.
-Setup.
-We use the standard vision dataset ImageNet (Russakovsky et al., 2015) containing ~1.3 million images and 1000 classes.
-We select 10% of the points at random and use them to train an initial ResNet-56 model $\bm{\theta}_{0}$ (He et al., 2016).
-Then we use this initial model to compute an uncertainty value and 2048-dimensional embedding for each example.
-The margin-based uncertainty value of $\mathbf{x}_{i}$ is given by $u_{i}=1-(\Pr(y=b\mid\mathbf{x}_{i};~{}\bm{\theta}_{0})-\Pr(y=b^{\prime}\mid\mathbf{x}_{i};~{}\bm{\theta}_{0}))$, which measure the difference between the probability of the best and second-best class label $b^{\prime}$ for an example.
-Finally, we use the fast maximum inner product search of Guo et al.
-(2020) to build a $\Delta$-nearest neighbor graph in the embedding space using $\Delta=10$ and cosine distance.
-We present all model training hyperparameters in Appendix D.
-We compare GIST with several state-of-the-art benchmarks:
-  • random: We select samples from the dataset uniformly at random (without replacement).
-While this approach is simple and lightweight, it creates diversity in many settings and provides good solutions.
-  • margin (Roth and Small, 2006): Margin sampling selects the top-$k$ points based on how hard they are to classify, i.e., using margin-based uncertainty values $u_{i}$.
-  • $k$-center (Sener and Savarese, 2018): We run the classic greedy algorithm for $k$-center on the $\Delta$-nearest neighbor graph.
-The distance between non-adjacent nodes is implicitly assumed to be very large, which is reasonable in the large budget regime $k=\Omega(n)$.
-  • submod: We select a subset by maximizing the submodular objective function $\alpha_{s}\sum_{i\in S}u_{i}-\beta_{s}\sum_{i,j\in S}s(i,j)$, under the budget constraint $|S|\leq k$.
-Similar pairwise submodular objective functions have been used in (Iyer et al., 2021; Kothawade et al., 2021; Ramalingam et al., 2021).
-Here $s(i,j)$ denotes the similarity between adjacent nodes using the same graph as in $k$-center.
-We tuned for best performance by selecting $\alpha_{s}=0.9$ and $\beta_{s}=0.1$.
-We evaluate GIST using both linear and submodular utility functions.
-Let $f(S)=\alpha\cdot g(S)+(1-\alpha)\cdot\texttt{div}(S)$ for $\alpha\in[0,1]$.
-In Table 1, GIST-linear uses the same linear utility as margin: $g(S)=\sum_{i\in S}u_{i}$ with $\alpha=0.9$.
-GIST uses the same submodular utility as submod: $g(S)=\alpha_{s}\sum_{i\in S}u_{i}-\beta_{s}\sum_{i,j\in S}s(i,j)$ with $\alpha=0.95$, $\alpha_{s}=0.9$, and $\beta_{s}=0.1$.
-We use the same graph as $k$-center for distances.
-Table 1: Top-1 classification accuracy (%) on ImageNet for different single-shot data downsampling algorithms.
-The cardinality constraint $k$ is expressed as a fraction of the ~1.3 million examples.
-The results are the average of 3 trials and the top performance for each $k$ is shown in bold.
-$k$ (of data) random margin $k$-center submod GIST-linear GIST
-30 66.27 65.97 66.70 66.52 66.89 67.23
-40 69.17 69.73 70.06 70.71 70.51 70.75
-50 71.05 72.33 73.09 72.69 72.67 73.14
-60 72.49 73.42 73.60 74.20 74.37 74.30
-70 73.70 74.49 74.24 75.24 75.41 75.32
-80 74.42 75.01 75.17 75.91 75.96 75.45
-90 75.16 75.10 74.97 75.84 76.12 76.03
-Results.
-We run each sampling algorithm with cardinality constraint $k$ on the full dataset to get a subset of examples that we then use to train a new ResNet-56 model.
-We report the top-1 classification accuracy of these models in Table 1.
-GIST with a linear or submodular utility is superior to all baselines.
-Interestingly, there is a cutover value of $k$ where GIST-linear becomes the best algorithm.
-Finally, observe that GIST-linear outperforms margin in all cases.
-This demonstrates how $\texttt{div}(S)$ encourages diversity in the subset of sampled points and ultimately improves downstream model quality.
-Conclusion
-This work presents a novel subset selection problem called MDMS that combines the total utility of the selected points (modeled by a monotone submodular function $g$) with the $\texttt{div}(S)=\min_{u,v\in S:u\neq v}\textnormal{dist}(u,v)$ diversity objective.
-We design and analyze the GIST algorithm, which achieves a $\nicefrac{{1}}{{2}}$-approximation guarantee by solving a series of maximum independent set instances on geometric intersection graphs with the GreedyIndependentSet bicriteria approximation algorithm.
-We complement our algorithm with a $0.5584$-hardness of approximation.
-From the theoretical perspective, it is an interesting open problem to bridge the gap between the algorithmic and hardness results of $0.5$ and $0.5584$ approximations.
-For the special case of a linear utility function $g$, we show that GIST has a $\nicefrac{{2}}{{3}}$-approximation guarantee, and we complement this with a $(\nicefrac{{2}}{{3}}+\varepsilon)$-hardness of approximation, for any $\varepsilon>0$.
-Finally, we present an empirical study showing that GIST steadily outperforms baseline methods for MDMS (in particular greedy) on a simple set of Gaussian points with random weights.
-We conclude by comparing the top-1 image classification accuracy achieved by GIST and three other state-of-the-art data sampling methods for ImageNet, demonstrating the benefit of optimizing for a blend of utility and diversity.
-Appendix A $f(S)$ is Not Submodular
-Consider the instance on four points $N=\{a,b,c,d\}\subseteq\mathbb{R}$ where $a=0$, $b=1$, and $c=d=2$ in the (one-dimensional) Euclidean metric, i.e., all the points are collinear.
-Let $g(S)=0$ for all $S\subseteq V$.
-For $f(S)$ to be submodular, it must hold that for every $S\subseteq T\subseteq N$ and $x\in N\setminus T$,
-$f(S\cup\{x\})-f(S)\geq f(T\cup\{x\})-f(T).$
-However, if $x=b$, $S=\{a,c\}$ and $T=\{a,c,d\}$, we have:
-$f(S\cup\{x\})-f(S)$ $=f(\{a,b,c\})-f(\{a,c\})=1-2=-1$
-$f(T\cup\{x\})-f(T)$ $=f(\{a,b,c,d\})-f(\{a,c,d\})=0-0=0,$
-so $f(S)$ is not submodular.
-In this instance, the function $f$ is not monotone. However, a similar monotone but still not submodular $f$ can be defined by setting $g(S)=\sum_{v\in S}w(v)$ where $w(v)=2$ for every $v\in N$.
-Appendix B Proof of Theorem 3.4
-Lemma B.1.
-[…]
-Proof.
-Let $k^{\prime}=|T|$ and $t_{1},t_{2},\dots,t_{k^{\prime}}$ be the points in $T$ in the order that GreedyIndependentSet selected them.
-Let $B_{i}=\{v\in V:\textnormal{dist}(t_{i},v)<d^{\prime}\}$ be the points in $V$ contained in the radius-$d^{\prime}$ open ball around $t_{i}$.
-First, we show that each $B_{i}$ contains at most one point in $S_{d}^{*}$.
-If this is not true, then some $B_{i}$ contains two different points $u,v\in S_{d}^{*}$.
-Since $\textnormal{dist}(\cdot,\cdot)$ is a metric, this means
-$\textnormal{dist}(u,v)$ $\leq\textnormal{dist}(u,t_{i})+\textnormal{dist}(t_{i},v)$
-$<d^{\prime}+d^{\prime}$
-$\leq d/2+d/2$
-$=d,$
-which contradicts the assumption that $S_{d}^{*}$ is an independent set of $G_{d}(V)$.
-Note that it is possible to have $B_{i}\cap B_{j}\neq\varnothing$, for $i\neq j$, since these balls consider all points in $V$.
-Now let $C_{i}\subseteq V$ be the set of uncovered (by the open balls) points that become covered when GreedyIndependentSet selects $t_{i}$.
-Concretely, $C_{i}=B_{i}\setminus(B_{1}\cup\cdots\cup B_{i-1})$.
-Each $C_{i}$ contains at most one point in $S_{d}^{*}$ since $|B_{i}\cap S_{d}^{*}|\leq 1$.
-Moreover, if $s\in C_{i}\cap S_{d}^{*}$, then $w(t_{i})\geq w(s)$ because the points are sorted in non-increasing order and selected if uncovered.
-Let $A=C_{1}\cup\cdots\cup C_{k^{\prime}}$ be the set of points covered by the algorithm.
-For each point $s\in S_{d}^{*}\cap A$, there is exactly one covering set $C_{i}$ corresponding to $s$.
-It follows that
-$\sum_{s\in S_{d}^{*}\cap A}w(s)\leq\sum_{i\in[k^{\prime}]:S_{d}^{*}\cap C_{i}\neq\varnothing}w(t_{i}).$ (3)
-It remains to account for the points in $S_{d}^{*}\setminus A$.
-Note that if we have any such points, then $|T|=k$ since the points in $S_{d}^{*}\setminus A$ are uncovered at the end of the algorithm.
-Further, for any $t_{i}\in T$ and $s\in S_{d}^{*}\setminus A$, we have $w(t_{i})\geq w(s)$ since $t_{i}$ was selected and the points are sorted by non-increasing weight.
-Therefore, we can assign each $s\in S_{d}^{*}\setminus A$ to a unique $C_{i}$ such that $C_{i}\cap S_{d}^{*}=\varnothing$.
-It follows that
-$\sum_{s\in S_{d}^{*}\setminus A}w(s)\leq\sum_{i\in[k^{\prime}]:S_{d}^{*}\cap C_{i}=\varnothing}w(t_{i}).$ (4)
-Adding the two sums together in Equation 3 and (4) completes the proof. ∎
-See 3.4
-Proof of Theorem 3.4.
-Let $d^{*}$ be the minimum distance between two distinct points in $S^{*}$.
-There are two cases: $d^{*}<\varepsilon d_{\max}$ and $d^{*}\geq\varepsilon d_{\max}$.
-In the first case, outputting the $k$ heaviest points (Line 2) achieves a $(1-\varepsilon)$-approximation.
-To see this, first observe that
-$\textnormal{OPT}\geq d_{\max}>\frac{d^{*}}{\varepsilon}\implies\varepsilon\cdot\textnormal{OPT}>d^{*}.$
-The sum of the $k$ heaviest points upper bounds $g(S^{*})$, so we have
-$\textnormal{ALG}\geq g(S^{*})=\textnormal{OPT}-d^{*}>(1-\varepsilon)\textnormal{OPT}.$
-Therefore, we focus on the second case where $d^{*}\geq\varepsilon d_{\max}$. GIST considers a threshold $d=(1+\varepsilon)^{i}\cdot\varepsilon d_{\max}/2\in D$ such that
-$\frac{d^{*}}{2}\in[d,(1+\varepsilon)d)\implies\frac{1}{1+\varepsilon}\cdot\frac{d^{*}}{2}<d\leq\frac{d^{*}}{2}.$
-[…]
-$f(T)$ $\geq g(S^{*})+d\geq g(S^{*})+\frac{d^{*}}{2(1+\varepsilon)}.$ (5)
-The potential max-diameter update to $S$ on Line 5 gives us
-$\textnormal{ALG}\geq d_{\max}\geq d^{*}.$ (6)
-Putting (5) and (6) together, for any $0\leq p\leq 1$, it holds that
-ALG $\geq p\cdot\left(g(S^{*})+\frac{d^{*}}{2(1+\varepsilon)}\right)+(1-p)\cdot d^{*}$
-$=p\cdot g(S^{*})+\left(1-p+\frac{p}{2(1+\varepsilon)}\right)\cdot d^{*}.$
-To maximize the approximation ratio as $\varepsilon\rightarrow 0$, we set $p=2/3$ by solving $p=1-p/2$. It follows that
-ALG $\geq\frac{2}{3}\cdot g(S^{*})+\left(1-\frac{2}{3}+\frac{1}{3(1+\varepsilon)}\right)\cdot d^{*}$ (7)
-$=\frac{2}{3}\cdot g(S^{*})+\frac{1}{3}\left(1+\frac{1}{1+\varepsilon}\right)\cdot d^{*}$
-$\geq\frac{2}{3}\cdot g(S^{*})+\frac{1}{3}\left(2-\varepsilon\right)\cdot d^{*}$
-$\geq\left(\nicefrac{{2}}{{3}}-\varepsilon\right)\cdot\textnormal{OPT},$
-which completes the proof. ∎
-Appendix C Missing Analysis for Section 4
-C.1 Proof of Theorem: 4.2
+D.2 Proof of Theorem 4.2
 See 4.2
 Proof.
-[…]
+First, recall that a clique is a subset of vertices in an undirected graph such that there is an edge between every pair of its vertices.
+Håstad [28] and Zuckerman [64] showed that the maximum clique problem does not admit an $n^{1-\theta}$-approximation for any constant $\theta>0$, unless $\textnormal{NP}=\textnormal{P}$.
+This implies that there is no constant-factor approximation algorithm for maximum clique.
+In other words, for any constant $0<\delta\leq 1$, there exists a graph $G$ and a threshold integer value $k$ such that it is NP-hard to distinguish between the following two cases:
   • YES instance: graph $G$ has a clique of size $k$.
@@
 The distance between a pair of points is $2$ if there is an edge between their corresponding vertices in $G$, and it is $1$ otherwise.
-Use the same threshold value of $k$ (in the YES and NO instance above) for the cardinality constraint on set $S$, and set each weight $w(v)=1$.
+Use the same threshold value of $k$ (in the YES and NO instance above) for the cardinality constraint on set $S$, and set each weight $w(v)=\nicefrac{{\alpha}}{{k}}$ for some parameter $\alpha$ that we set later in the proof.
+We also set $\lambda=1-\alpha$.
 In a YES instance, selecting a clique of size $k$ as set $S$ results in the maximum possible value of the objective:
@@
 $=\max\left\{1,2-(2-\delta)\alpha\right\}.$
-These two terms become equal for $\alpha=1/(2-\delta)$. Thus, the gap between the maximum value any algorithm can achieve in the NO case and the optimum value in the YES case is
+We make these two terms equal by setting $\alpha=1/(2-\delta)$. Thus, the gap between the maximum value any algorithm can achieve in the NO case and the optimum value in the YES case is
 $\frac{1}{2-\alpha}=\frac{1}{2-\nicefrac{{1}}{{(2-\delta)}}}=\frac{2-\delta}{3-2\delta}.$
@@
 By setting $\delta<9\varepsilon/(1+6\varepsilon)$, we satisfy the required inequality and achieve the inapproximability gap in the theorem statement. ∎
-C.2 Proof of Lemma 4.4
+D.3 Proof of Lemma 4.4
 Let $G=(V,E)$ be a simple undirected graph.
@@
 Now suppose that $\{u,v\}\in E$. It follows that
-$\lVert h_{G}(u)-h_{G}(v)\rVert_{2}^{2}$ $=\sum_{e\in N^{\prime}(u)\setminus\{v\}}\frac{1}{2\deg^{\prime}(u)}+\sum_{e\in N^{\prime}(v)\setminus\{u\}}\frac{1}{2\deg^{\prime}(v)}+\left(\sqrt{\frac{1}{2\deg^{\prime}(u)}}-\sqrt{\frac{1}{2\deg^{\prime}(v)}}\right)^{2}$
+$\lVert h_{G}(u)-h_{G}(v)\rVert_{2}^{2}$
+$=\sum_{e\in N^{\prime}(u)\setminus\{v\}}\frac{1}{2\deg^{\prime}(u)}+\sum_{e\in N^{\prime}(v)\setminus\{u\}}\frac{1}{2\deg^{\prime}(v)}+\left(\sqrt{\frac{1}{2\deg^{\prime}(u)}}-\sqrt{\frac{1}{2\deg^{\prime}(v)}}\right)^{2}$
 $=\frac{\deg^{\prime}(u)-1}{2\deg^{\prime}(u)}+\frac{\deg^{\prime}(v)-1}{2\deg^{\prime}(v)}+\left(\frac{1}{2\deg^{\prime}(u)}+\frac{1}{2\deg^{\prime}(v)}-2\sqrt{\frac{1}{4\deg^{\prime}(u)\deg^{\prime}(v)}}\right)$
@@
 which completes the proof. ∎
-C.3 Proof of Theorem 4.5
+D.4 Proof of Theorem 4.5
 See 4.5
 Proof.
-[…]
+We build on the hardness of approximation for the maximum independent set problem for graphs with maximum degree $\Delta=3$.
+Alimonti and Kann [4, Theorem 3.2] showed that this problem is APX-complete, so there exists an $\varepsilon_{0}>0$ such that there is no polynomial-time $(1-\varepsilon_{0})$-approximation algorithm unless $\textnormal{NP}=\textnormal{P}$.
+Hence, there exists a graph $G$ with max degree $\Delta=3$ and a threshold integer value $k$ such that it is NP-hard to distinguish between the following two cases:
   • YES instance: graph $G$ has an independent set of size $k$.
@@
 We reduce this instance of bounded-degree maximum independent set to MDMS with objective function (2) as follows.
-Embed each node of the graph $G$ into Euclidean space using the function $h_{G}(v)$ in Lemma 4.4.
-We use the same threshold value of $k$ (between YES and NO instances above) for the cardinality constraint on set $S$, and we set each weight $w(v)=1$.
+Embed each node of the graph $G$ into Euclidean space using the function $h_{G}(v)$ in Lemma˜4.4.
+We use the same threshold value of $k$ (between YES and NO instances above) for the cardinality constraint on set $S$, and we set each weight $w(v)=\nicefrac{{\alpha}}{{k}}$ for some parameter $\alpha$ that we set later in the proof.
+We also set $\lambda=1-\alpha$.
 In a YES instance, selecting an independent set of size $k$ as the set $S$ results in the maximum value of objective (2):
@@
 since $\lVert h_{G}(u)-h_{G}(v)\rVert_{2}=1$ for any two distinct points $u,v\in S$ since there is no edge between $u$ and $v$ in graph $G$.
-In a NO instance, the best objective value that can be achieved in polynomial-time is the maximum of the following two scenarios: (a) selecting $k$ points with minimum distance at most $81-1/(2(\Delta+1))=1-1/$ 8, or (b) selecting at most $(1-\varepsilon_{0})k$ vertices forming an independent set with minimum distance equal to $1$.
+In a NO instance, the best objective value that can be achieved in polynomial-time is the maximum of the following two scenarios: (a) selecting $k$ points with minimum distance at most $1-1/(2(\Delta+1))=1-1/8$, or (b) selecting at most $(1-\varepsilon_{0})k$ vertices forming an independent set with minimum distance equal to $1$.
 The maximum value obtained by any polynomial-time algorithm is then
@@
 $=\max\left\{1-\varepsilon_{0}\cdot\alpha,(7+\alpha)/8\right\}.$
-These two terms become equal for $\alpha=1/(1+8\varepsilon_{0})$.
+We make these two terms equal by setting $\alpha=1/(1+8\varepsilon_{0})$.
 Therefore, the gap between the maximum value any algorithm can achieve in the NO case and the optimum value in the YES case is upper bounded by
@@
 Since $\varepsilon_{0}>0$ is a constant, $\varepsilon_{1}\coloneqq\varepsilon_{0}/(\nicefrac{{1}}{{\varepsilon_{0}}}+8)>0$ is also a constant. This completes the proof of APX-completeness. ∎
-Appendix D Additional Details for Section 5
-D.1 Synthetic Dataset
-We extend our comparison of baseline algorithms in Section 5.1 by sweeping over values of $\alpha,\beta$ in the objective function.
+Appendix E Additional details for Section 5
+E.1 Synthetic dataset
+We extend our comparison of baseline algorithms in Section˜5.1 by sweeping over values of $\alpha,\beta$ in the objective function.
 Figure 2: Baseline comparison with $\alpha\in(0.85,0.90,0.95,1.00)$ and $\beta=0.75$.
 Figure 3: Baseline comparison with $\alpha=0.95$ and $\beta\in(0.60,0.70,0.80,0.90)$.
-D.2 Image Classification
+E.2 Image classification
 Hyperparameters for ImageNet classification.
-We generate predictions and embeddings for all points using a coarsely-trained ResNet-56 model (He et al., 2016) trained on a random 10% subset of ImageNet (Russakovsky et al., 2015).
+We generate predictions and embeddings for all points using a coarsely-trained ResNet-56 model [29] trained on a random 10% subset of ImageNet [54].
 We use SGD with Nesterov momentum 0.9 with 450/90 epochs.
@@
 We extract the penultimate layer features to produce $2048$-dimensional embeddings of each image.
-We use the same hyperparameters as the original ResNet paper (He et al., 2016) with budgets and one-shot subset selection experiments designed in the same manner as Ramalingam et al.
-(2021).
+We use the same hyperparameters as the original ResNet paper [29] with budgets and one-shot subset selection experiments designed in the same manner as [49].
+Running times.
+The end-to-end running time is dominated by ImageNet model training, which takes more than a few hours even with accelerators (e.g., GPU/TPU chips).
+The subset selection algorithms that use margin and submodular sampling range between 3–4 minutes per run on an average.
+GIST subset selection is similar to the margin or submodular algorithms for a given distance threshold $d$.
+By using parallelism for different $d$ values, we can keep the GIST-submod subset selection runtime the same as the submod algorithm.
+In summary, the actual subset selection algorithm step is extremely fast (nearly negligible) compared to the ImageNet training time.
+Furthermore, we can exploit distributed submodular subset selection algorithms that can even handle billions of data points efficiently [15].